A Bibliography of Papers in Lecture Notes in Computer Science (2012): Volumes 7250–7299

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.01

Title word cross-reference

2 [108, 246, 122, 51]. 2^{n} [122]. 3 [346, 246, 351, 605, 659, 653, 677]. 4 [256, 554]. ^{e} [838]. ^{F} [191]. ^{i} [838]. b [485]. C+ [270]. E_{HQ^{+}}^{DDL}SHIQ [907]. k [723]. F_{2^{2k}} [554]. N [326]. O(1/ϵ^{2})^{n} [123]. p [553]. s [923]. T OY [838].

*and [815].

-ary [553]. -Center [108]. -Coloring [351]. -Connected [659]. -convexity [723]. -D [51]. -Disjoint [122]. -DNNF [923]. -Exemplar [807]. -Finger [605]. -Gram [326]. -Matching [485]. -modems [723]. -Satisfiability [677]. -Time [123]. -Uniform [554].

000* [815].

10th [978, 952, 957, 977]. 11th [989, 990, 976, 967, 968, 993]. 12th [972]. 14th [973, 960, 974]. 160 [163]. 16th [963].

23rd [958]. 2S [350].

32nd [973].

4th [966].

60th [961]. 65th [261]. 6th [988, 964, 979].

7th [984, 962, 980, 995, 996, 986].

802.11 [286]. 802.16 [506]. 8th [954, 992].

9th [987, 997, 994].

A* [156]. AAIM [985]. AAMAS [955]. Ablation [249]. above [121]. Abstract [187, 83, 409, 192, 885]. Abstracting [381]. Abstraction [12]. Abstractions [378]. Acceleration [536]. Accelerometer [879]. Access [779, 556, 507]. Accuracy [229, 630]. Accurately [820]. ACO [900]. Acoustics [614]. Acquisition [583, 613]. across [228]. Act [371, 643]. ACTHEX [275]. Actigraphy [41]. Action [270, 272, 276, 695]. Actions [270, 645, 643]. ActionScript [127]. Activation [308, 17]. Active [458, 395, 864, 46, 603]. Activities [37, 933]. Activity [32, 879, 925, 641]. Activity-Based [925]. Actor [433, 422]. Actors [74, 432]. Actuated [599]. Actuator [882]. Ad [423, 501, 503, 441, 500]. Ad-Hoc [503, 441, 500]. Adapt [606]. Adaptation [79, 763]. Adaptive [495, 370, 348, 485, 88, 556, 877, 534, 48, 441, 870, 215]. Addition [462]. Adhoc [289]. Admission [533]. Admixtures [226]. Adopting [590]. Adoption [24, 27, 776]. Adult [639]. Adults [28]. Advanced [438, 939]. Advances [986, 995]. Adversative [143]. Advertising [690]. Advice [111, 436]. Advising [229]. AE [964]. AE-CAI [964]. AES [448]. AES-Enabled [448]. Affect [481, 633]. Affective [45]. Affects [635]. Affordances [692]. after [772]. Again [404]. against [391, 178, 451]. Age [854]. Agency [644]. Agent [955, 374, 368, 87, 89, 88, 373, 644, 375, 887, 86]. Agent-Based [89]. Agents [93, 368, 367, 442, 91]. Agglomeration [47]. Aggregation [289]. Agreeable [107]. Agreement [366, 118]. Agreements [773]. Agricultural [882]. AI [997, 965, 996]. Aids [258]. AIMS [979]. Aircraft [655]. Alarm [634]. Alert [295]. Algebraic [559, 178, 463]. Algorithm [107, 264, 917, 811, 478, 123, 349, 350, 351, 233, 232, 156, 176, 354, 335, 312, 314, 238, 871, 926, 877, 909, 358, 870, 890, 500, 324]. Algorithm-Based [877]. Algorithmic [813, 985]. Algorithmics [985]. Algorithms [951, 502, 44, 120, 814, 328, 482, 483, 721, 816, 716, 666, 976, 988, 591, 844, 668, 469, 709, 155, 158]. Aligning [652]. Alignment [234, 256]. Alignments [229]. All-Different [914]. Allocation [495, 880, 784, 870]. Almost [553]. ALSIP [959, 155]. Alteration [240]. Alternative [824]. Ambient [40, 22]. Ambients [435]. American [957]. among [267]. Analysis [322, 411, 798, 24, 516, 517, 220, 66, 210, 363, 170, 809, 215, 519, 44, 120, 520, 978, 952, 521, 56, 248, 491, 523, 901, 52, 102, 151, 217, 714, 574, 538, 336, 820, 177, 528, 60, 451, 549, 437, 242, 460, 81, 161, 162, 975, 944, 10, 402, 172, 165, 457]. Analytics [581, 815]. Analyzing [11, 653]. Ancestry [226]. ANFIS [349]. Angiography [256]. Angle [597]. Annotation [57, 867]. Annual [987, 963]. Anomalies [545]. Anomaly [544]. Anonymisation [541]. Anonymity [541]. Answer [263]. Answering [892]. Answers [141, 835, 273]. Ant [357, 890]. Ants [344, 437]. Aortic [254]. Aperiodic [565]. Apicalis [344]. APIs [681]. Application [429, 264, 300, 913, 57, 958, 614, 151, 310, 398, 45, 872, 890]. Applications [222, 987, 346, 992, 117, 403, 276, 769, 972, 980, 212, 883, 895, 996, 709, 994, 646, 586, 402, 887, 951]. Applied [605, 359]. Applying [687]. Approach [806, 595, 263, 34, 912, 265, 187, 143, 364, 101, 482, 279, 533, 542, 234, 506, 524, 905, 79, 371, 654, 872, 893, 534, 575, 690, 711, 795]. Approaches [815, 333]. Approximate [123]. Approximating [515]. Approximation [370, 110, 300, 673, 114, 273, 716, 340, 719]. Approximations [718]. Apps [27]. April [963, 957, 967, 968, 969, 966]. Apriori [806]. Arabic [943]. Arc [479]. Arc-Flags [479]. Architectural [294, 451, 82]. Architecture [33, 293, 605, 240, 687, 296, 578, 579, 287]. Architectures [574]. Area [495, 732, 617]. Aren’t [513]. Arequipa [957]. Arithmetic [557]. Arrays [912]. Arrows [839, 137]. Art [618, 723]. Articulation [518, 484]. Artifacts [832, 601, 690]. Artificial [95, 351, 609, 309, 644, 959, 996, 967, 968]. Artiminio [952]. ary [553]. ASD [642]. ASes [774]. Asia [998]. Asian [933]. Asil [127]. ASP [280]. Aspect [380, 971, 382, 379]. Aspect-Aware [382]. Aspect-Oriented [380, 971]. Aspects [954, 201, 985]. Assembly [981, 379]. Assess [820]. Assessing [860]. Assessment [679, 24, 37, 36, 795]. Assessments [687]. Assignment [506]. Assist [639]. Assistance [606]. Assistant [77]. Assisted [40, 349, 742, 964, 22]. Assistive [38]. Association [330, 775, 244]. Associations [7, 197]. Associative [309]. Assume [712]. Assume/Guarantee [712]. Assumption [179]. Assured [932]. Astronomical [302]. Asymptotic [567]. Asynchronous [719]. Asynchronously [73]. Attack [458, 472, 454, 176, 456, 463, 462, 461]. Attacking [453, 448]. Attacks [752, 178, 450, 166, 177, 451, 455, 449, 165]. Attribute [188]. Attributes [335]. Audience [854]. Audiences [849]. Audition [625]. Auditory [633]. Augmentation [610]. Augmented [248, 249, 630, 964]. August [955]. Autism [638, 602, 893]. Auto [313]. Auto-selection [313]. AutoDunt [170]. Automata [210, 410, 414, 375]. Automated [206, 828]. Automatic [866, 295, 467, 713, 57]. Automation [538]. Automotive [283]. Autonomic [429, 877]. Autonomous [543, 979]. Autoregressive [55]. Availability [494]. Average [121]. Aware [496, 671, 793, 382, 756, 788, 91, 759, 587, 887]. Awareness [142, 390, 586, 399]. Axelrod [96]. AZP [900]. AZP-ACO [900].

B [401]. Back [62, 84, 98, 104, 138, 202, 259, 297, 376, 385, 407, 425, 445, 465, 550, 593, 804, 949, 172]. Backdoors [775]. Backhaul [788]. Backup [747]. Bacterial [353, 240]. Bagged [909]. Balance [29]. Balancing [535]. Ball [233]. Ball-Based [233]. Ballast [233]. band [473]. Bandwidth [495, 668, 547]. Bankruptcy [898]. Barcelona [963]. Based [322, 496, 285, 470, 911, 595, 400, 548, 300, 323, 265, 458, 416, 66, 917, 880, 363, 809, 640, 919, 856, 920, 2, 801, 965, 368, 605, 59, 533, 250, 781, 816, 876, 624, 52, 233, 924, 540, 613, 542, 381, 943, 100, 777, 89, 886, 77, 414, 333, 134, 150, 335, 442, 905, 314, 27, 70, 879, 871, 942, 168, 337, 926, 877, 451, 927, 925, 242, 632, 296, 135, 929, 893, 428, 286, 873, 867, 763, 861, 153, 130, 8, 218, 289, 214, 61, 620, 505, 712, 375, 172, 768, 887, 795, 806]. Basic [566, 831, 141, 517]. Batch [820]. Baxter [658]. Bayesian [49, 819, 46]. BCNF [188]. BDDs [913]. BDI [93]. Beacons [777]. Beam [732, 930]. Bed [488]. Bee [351]. Before [503]. Begging [141]. Behavior [650]. Behavioral [392]. Behaviour [248, 637]. Behavioural [412, 444, 79]. Beijing [987, 985]. Belgium [978]. Belief [272]. bent [555]. Bernhard [961]. Best [608, 355]. Better [110, 475, 620]. Between [604, 622, 156, 525, 662]. Between-Person [604]. Beyond [414]. BGP [761]. Bi [278]. Bi-state [278]. Biased [822]. Bichromatic [108]. Biclustering [338]. Bidirectional [55]. Bigraphical [82]. Bilinear [179]. Bimanual [616]. Binary [568, 658, 312, 560]. Binder [818]. Binding [818]. Binomial [360]. Bioinformatics [992]. Biological [337]. Biology [963, 813, 715]. Biomedical [346]. Biometric [651]. Biomolecular [227]. Bipartite [667, 669]. BiQL [11]. Birthday [261, 961]. Bisimulations [833]. Bisociation [2]. Bisociations [18]. Bisociative [1, 20, 3, 951]. BisoNet [12]. BisoNets [5]. BitTorrent [770, 767]. Black [927, 925]. Black-Box [925]. Blaming [144]. Blinding [453]. Block [456]. Bluetooth [498]. Board [283]. Body [495, 600, 355]. Boolean [559, 827, 178, 557, 560, 415, 558]. Boot [166]. Bordeaux [976]. Border [719]. Botnet [169]. Botnets [948]. Bound [125, 677]. Bounded [113, 210, 921, 414, 218, 859]. Bounded-Degree [113]. Bounds [717]. Box [725, 927, 925]. BPR [406]. BPR-Framework [406]. BPRules [406]. Braille [635]. Brain [521, 606, 258]. Brain-Computer [606]. Branch [481, 734, 467]. Branch-and-Price [467]. Break [727]. Breakage [237]. Breakage-Fusion-Bridge [237]. Brides [733]. Bridge [237]. Bridges [484]. Bridging [6, 534]. Brief [20]. Bringing [850]. Broadband [787]. Broadcasting [500]. Broker [393]. Buffer [672]. Bugs [206]. Build [546]. Building [538, 757, 900, 697, 372]. Bumper [631]. Bus [291]. Business [792, 413, 205, 528, 687, 796, 685, 849]. Butterfly [462]. Butterfly-Attack [462]. By-Example [196]. Byzantine [733].

Cache [869, 744, 120]. Caching [745, 532]. CAI [964]. Cake [725].

Calculational [844]. Calculi [832]. Calculus [264, 840, 705, 712]. Call [830, 832, 681, 755]. Call-by-Need [832]. Call-by-Value [830]. Call-Stateful [755]. Camera [59]. Can [848, 511, 633, 219, 603]. Canada [980, 964]. Cancer [228, 249, 237, 241]. Canonical [131]. Capturing [840]. Car [908, 287]. Car2X [284]. Cards [449]. Care [26, 761]. Carlo [352]. Carriers [792]. Cartification [512]. Case [150, 461, 631, 27, 402, 802, 677, 683]. Catalog [135]. Categorization [944]. Causal [268]. Cause [885]. Caused [548, 621]. Cayley [109]. CCA [179]. CDCL [211]. Cellular [813, 282, 491, 296, 375]. Center [108]. Centers [651, 756]. Centrality [769]. Centred [22]. Centres [696]. Centric [34, 744, 745, 743, 574, 742, 932]. Certain [553]. CFLP [836]. Chain [226, 895, 853]. Challenge [788]. Challenge-Aware [788]. Challenged [642]. Challenges [384, 282, 530, 198, 130, 207, 514, 22, 361]. Chance [919]. Change [650, 637]. Channel [601, 450, 447, 494, 975, 463]. Channels [494, 507]. Characteristics [801, 219]. Characterization [915, 771, 172, 165]. Characterizing [772, 558]. Charge [227]. Checking [414, 413, 81, 218]. Chervonenkis [333]. CHF [39]. Children [602]. China [987, 995, 985]. Chinese [942]. CHIRON [31]. Choice [608]. Choosing [685]. Choreographies [73]. Chronic [24, 952, 27]. CiAO [382]. Cipher [456]. Ciphers [451]. Ciphertext [946]. Circuit [893]. Circuits [923]. Circular [114]. Cities [588]. Citizen [153]. Clairvoyance [208]. Class [383, 337, 46, 90]. Classic [640]. Classical [903, 832]. Classification [749, 115, 329, 47, 371, 909, 46]. classifier [371]. Classifiers [898, 897]. Classroom [135]. Clauses [677]. Client [391, 491, 777]. Client-Side [391]. Clinical [35, 253, 31]. CLIPS [896]. Cloaking [194]. Clock [634]. Clone [525]. Closed [157]. Closer [459]. Closing [648, 575]. Cloud [285, 429, 869, 754, 880, 881, 491, 755, 217, 393, 871, 872, 873, 932, 532, 874]. Cloud-Centric [932]. Clouds [870]. CLP [837]. Cluster [289, 341]. Clustering [331, 332, 344, 335, 442, 823, 341, 359]. Clusters [670]. Co [857]. Co-solvers [857]. Coalitional [786]. Code [221, 401]. Coefficient [562]. Coevolutionary [354]. Coffee [737]. Cognitive [363, 276, 503, 909, 535, 592, 603, 786]. Coherence [78]. COIN [955]. Coinductive [846]. Cold [166]. Collaboration [543, 89, 427]. Collaborative [745, 686, 906]. Collection [7]. Collections [5]. Collective [703, 970, 643]. Colloquial [749]. Colonoscopy [252]. Colony [351, 357, 890]. Coloring [914, 116, 351]. Colour [246]. Colour-Consistency [246]. Colouring [431]. Column [927]. Combination [682]. Combinations [895]. Combinatorial [997, 919, 930]. Combinatorics [237]. Combined [912]. Combining [606, 865, 688]. Commerce [367]. Commitment [945]. Committee [897]. Committer [945]. Communication [293, 423, 982, 983, 977, 689, 966]. Communications [571]. Community [746, 520, 709]. Compact [474, 353]. Company [546]. Comparative [337]. Compare [817]. Comparison [471, 544, 856, 498, 333, 824]. Compatible [327]. Compensable [434].

Compensator [598]. Competing [359]. Competition [779]. Competitive [890]. Compiler [130]. Compiling [831]. Complete [915, 414, 472]. Complex [226, 753, 241, 802]. Complexity [107, 111, 568, 121, 333, 334, 64, 569, 567, 849]. Compliance [92]. CompNets [991]. Component [954, 66, 809, 70, 48]. Component-Based [66, 70]. Components [80, 68, 83, 71, 77]. Composing [65]. Composite [270]. Composition [877]. Compositional [712]. Compositionally [413]. Compositions [383, 406]. Compound [54]. Comprehensive [679, 406]. Compressed [162]. Compression [14]. Computability [855]. Computation [987, 701, 778, 702, 843, 969, 146]. Computational [917, 179, 963, 813, 970, 953]. Computationally [314]. Compute [401]. Compute-Intensive [401]. Computer [596, 735, 606, 964, 750]. Computer-Assisted [964]. Computers [704]. Computing [429, 400, 396, 559, 880, 667, 519, 403, 476, 811, 484, 156, 995, 877, 708, 872, 967, 968, 197, 192, 158, 874, 706]. Concentrated [450]. Concentration [36]. Concept [515, 516, 517, 951, 519, 856, 520, 978, 2, 521, 331, 6, 526, 16, 527, 528, 861]. Concept-Based [861]. Concepts [517, 511, 522, 513, 195]. Conceptual [294, 199, 198, 197, 961]. Concurrency [206]. Concurrent [411, 216]. Conditional [336, 58]. Conference [987, 984, 997, 989, 990, 963, 978, 952, 962, 973, 972, 980, 982, 983, 960, 977, 988, 995, 996, 986, 967, 968, 979, 994, 974, 985]. Configuration [866, 926, 862]. Conflict [416, 729]. Conflict-Free [729]. Conformance [72, 415]. Confounded [820]. Congestion [533]. Conjoined [145]. Conjunction [918, 964, 969]. Conjunctive [946]. Connect [704]. Connected [122, 659]. Connection [525, 505]. Connectivity [296]. Connector [431]. Conquer [211]. Consecutive [668]. Consensus [814, 118]. Conservation [645]. Conservative [263, 881]. Conserved [244]. Consideration [870]. Considerations [272, 886]. Consistency [246]. Consistent [152]. Consistently [392]. Constant [946, 716]. Constant-Time [716]. Constrained [113, 886, 882, 844]. Constraint [916, 836, 921, 905, 925, 929, 891, 846]. Constraint-Based [929]. Constraints [912, 269, 729, 918, 919, 433, 183]. Constructing [6]. Construction [109, 718]. Constructions [174]. Constructive [975]. Constructs [132]. Consultation [24]. Consumption [760]. contact [41]. Contactless [458]. Content [322, 798, 745, 766, 747, 779, 770, 336, 742, 859, 586]. Content-Based [322]. Contention [507]. Contests [680]. Context [749, 381, 586, 653, 887, 274]. Context-Aware [887]. Context-Sensitive [749]. Contexts [522]. Contiguous [360]. ContikiRPL [537]. Continuous [468, 357]. Continuum [346, 26]. Contractor [911]. Contracts [443]. Contraint [997]. Control [429, 833, 605, 533, 538, 314]. Controller [908]. Controlling [78]. Convergence [305, 306, 304, 707, 318, 317]. Conversion [452]. Convex [106, 911, 667, 475, 669, 659]. convexity [723]. Cooperation [33, 785, 769, 780]. Cooperative [503, 532]. Cooperativity [141]. Coordinated [83, 432]. Coordinating [435, 444]. COORDINATION [974, 797, 531, 438, 88, 882, 427, 86, 440, 428, 436, 441, 955, 974]. Core [891]. Coronary [256]. Corporate [898, 692]. Corpus [326]. Correct [839, 965]. Correcting [313, 166]. Correction [255]. Correlated [185]. Correlation [561, 571, 552, 565]. COSADE [975]. Cost [638, 81, 505, 897]. Cost-Sensitive [897]. Costs [492]. Counter [449]. Countermeasure [451]. Countermeasures [454]. Counting [740, 815, 103]. Coupled [230]. Coupling [612]. Cover [471, 110, 8]. Coverage [220]. Covering [676, 663, 333]. CPAIOR [997]. Crawling [746]. Creation [4, 45, 194]. Creative [2]. Creativity [691]. Credibility [935]. Crete [994]. Criminals [511]. Critical [530]. Cross [561, 577, 884]. Cross-Correlation-II [561]. Cross-Disciplinary [577]. Cross-Layer [884]. crossing [667]. Crossings [727, 295]. Crossover [360]. Crossroads [848, 618]. Crossword [735]. CRT [453]. Cryptanalysis [163]. Cryptographic [172]. Cryptographically [940]. Cryptology [960]. CSP [712]. CT [254]. CT-US [254]. CTA [256]. Cube [211]. CubeHash [177]. Cubes [522]. CUDA [303]. Cues [609]. Culture [201]. Cumulative [916]. Current [598, 171]. Cursor [596]. Curvature [595, 621]. Curves [555, 460, 175]. Custom [37]. Customer [608]. Cut [660]. Cutaneous [617, 620]. Cutting [854]. Cyber [776]. Cyber-Insurance [776]. Cybercrime [934]. CybercrimeIR [934]. Cyberspace [933]. Cycles [941]. Cyclic [916, 839]. Cylindrical [627]. Czech [991, 989, 990].

D [246, 346, 246, 51, 256, 653]. D-RAB [653]. D/ [246]. Dag [472]. Daily [32]. DAIS [972]. Dallas [992]. Dark [509]. Darmstadt [975]. Data [495, 222, 68, 940, 302, 856, 510, 328, 939, 288, 20, 860, 330, 855, 232, 236, 100, 331, 332, 185, 526, 895, 344, 527, 905, 757, 863, 337, 45, 756, 194, 853, 713, 654, 845, 861, 289, 195, 857, 858, 192, 874]. Data-Based [856]. Data-Driven [654]. Data-Parallel [192]. Database [184, 190, 938, 532]. Databases [189, 850, 157, 185]. Datacenter [490]. Dataflow [78]. Datalog [266]. Day [170]. Deadlines [107]. Dealing [849]. Debate [854]. Debugging [835]. December [962, 960, 959, 953]. Decentralized [781, 852]. Decision [696, 685, 505, 897]. Declarative [834, 835, 510]. Declaratives [144]. Decomposition [467, 356]. Dedicated [961]. Dedication [182]. Deductive [865]. Default [274]. Defining [419]. Definitions [590]. Deflection [612]. Deformable [613]. Deformation [250]. Deformity [255]. Degradation [541]. Degree [113, 740, 753]. Deimmunization [243]. Delays [357, 164]. delegatable [175]. Deliberation [153]. Delimited [833]. Delimited-Control [833]. Delivery [40]. Demand [880, 764, 759]. Denotational [66]. Dense [673]. Density [115, 125]. Dependable [979]. Dependence [170]. Dependencies [795]. Dependency [322, 242]. Dependent [193]. Dependently [829]. Dependently-Typed [829]. Deployment [400, 75, 217]. DEPSO [350]. DEPSO-2S [350]. Deques [834]. Derived [568, 560]. Describing [136]. Descriptions [718]. Descriptively [102]. Design [875, 384, 984, 754, 293, 187, 618, 693, 459, 792, 681, 627, 682, 883, 448, 294, 683, 684, 239, 382, 686, 398, 694, 877, 137, 687, 654, 578, 975, 698, 587, 689, 691, 653, 695, 180, 986]. Designated [175]. Designed [641]. Designing [638, 403, 680, 682, 697, 646, 688]. Designing-in-the-Large [688]. Desktop [875]. DESRIST [986]. Detecting [746, 606, 904, 774, 775, 169, 545]. Detection [544, 325, 170, 939, 540, 542, 938, 943, 354, 937, 238, 206]. Determinants [637, 654]. Determine [251]. Determining [637]. Deterministic [109]. Develop [252, 684]. Developing [642, 490]. Development [580, 247, 607, 800, 971, 683, 398, 656, 872, 22]. Developments [813, 844]. Deviation [567]. Device [605, 608]. Devices [598, 982, 983, 886, 50, 800]. DFS [210]. Diabetic [639]. Diagram [661]. Dial [928]. Dial-a-Ride [928]. Dialectics [103]. Dialogue [371]. Diameter [476, 124]. Dickson [555]. Difference [447]. Differences [623, 603]. Different [914, 856, 334]. Differential [345, 348, 349, 350, 820, 356, 358, 359, 360, 361, 457]. Differentially [554]. Differentiation [674]. Diffie [179]. Digit [736]. Digital [680, 892, 854]. Dimension [333]. Dimensional [116, 124, 240, 522]. Direct [506]. Directed [476]. Direction [632, 621]. Directionally [51]. Directions [813]. Disambiguation [894]. Disaster [285]. Disciplinary [577]. Disconnected [396]. Discourse [143]. Discovery [806, 951, 1, 531, 72, 20, 7, 526, 16, 705, 241, 18, 862, 405, 155, 378, 23]. Discrete [346]. Discrete-Continuum [346]. Discriminant [901]. Discrimination [622, 609, 629]. Disease [24, 952, 27]. Dishonest [443]. Disjoint [122]. Display [627, 632]. Displays [600]. Dissemination [388]. Dissociation [629]. Distance [114, 807, 811, 156, 660, 319, 662, 158]. Distances [419]. Distinguishability [335]. Distress [937]. Distributed [789, 80, 748, 300, 543, 416, 531, 721, 485, 973, 747, 88, 432, 398, 387, 129, 64, 907, 389, 402, 874, 972]. Distributed-Memory [129]. Distribution [394, 742]. Distributions [567]. Diversity [791]. DLV [277]. DNA [100]. DNNF [923]. DNS [549]. Do [704]. Document [7, 151, 5]. Documents [943, 6]. Does [32, 364]. Domain [326, 589, 743, 791, 378, 797, 782, 772]. Domain-Driven [378]. Domain-Specific [326]. Domains [912, 228]. Domination [669]. Don’t [481]. Downloading [770]. DPL [450]. DPOR [214]. DPOR-Based [214]. Drawing [659]. Dreamland [634]. Drift [331]. Drill [628]. Driven [322, 531, 948, 654, 378]. Driver [287, 286]. Dual [475]. Duality [525]. Dummett [840]. Duplication [234, 821]. Dutch [696]. Dynamic [222, 380, 397, 531, 434, 170, 215, 479, 71, 622, 755, 100, 88, 284, 422, 146, 786, 507]. Dynamics [230, 142, 722, 802]. Dynamizing [486].

E-Commerce [367]. e-health [35, 638]. E-Learning [372]. EAGLE [864]. Earth [114]. Earthquake [772]. Ease [194]. EASEA [345]. Easier [739]. Eating [654]. EC [969]. ECG [36]. ECL [838]. Edge [727, 477, 14, 13]. Edged [597]. Edges [727]. Edit [156, 158]. Editing [54]. Educate [655].

Education [24]. Effect [626, 499, 154, 617]. Effective [280, 283, 872]. Effectively [854]. Effectiveness [649]. Effects [597, 28]. efficacy [28]. Efficiency [401, 288, 886, 787, 790]. Efficient [470, 114, 210, 880, 946, 664, 283, 314, 742, 879, 864, 164, 165]. Efficiently [761]. Efforts [35]. Egress [375]. Eigenvectors [753]. Elastic [872]. Elastically [613]. Elder [38]. Elderly [33, 639]. Electricity [645]. Electrocutaneous [627]. Electrode [258]. Electromagnetic [458]. Electron [715]. Electronic [35, 90]. Element [599]. Eliminating [164]. Elimination [760, 714]. Ellipsoidal [718]. Elliptic [460, 175]. Else [726]. Embedded [132, 316, 164]. Embedding [488]. Emerge [902]. Emerge-Sort [902]. Emergence [96]. Emergency [689, 935]. Emotional [602, 367, 937]. Emotions [602]. EmotiWord [45]. Emphatic [143]. Empirical [401, 656]. Employing [348, 655, 52]. Empowering [638]. Empowers [643]. eMule [169]. eMule-like [169]. Enabled [448, 535]. Enabling [702, 590]. Encourages [640]. Encryption [946, 176, 167, 168]. End [754, 726]. End-to-End [754]. Endoscopic [246, 257]. Energy [496, 107, 880, 881, 879, 787, 790, 759]. Energy-Aware [496, 759]. Engage [704, 641, 854]. Engine [896]. Engineering [140, 478, 368, 201, 694, 469, 759, 590]. Engines [283]. English [264]. Enhance [296]. Enhanced [255, 199, 463]. Enhancement [600, 696]. Enhancing [337, 767]. Enlargement [54]. Enriched [585]. Enriching [326]. Enrollment [947]. Ensemble [898]. Ensuring [651, 25]. Enterprise [687, 688]. Enterprise-Wide [688]. Entities [585]. Entropy [186]. Envelopes [475]. Environment [33, 580, 247, 305, 304, 310, 311, 38, 439]. Environments [638, 880, 871, 964, 656, 587, 532, 441, 23]. Envy [664]. Envy-Free [664]. EPE [450]. EPE-Immune [450]. Epigenetic [228]. Epileptic [258]. EQR [496]. Equal [620]. Equality [269]. Equational [279]. Equations [162]. Equilibrium [354]. Equivalence [271]. Equivalences [412]. Equivariance [117]. Era [815]. Erdos [109]. Erlang [424]. ERP [686]. Error [812, 463]. Errors [166]. Essays [961, 965]. Establishing [97]. Estimating [229]. Estimation [394, 547, 799]. Estimator [324]. ESWC [994]. ETICS [991]. EuroHaptics [982, 983]. Europe [853]. Evacuations [655]. Evaluating [69, 681, 41, 796]. Evaluation [148, 374, 20, 612, 883, 152, 504, 787, 497, 180]. Event [401, 381, 393]. Event-B [401]. Event-Based [381]. Events [128]. Everyday [37]. Everyone [726]. Evidence [600, 244]. Evolution [345, 348, 810, 349, 773, 350, 234, 803, 356, 878, 358, 359, 360, 361]. Evolutionary [811, 908, 821, 969]. Exact [814, 677]. Examination [689]. Examining [944]. Example [196]. Examples [135]. Exclusion [501, 842, 871]. Execution [420, 83, 418, 194]. Exemplar [807]. Exercise [29]. Exergame [640]. Existential [269]. Expanding [109]. Expectable [799]. Expectations [94]. Expected [31]. Experience [380, 794]. Experimental [498, 751, 612, 905, 976]. Expert [34]. Explain [603]. Explaining [920, 923]. Explanation [87]. Exploitation [463]. Exploiting [901, 501, 447]. Exploration [2, 522, 20, 19]. Exploratory [933]. Exploring [509, 650, 630]. Exponent [453]. Expression [820, 822, 194]. Expressions [145]. Extant [821]. Extended [994]. Extending [263, 270, 838]. Extension [785]. Extensional [200]. Extraction [865].

Facets [914]. Facial [225]. FACS [954]. Fact [193]. Fact-Oriented [193]. Factor [523]. Factorization [516]. Factors [242, 153]. Failures [548, 423, 79]. Fairness [128, 784]. Families [432, 554]. Family [382]. FAMoS [395]. Fan [632]. Fan-Based [632]. Fast [789, 470, 325, 811, 178, 895, 825]. Faster [122]. Fat [347]. Fault [80, 456, 219, 457]. Fault-Proneness [219]. Fault-Tolerant [80]. Faults [415]. FAW [985]. FAW-AAIM [985]. FCA [524, 525, 514]. FCI [143]. FD [836]. fdcc [912]. Feasibility [927]. Features [904, 942, 61, 53]. Federated [590]. Federating [591]. Federations [782]. Feedback [599, 645, 628, 648]. Feel [611]. FEmtocells [296, 505]. Few [184]. FIA [573]. Fields [58, 564, 558]. Fifth [149]. Fight [934]. File [888, 219, 389]. Filesystem [400]. Filesystem-Based [400]. Financial [897]. Finding [724, 737, 15, 857]. Finger [605, 635]. Fingering [349]. Fingertip [623]. Finite [912, 599, 561, 564, 558, 567]. Finland [982, 983]. Firefighters [739]. First [264, 383, 434, 196, 90]. First-Class [383, 90]. First-Order [434]. Fits [725]. Fitted [37]. Fitting [608]. Fixation [391]. Fixed [728, 291]. Fixpoint [273]. Flags [479]. Fleet [369]. Flexible [757, 395, 592]. Flexub [397]. Floating [223]. Floating-Point [223]. Floorplans [658]. FLOPS [993]. Flow [548, 915, 919, 920, 540, 542, 316, 845, 547]. Flow-Based [548, 919, 920, 540, 542]. Flow-Level [547]. Flowshops [358]. Fluid [437]. FMOODS [973]. fMRI [521]. FOL [264]. FootGlove [608]. Foraging [353, 437]. Force [616, 612]. Forces [623]. Forecast [880]. Forecasting [898, 897]. Foreground [59]. Forgetful [721]. Form [833, 559, 189]. Formal [517, 519, 520, 521, 75, 526, 528, 64, 954, 978, 973]. Formalization [193, 77]. Formalizing [270, 367]. Formation [468]. Forming [701]. Formulae [567]. Formulation [110]. FORTE [973]. Forward [601]. Forwarding [745]. Fostering [26]. Foulds [817]. Foundations [961, 189, 338, 147]. Fourth [189]. FPGA [450]. Fractional [51]. Frame [854]. Framework [584, 406, 89, 490, 528, 91, 697, 46, 907, 587, 646, 23, 837]. France [997, 955, 976]. Fraudulent [897]. Free [729, 478, 774, 664]. Freedom [309]. Frequencies [512]. Frequency [570, 629, 565]. Frequency-Hopping [570, 565]. Friction [611]. Front [21, 42, 426, 446, 466, 487, 508, 529, 551, 572, 594, 615, 636, 657, 678, 699, 720, 741, 762, 783, 805, 826, 847, 868, 889, 910, 931, 950, 63, 85, 99, 105, 126, 139, 160, 181, 203, 224, 245, 260, 281, 298, 299, 320, 321, 342, 343, 362, 365, 377, 386, 408]. Frontiers [959, 985]. Fruitful [25]. FSR [451]. FSR-Based [451]. FSToolkit [590]. Full [883, 355, 457]. Full-Body [355]. Fully [479, 168, 885]. FUN [988, 988]. Function [370, 559, 55, 371, 340, 218, 545]. Functional [831, 958, 127, 828, 136, 993]. Functionals [186]. Functions [263, 394, 178, 555, 308, 557, 560, 558, 565]. Fusion [237]. Future [981, 583, 580, 813, 582, 792, 589, 205, 577, 578, 25, 796, 405, 586, 573]. Fuzzifications [524]. Fuzzy [363, 533, 331, 332, 909, 861].

Gains [790]. Galleries [723]. Game [798, 724, 640, 730, 645, 786, 641]. Gamepad [602]. Games [903, 655, 722]. Gap [534, 575]. Garbage [848]. Garments [37]. Gathering [726, 347]. Gaussian [250]. Gaze [56]. Gb [536]. Gb/s [536]. Gearing [280]. Gene [816, 825]. General [837]. Generalization [106]. Generalized [137, 209]. Generate [688]. Generated [401, 5]. Generation [789, 875, 222, 922, 295, 926, 927, 131, 592, 455]. Generator [458, 718, 339]. Generators [223, 307, 459]. Generic [174, 442, 132]. Genetic [363, 877, 240, 864]. Genome [234, 240, 198]. Genomes [815, 231]. Genomic [241]. Genomics [237]. Geometric [52]. Geometrical [48]. Geometrically [199, 316]. Geometry [917]. Geometry-Based [917]. Germany [975]. Gestures [252]. GHC [829]. Giants [774]. Given [718]. Glass [518]. Global [347, 916, 364, 392]. GML [199]. Gödel [840]. Goes [765]. Golomb [564]. Good [737]. Goodbye [726]. Gossip [871]. Gossip-Based [871]. Government [696]. GPC [995]. GPU [346]. Gradable [147]. Gradable-Predicate [147]. Gradient [895]. Gram [326]. Graph [914, 729, 351, 863, 469]. Graphs [471, 109, 667, 476, 125, 480, 484, 740, 413, 15, 668, 316, 669, 659]. Great [706]. Greece [977, 996, 994]. Greedy [480, 358, 899, 659]. Grid [875, 400, 995]. Ground [626]. Grounder [277]. Group [227, 173, 439, 372]. GSM [801]. GSM-Based [801]. GTP [816]. Guarantee [712]. Guards [662]. Guidance [254, 25]. Guided [210, 249, 243]. Guiding [211].

Haifa [962]. Hamming [660, 565]. Hand [627, 162, 621]. Handheld [626]. Handle [392]. Handling [306, 288, 318, 317]. Handover [505]. Handovers [504]. Hands [634]. Hanging [731]. Hansel [683]. Hap [232]. Hap-seq [232]. Haplotype [232]. Happen [411]. Haptic [616, 598, 599, 601, 618, 619, 605, 606, 608, 622, 624, 627, 628, 630, 634, 621, 595, 597]. Haptics [596, 604, 606, 625, 982, 983]. Hard [738, 341]. Hardness [814, 719]. Hardware [752, 757, 137, 536, 207, 180, 962]. HAS-160 [163]. Hash [752]. Hashing [325]. Haskell [142, 129, 130]. Having [184]. healing [148]. Health [984, 651, 638, 36, 952, 637, 646, 35]. Healthcare [34, 30]. Healthy [654]. Heart [37]. Hebbian [909]. Held [991, 964, 969]. Hellenic [996]. Hellman [179]. Help [32, 219, 857]. Heraklion [994]. Herbrand [263]. Here [267]. Hereditary [112]. Heterogeneous [789, 83, 785, 778, 483, 16, 790, 592, 23, 776, 670]. Heterozygosity [808]. HetsNets [991]. Heuristic [480, 927, 825]. Heuristics [471, 714]. Hidden [308]. Hiding [173]. Hierarchial [324]. Hierarchical [67, 710, 47]. Hierarchy [673, 191]. High [922, 129, 241]. High-Level [129]. High-Throughput [241]. Higher [828]. Higher-Order [828]. Highly [284]. History [781, 821]. Hit [570]. HIV [242]. Hoc [423, 501, 503, 441, 500].

Home [33, 24, 639, 36, 28]. Homes [952]. Homicide [464]. Homomorphisms [859]. Homothets [116]. Hong [995]. Honour [965]. hop [503, 884]. Hopping [570, 565]. HotWave [380]. Household [645]. HTTP [765]. HTTP-Live [765]. Huge [471]. Human [521, 56, 904, 87, 291, 198, 61, 941]. Human-Agent [87]. HVC [962]. Hybrid [210, 302, 351, 922, 410, 937, 845, 712, 500]. Hybridization [350]. Hybrids [924]. Hyper [555]. Hyper-bent [555]. Hyperelliptic [555]. Hypergraphs [116]. Hyperintensions [200].

I-SEARCH [584]. I/O [870]. IAT [955]. IB [179]. IB-KEM [179]. ICAISC [967, 968, 969]. ICDMS [285]. ICFCA [978]. ICISC [960]. ICOST [952]. identifiable [819]. Identification [548, 818, 6, 357]. Identified [597]. Identify [219]. Identifying [814, 941]. Identity [168]. Identity-Based [168]. Idle [358]. IDS [543]. IEEE [506, 286]. IFIP [991, 989, 990, 973, 972, 979]. IFL [958]. II [990, 561, 983, 968]. IIS [210]. IIS-Guided [210]. Illicit [933]. Illusion [624, 644]. Image [44, 249, 52, 102, 60]. Images [47, 313, 61]. Immune [450]. Impact [24, 751, 700, 630, 952]. Impacts [796]. Impaired [596, 28]. Impatient [738]. Impedance [612]. Implantation [254]. Implementation [471, 450, 142, 382, 91, 180, 874, 837, 958]. Implementations [172]. Implementing [686, 129]. Implications [693, 448]. Implicative [212]. Importance [197]. Improve [29, 364, 501, 494, 884]. Improved [125, 801, 176, 177, 161]. Improvement [598]. Improving [836, 319]. Imputation [232]. In-Home [639, 28]. Inapproximability [807]. Inapproximable [674]. Incidence [812]. Incidental [307]. Incidents [944]. Incomplete [527]. Incompressibility [675]. Incorporating [801]. Increased [228, 758]. Increasing [787]. Incremental [329]. Independent [24, 913, 638, 326, 922, 190, 669]. Index [470, 564, 341]. Index-Based [470]. India [651]. Indian [294]. Indicating [632]. Indiscernibility [327, 102]. Individual [266, 373]. Indoor [498]. Induced [112]. Induction [364]. Industrial [454]. Industry [361]. Inference [226, 186, 705]. Inferring [825]. Infinite [115, 561]. Influence [936, 644, 649, 776]. Infomobility [60]. Infon [266]. Informatics [998, 957, 149]. Information [595, 746, 744, 693, 11, 2, 901, 309, 291, 865, 747, 6, 743, 574, 3, 16, 461, 742, 694, 986, 985, 859, 932, 155, 906, 695, 935, 960]. Information-Centric [744, 743, 574, 742]. Informative [336]. Informing [701]. Infrared [52]. Infrastructure [583, 881, 939, 530, 979]. Infrastructures [393, 591, 590]. Initiative [442, 284]. Injecting [223]. Innovation [856, 680, 205, 25]. Innovations [24]. Innovative [302, 36]. Inprocessing [204]. Inputs [620]. Instance [329, 926, 907, 795]. Instance-Based [795]. Instance-Specific [926]. Instances [675]. Instant [541]. Institutionally [93]. Institutionally-Situated [93]. Institutions [955, 90]. Instrumentation [127]. Insurance [776]. Integer [808, 123, 924, 894]. Integers [838]. Integral [161]. Integrated [580, 735]. Integrating [698]. Integration [44, 674, 35, 198, 688, 997]. Intelligence [346, 309, 902, 703, 528, 996, 970, 959, 967, 968, 953, 998]. Intelligent [285, 581, 277, 464, 294]. Intensity [623, 626, 649, 641]. Intensive [222, 401]. Inter [797, 89, 743, 683, 772, 791]. Inter-Domain [743, 791, 797, 772]. Inter-Organizational [683, 89]. Interacting [73]. Interaction [159, 269, 781, 703, 45]. Interaction-Based [781]. Interactions [583, 416, 56, 86]. Interactive [638, 581, 631, 70, 402]. Interconnected [490]. Interconnection [794]. Interdependent [94, 190]. Interdomain [771]. Interface [400, 421, 627]. Interfaces [606, 378]. Interleaved [450]. International [954, 984, 991, 997, 989, 990, 992, 963, 955, 978, 952, 962, 973, 958, 972, 980, 982, 983, 960, 976, 977, 988, 995, 964, 149, 986, 967, 968, 969, 979, 953, 975, 993, 974, 985, 966]. Internet [981, 977, 583, 580, 57, 581, 773, 582, 792, 589, 703, 791, 577, 578, 579, 796, 405, 643, 586, 573]. Interoperable [972]. Interpolation [842, 218]. Interpolation-Based [218]. Interrelations [588]. Interval [911, 918, 730]. Intervention [652, 654]. Interventions [964, 255]. Intrinsic [622]. Introduced [601]. Introduction [951, 573]. Intrusion [939, 540, 542, 938]. intrusive [40, 38]. Intuitionistic [154]. Intuitive [607]. Invasive [248, 39]. Inverse [264, 660]. Investigating [617]. Investigation [154]. Involvement [25]. IoT [295]. IP [789, 110, 758]. IPTV [494]. IPv6 [789, 548]. Irreducible [562]. isAI [959]. ISBRA [992]. Isolation [563]. ISP [780]. Israel [962]. Issue [452]. Issues [544]. Italy [952, 988, 953]. Item [8]. Items [665]. Itemset [512]. Itemsets [157]. Iterated [358]. Iteratees [841]. Iteration [78]. Iteration-Wise [78]. Iterative [820]. iThink [903]. Itoh [461]. Itoh-attack [461]. IX [971].

JADE [368]. Japan [959, 993, 772]. Japanese [145, 152, 144]. Java [380]. Jobs [670]. Joint [973, 985]. JSAI [959]. JSAI-isAI [959]. Judge [153]. Judgments [151, 620]. Junction [238]. June [984, 997, 952, 973, 972, 980, 982, 983, 976, 977, 988, 979, 974]. Juris [149]. Juris-Informatics [149]. JURISIN [959, 149]. Just [623, 130]. Just-In-Time [130]. Justifying [92].

Kakeya [106]. Kal [724]. Kal-Toh [724]. KEM [179]. Kernel [305, 306, 304, 310, 899, 545]. Kernels [676, 863]. Key [167, 455, 795]. Keyframes [57]. Keys [184, 166]. Keywords [946]. Kinesthetic [620]. Kinetic [611]. Kisses [726]. Kissing [726]. Knapsack [111, 716]. Knowledge [223, 416, 951, 1, 20, 150, 155, 692]. Knowledge-Based [416]. Known [176]. Kobe [993]. Koestler [2]. KoFAS [284]. Kong [995]. Korea [960]. Kronecker [119]. KS [958]. Kuala [998].

Labeled [156]. Labeling [58]. Lack [619]. Lambda [264, 840]. Lamia [996]. Language [831, 140, 326, 270, 11, 922, 845, 136]. Languages [276, 958, 132, 845, 974]. Large [811, 433, 484, 851, 549, 489, 155, 197, 688]. Large-Scale [489, 155]. Lassos [414]. Last [308]. Latency [541]. Latent [170]. Latin [957, 957]. Lattice [945]. Lattices [516, 518, 717, 115, 527]. Lawrence [958]. Layer [301, 308, 936, 884]. Lazy [133, 131]. LBlock [457]. LD [848]. Leakage [452, 461, 172, 165, 463]. Leakages [447]. Lean [734]. Learning [370, 301, 502, 510, 533, 142, 310, 311, 865, 851, 334, 312, 336, 905, 793, 79, 648, 864, 862, 909, 43, 319, 340, 46, 899, 906, 372]. Least [311]. LED [456]. Legal [151, 152]. Length [226, 619, 625, 620]. Lengths [566, 717, 567]. LENLS [959, 140]. Lesamnta [161]. Less [744, 493]. Lessons [577]. Leuven [978]. Level [517, 922, 295, 129, 219, 547]. Leveraging [766]. Lexicon [45]. Libraries [892]. Library [903]. Life [739]. Lifespan [768]. Lifschitz [965, 262, 261]. Lifting [51]. Light [251]. Lightweight [79]. Like [189, 169]. Limited [347, 666]. Linda [441]. Line [52, 314, 620, 670]. Line-Based [52]. Linear [210, 568, 812, 816, 334, 820, 177, 894, 357, 787, 569, 567]. Linearisation [314]. Lines [81]. Linguistic [867]. Link [506, 864, 862]. Linkage [226]. Linked [856, 860, 855, 857, 858]. Linköping [984]. Links [5]. Lithuania [966]. Little [857]. Live [769, 764, 765]. Liveness [213]. Liver [249]. Living [40, 24, 638, 22]. Load [442, 535, 603]. Loading [888]. lobSTR [231]. Local [746, 917, 816, 468, 340, 891, 688, 885]. Localisation [498]. Locality [390, 767]. Locality-Awareness [390]. Localization [801, 895, 800, 499, 258]. Location [502, 917, 32, 640, 239]. Logic [831, 264, 266, 267, 834, 434, 140, 268, 193, 273, 274, 200, 965, 279, 191, 840, 9, 76, 846, 993, 147, 136, 278]. Logic-Based [965]. Logical [516, 134, 705]. Logics [267]. Logit [722]. Long [803, 244]. Long-Range [244]. Longest [929]. Look [459]. Lookaheads [211]. Looking [518]. Loop [478, 235, 648]. Loop-Free [478]. Looping [839]. Loops [429]. LoopWeaver [235]. Loss [808, 234]. Losses [601]. Loudness [633]. Low [107, 638, 570, 623, 879, 387, 571, 541, 552]. Low-Cost [638]. Low-Hit-Zone [570]. Low-Intensity [623]. Low-Latency [541]. Low-Memory [387]. Lower [717]. LTE [504]. Lumpur [998]. Luxembourg [979]. Lyon [955].

Machine [915, 502, 701, 510, 464, 865, 409, 315]. Machines [187, 83, 880, 700, 666, 134, 334, 336, 192]. Macrocells [505]. Macroscopic [524]. Mad [727]. Maintenance [479]. Makes [692]. Makespan [672]. Making [724, 739, 706]. Malaysia [998]. Mammalian [244]. Management [285, 34, 952, 886, 27, 756, 697, 979, 985, 750, 887, 30]. Managing [406]. MANET [502, 90]. MANETs [397, 396, 556]. Manipulation [616, 922, 605]. Manufacturing [735]. Map [118]. Mappings [860]. Maps [909, 319]. Market [776]. Marketing [890]. Marketplace [582]. Markov [226, 252]. Markovian [54, 53]. Marshalling [728]. MAS [363]. Mass [823]. Masses [850]. Massively [345]. Matching [667, 866, 482, 485]. Matchings [480, 740]. Material [610]. Mathematics [103]. Matrices [812]. Matrix [771]. Matter [62, 84, 98, 104, 138, 202, 259, 297, 376, 385, 407, 425, 445, 465, 550, 593, 804, 949, 21, 42, 426, 446, 466, 487, 508, 529, 551, 572, 594, 615, 636, 657, 678, 699, 720, 741, 762, 783, 805, 826, 847, 868, 889, 910, 931, 950, 63, 85, 99, 105, 126, 139, 160, 181, 203, 224, 245, 260, 281, 298, 299, 320, 321, 342, 343, 362, 365, 377, 386, 408]. Maturity [685]. Maude [69]. Maximum [913, 112, 667, 480, 844, 662]. Maximum-Sum [844]. MaxSat [121]. May [987, 411, 991, 997, 989, 990, 992, 998, 955, 978, 995, 996, 986, 967, 968, 969, 975, 993, 994, 985]. May-Happen-in-Parallel [411]. McErlang [424]. MDASCA [463]. MDE [77]. Meaning [366]. Measure [817]. Measurement [613, 538, 287, 147]. Measurement-Theoretic [147]. Measurements [536, 547]. Measures [860, 216]. Measuring [323, 773]. Mechanical [612, 626]. Mechanism [384, 785, 237]. Mechanisms [548]. Mechanoreceptor [624]. Mechatronic [438]. Mediated [600]. Medical [34]. Meets [576]. Members [947]. Memetic [364]. Memory [387, 129, 873, 430]. Mental [606]. Mergers [14]. Mergesort [481]. Mesh [388, 799]. Message [396]. Messages [602, 689]. MEST [34]. Meta [324]. Meta-algorithm [324]. Metadata [389]. Metanorm [96]. Metaoptimization [356]. Metastability [722]. Metastatic [249]. Metering [939]. Method [187, 302, 876, 624, 236, 311, 312, 494, 926, 41, 900, 319, 164, 695, 165]. Methods [856, 20, 787, 944]. Methylation [228]. Metric [477, 374, 319, 823]. Metrics [769, 152, 763]. MICCAI [964]. Microarray [100]. Microblog [936]. Mid [621]. Middleware [875, 398, 405, 399, 887]. Migrating [412]. Migration [348]. MiMI [959, 159]. Mini [340]. Mini-models [340]. Minimal [184]. Minimally [248]. MiniMax [354]. Minimization [760, 719, 662]. Minimizing [107]. Minimum [663, 672, 660]. Mining [940, 509, 510, 939, 483, 330, 775, 938, 157, 895, 337, 8]. MIP [924]. MIP-Based [924]. MIPS [873]. Misjudgment [621]. Misperception [625]. Mispredictions [481, 734]. Miss [120]. Missing [835, 302, 16]. Mixed [808, 247, 924]. Mixture [250, 905]. Mobile [789, 435, 881, 640, 721, 883, 50, 852, 27, 398, 788, 439, 872, 787, 690, 441, 296]. Mobility [583, 982, 983, 75, 28]. Modal [68, 38]. Model [66, 69, 348, 808, 246, 179, 452, 364, 650, 652, 922, 792, 250, 624, 55, 54, 613, 151, 100, 75, 410, 196, 432, 414, 134, 704, 199, 96, 926, 654, 867, 81, 861, 218, 784, 90, 685, 441, 653, 463, 768]. Modeling [599, 193, 510, 235, 252, 237, 684, 255, 198, 402, 375]. Modelling [417, 94, 961, 182, 780, 82]. Models [797, 987, 226, 326, 813, 401, 314, 70, 709, 796, 974, 195, 369, 340]. modems [723]. Moderate [641]. Moderate-Intensity [641]. Modern [402]. Modification [265]. Modification-Based [265]. Modifications [470, 477]. Modifying [597]. Modular [564, 462]. Modulating [610]. Module [800, 180]. MOFETA [296]. Molecular [963]. Monadic [132, 146]. mong [918]. Monitoring [40, 583, 651, 639, 36, 94, 732, 393, 395, 41, 287, 39, 874]. Monolithic [209]. Monoplane [256]. Monte [352]. Morphing [23]. Morphogenesis [701]. Mosaic [658]. Motif [806]. Motifs [233]. Motion [596, 250, 355]. Motion-Impaired [596]. Motivational [646]. Motor [598]. Movement [635]. Movements [621]. Mover [114]. Moving [59]. MPLS [488]. Multi [782, 420, 301, 473, 730, 374, 274, 778, 101, 522, 368, 936, 373, 86, 503, 38, 371, 39, 945, 375, 884, 665]. Multi-Agent [374, 368, 373, 375, 86]. Multi-band [473]. Multi-classifier [371]. Multi-Committer [945]. Multi-context [274]. Multi-dimensional [522]. Multi-domain [782]. Multi-Execution [420]. Multi-hop [503, 884]. Multi-interval [730]. Multi-Layer [936]. Multi-modal [38]. Multi-parametric [39]. Multi-Party [778]. Multi-type [665]. Multi-valued [101, 522]. Multiaccess [535]. Multicast [497]. Multicolorings [668]. Multicriteria [930]. Multilayer [308]. Multilinear [475]. Multimedia [45, 953, 402]. Multimodal [584]. Multimodality [159]. Multiobjective [330]. Multiparty [416]. Multipath [499]. Multiple [429, 477, 229, 482, 876, 942, 825, 415, 899, 670, 507, 361]. Multiple-Access [507]. Multiplexing [876]. Multiprocessor [671]. Multiscale [715]. Multisequences [567]. Multispace [159]. Multithreaded [215, 120]. Multivariate [674]. MURPAR [825]. MUSCLE [953]. Music [349]. Musubi [852]. Mutation [356, 361]. Mutual [901, 501, 842, 871]. Mutualistic [86]. myFitnessCompanion(R) [27].

Name [743, 942]. Nantes [997]. Narrative [587]. Narrative-Aware [587]. Nash [354]. Natural [140]. Navigation [596, 347]. Navigational [607]. Near [915, 101, 102, 629]. Near-Optimal [915]. Near-Threshold [629]. Need [832, 652, 449, 585]. Negabent [558]. Negotiation [366, 367, 793]. Netflow [545]. Nets4Cars [966]. Nets4Cars/Nets4Trains [966]. Nets4Trains [966]. Network [754, 300, 544, 4, 809, 801, 908, 751, 760, 860, 310, 311, 637, 313, 742, 936, 821, 97, 756, 488, 242, 296, 708, 489, 318, 317, 390, 289, 10, 14, 750, 799, 158, 172, 795]. Network-Aware [756]. Network-Based [300, 242]. NETWORKING [991, 989, 990, 282, 574, 404, 991, 989, 990]. Networks [495, 789, 496, 285, 748, 746, 710, 394, 417, 744, 809, 303, 785, 479, 423, 732, 11, 305, 306, 304, 483, 307, 388, 230, 491, 745, 501, 506, 283, 74, 410, 6, 743, 803, 9, 707, 895, 3, 16, 292, 538, 882, 753, 395, 788, 499, 315, 504, 825, 709, 440, 316, 536, 169, 535, 286, 787, 979, 790, 784, 284, 592, 758, 497, 799, 786, 776, 500, 802, 884, 13]. Neural [370, 300, 303, 305, 306, 304, 652, 307, 908, 310, 311, 895, 313, 314, 315, 316, 318, 317, 893, 172]. Neuro [302]. Neuro-wavelet [302]. Neuromorphic [624]. Neurosurgery [247]. Next [789, 875, 592]. NFC [448]. NNI [816]. No [358]. No-Idle [358]. Node [423, 440, 17, 14]. Nodes [15]. Noise [306, 318, 317, 854]. Nominal [145, 335]. Nomograms [307]. Non [40, 263, 667, 305, 330, 310, 51, 819, 926, 41, 38, 823, 39, 175]. Non-contact [41]. Non-crossing [667]. Non-delegatable [175]. Non-Herbrand [263]. Non-identifiable [819]. Non-intrusive [40, 38]. Non-metric [823]. Non-Model-Based [926]. Non-redundant [330]. Non-separable [51]. Non-stationary [305, 310]. Nonexistence [553]. Noninterference [418]. Nonlinear [909]. Nonstationary [318]. Norm [92, 96, 91, 565]. Norm-Aware [91]. Normal [833, 559, 189]. Normative [93, 91, 649].

Norms [955, 97]. Norway [954]. Noticeable [623]. Notification [533]. Notions [167]. Novel [595, 639, 624, 7, 861, 422]. November [960]. NOVI [591]. NP [472, 814]. NP-Complete [472]. NP-Hardness [814]. NPI [143]. NPI/FCI [143]. NT [888]. NTRU [717]. Number [435, 458, 173, 459, 666, 677]. Numbers [333]. Numerical [157]. NV [986].

O [870]. Object [193, 604, 613, 626, 75, 70, 896]. Object-Oriented [70, 896, 193]. Objective [763]. Objects [411, 59, 313]. Oblique [339]. Obscure [774]. Observation [413, 61]. Observational [31]. Obtained [166]. Occasion [961]. October [958]. Odd [552]. Off [546, 787]. Offline [643]. Offloading [505]. Offloadings [883]. Old [404]. Older [25]. Oldest [404]. Omniscience [96]. On-Board [283]. On-Demand [764]. On-Line [314, 670]. ON-OFF [787]. One [725, 452]. One-Size-Fits-All [725]. Online [679, 746, 637, 672, 937, 369, 665]. Ontologies [849]. Ontology [322, 531, 269, 44, 866, 878, 43]. Ontology-Driven [322, 531]. Open [856, 249]. OpenCL [221]. Opera [402, 89]. Operating [382]. Operating-System [382]. Operationalization [95]. Operator [356]. Operators [833]. OPNET [290]. Opportunistic [396, 501]. Opportunities [282]. Optical [592]. Optimal [915, 732, 232, 658, 766, 571]. Optimising [136]. Optimization [473, 363, 350, 921, 908, 353, 332, 355, 357, 930, 890]. Optimized [207]. Optimizing [823]. Optimzation [997]. Orchestrated [72]. Orchestration [439]. Order [264, 434, 191, 310, 422, 828]. Ordered [327]. Ordering [913]. Ordinal [523]. ORG [339]. ORG- [339]. Organisational [698]. Organising [319]. Organization [234, 942, 369]. Organizational [683, 89]. Organizations [955, 88]. organized [543]. Orientation [251]. Orientations [729]. Oriented [380, 187, 83, 519, 193, 971, 70, 152, 878, 896, 569]. Orthogonal [305, 306, 304]. Oscillator [458]. Oscillators [230]. Oslo [954]. OSM [193]. OSM-Logic [193]. OSPF [759]. Overcoming [96]. Overhead [409]. Overhearing [501]. Overlay [785, 249, 780, 497]. Overlay-ISP [780]. Overlays [485]. Oversubscription [796]. Overview [4, 20, 19, 713, 10]. Ownership [600].

P [818]. P-Binder [818]. P2P [764, 169, 763]. P=NP [674]. PA [450]. PA-DPL [450]. Pachycondyla [344]. Pacific [998]. Package [384]. Packet [601]. Packing [673, 676]. Page [323]. Pages [336]. Pairing [946]. Pairs [108, 563]. PAISI [998]. Papers [954, 955, 962, 958, 960, 964, 959, 953]. PARADE [345]. Parallel [411, 435, 345, 301, 348, 666, 355, 843, 129, 844, 192, 670]. Parallelism [895]. Parameter [300, 728, 229, 612, 677]. Parameterized [265, 121, 675]. Parameters [570, 792]. parametric [39]. Parasite [169]. Parcellation [258]. Parentally [822]. Parentally-Biased [822]. Pareto [664]. Paris [737]. Parity [729]. Parsers [151]. Parsimonious [808, 825]. Part [989, 990, 982, 983, 967, 968]. Partial [431, 883, 461, 422]. Partial-Order [422]. Partially [327]. Participate [643]. Participatory [291]. Particle [350, 355, 144]. Particles [355]. Partition [872]. Partitional [359]. Partitioning [227, 469]. Party [778]. Parzen [310, 318, 317]. Parzen-Type [310, 318, 317]. Passenger [655]. Passive [41, 603]. Patches [599]. Path [654, 929]. Paths [478, 761]. Patients [651, 639, 39, 258]. Patrolling [662]. Pattern [442, 877, 38, 635]. Patterns [470, 71, 438, 607, 392, 766, 217, 9, 64, 82]. Pay [794]. PCA [300]. PCI [803]. pCloud [870]. Pearl [834]. Pedigrees [819]. Peep [518]. Peer [679, 769, 440, 390, 765, 768]. Peer-to-Peer [440, 390, 765, 768]. Peering [773]. People [33, 638, 295, 25, 641]. Peptide [812]. Peptide-Protein [812]. Perceived [650, 652, 626]. Perception [616, 619, 612, 982, 983, 617, 854]. Perceptron [301, 308, 312]. Perceptual [601, 102]. Perfect [566, 118, 740, 553, 552]. Performance [69, 836, 491, 501, 319, 379, 402, 497, 884]. Performances [629]. Periodic [492, 552]. periodicity [169]. Peripheral [634]. Permutation [758]. Permutations [658, 554]. Persistent [834]. Person [34, 604]. Person-Centric [34]. Personal [231]. Personalized [30]. Perspective [934, 810]. Perspectives [688]. Persuade [655, 641]. Persuasion [652, 30, 648, 653]. PERSUASIVE [984, 642, 651, 639, 645, 644, 656, 647, 649, 653, 984]. Persuasiveness [650]. Perturbation [240]. Peru [957]. Pervasive [576, 995, 399, 887]. Phasing [232]. Phenomenon [709]. Pheromone [927]. Pheromone-Based [927]. Phone [641]. Phones [879]. Photometric [302]. Phylogenetic [825]. Physical [751, 641]. Pico [490]. Pico-Datacenter [490]. Picture [731]. Picture-Hanging [731]. Piecewise [820]. Piezo [599]. Pilot [633]. Pioneering [205]. Pipeline [822]. Pisa [953]. PKZIP [176]. Placement [766]. Plagiarism [325, 943]. Plaintext [176]. Plan [374]. Planar [724, 917, 116]. Plane [113, 757, 659, 621]. Planning [903, 588, 280, 803, 255, 91]. Plant [357]. Platform [428, 39, 404]. Platforms [800]. Plausible [54]. Play [142]. Playback [764]. Players [142]. Point [223]. Pointcut [378]. Points [108, 518, 484, 663, 460]. Poland [967, 968, 969]. Policies [388, 764]. Policy [352, 283, 886]. Polygonal [662]. Polynomial [435, 474, 718, 568, 674, 191]. Polynomial-Time [191]. Polynomials [555, 333]. Pool [355]. POPE [822]. Population [29, 337, 361]. Portfolio [926]. Post [29]. Post-stroke [29]. Potassium [36]. Power [760, 882, 451, 644, 460, 164, 165]. Power-Constrained [882]. Practical [519, 213, 91, 25, 172]. Practice [472, 253, 770, 986]. Practices [590]. Prague [991, 989, 990]. Pre [328, 533]. Pre-Congestion [533]. Pre-processing [328]. Precision [619]. Predicate [147]. Predict [769]. Predicting [214]. Prediction [370, 502, 865, 242, 315, 787]. Predictive [314, 654]. Preemptions [666]. Preferences [485]. Preliminary [29]. Premature [843]. Preoperative [246]. Preprocessing [204]. Prerequisite [154]. Prerequisite-Effect [154]. Presence [454, 499]. Preservation [274]. Preserving [940, 859]. Prevention [24, 952, 539]. Preventive [646]. Price [467]. Pricing [664, 786, 665]. Primal [266, 475]. Principal [753]. Principles [295, 683, 686, 708, 578]. Priorities [416, 573]. Priority [144]. Privacy [940, 274]. Privacy-Preserving [940]. Private [166]. Probabilistic [417, 143, 306, 311, 217, 895, 15, 318, 317]. Probability [539]. Probe [251]. Problem [106, 435, 782, 471, 110, 726, 472, 913, 915, 111, 808, 122, 733, 716, 660, 894, 929, 719, 46, 928]. Problems [671, 997, 112, 474, 917, 676, 921, 468, 344, 927, 928, 844, 930, 369]. Proceedings [989, 990, 982, 983, 967, 968, 987, 984, 991, 997, 992, 998, 963, 978, 952, 957, 973, 972, 980, 976, 977, 988, 995, 996, 986, 969, 979, 975, 993, 994, 974, 985, 966]. Process [95, 366, 419, 938, 682, 537, 153, 685]. Processes [412, 434, 813, 751, 413, 427, 513, 849]. Processing [748, 155, 858, 328]. Processor [873]. Product [81]. Production [275, 896]. Profiled [165]. Profiler [231]. Profiling [866]. Program [420, 29, 714]. Programming [380, 263, 997, 267, 268, 474, 808, 827, 123, 279, 924, 925, 864, 894, 845, 846, 993, 829]. Programs [215, 271, 273, 424, 734, 127, 415, 131, 422, 828]. Progress [216]. Project [31]. Projection [249]. Prometheus [490]. Promises [981]. Proneness [219]. Proof [77]. Proofs [452]. Propagation [812, 498, 920, 936, 891, 172]. Propagators [923]. Properties [323, 622, 134]. Property [173, 610, 843, 135]. Property-Based [135]. Prophecy [711]. Proposal [374, 792]. Proprioception [609]. Proprioceptive [617]. Protection [391, 788]. Protein [225, 812, 818]. Protein-Protein [818]. Proteins [235, 243]. Protocol [496, 367, 947]. Prototype [293]. Prototyping [875]. Provide [29]. Provided [525]. Provider [774]. Provider-Free [774]. Proxy [168]. Pruning [13]. PS [838]. Pseudo [595, 597]. Pseudo-haptic [595, 597]. Pseudorandom [560]. Pseudorandomness [119]. Public [167, 870, 854]. Public-Key [167]. Publication [520]. Publish [397, 390]. Publish/Subscribe [397, 390]. Published [449]. Publishing [770]. Pursuit [710]. Puzzle [735]. Puzzles [731, 777]. PXI [605].

QoE [763]. QoS [406]. Quadratic [901, 557]. Quality [763, 794, 799, 887]. Quantification [236]. Quantifier [714]. Quantifier-Elimination [714]. Quantifying [230]. Quantitative [217, 795]. Quantum [706]. Quartet [811]. Quasi [169]. Quasi-periodicity [169]. Quaternions [566]. Queries [124]. Query [496, 11, 892, 194, 845, 858]. Query-Based [496]. Querying [269]. Questions [141]. Queues [834]. Quotients [568].

R [705]. R-Calculus [705]. RAB [653]. Rabbit [177]. Racing [908]. Rademacher [333]. Radio [503, 786]. Railway [294]. Random [458, 717, 119, 125, 480, 459, 777, 185, 58, 164]. Range [124, 244]. RASP [271]. Rate [370, 37]. Ratio [110, 539]. Rationale [31]. Rationalized [735]. ray [256]. RBF [315]. RDF [863, 183, 867, 859]. RDF-Based [867]. Re [168]. Re-encryption [168]. Reachability [210]. Reaching [366, 634].

Reactivity [209]. Readiness [652]. Real [544, 834, 36, 476, 438, 542, 74, 448, 355, 60, 713, 361]. Real-Time [544, 834, 36, 438, 542, 74, 355, 713]. Real-World [476, 448]. Realisation [301]. Realised [313]. Realistic [247]. Reality [247, 249, 630, 981]. Realizability [443, 73]. Realization [472]. Really [761]. Rearrangements [241]. Reasoning [93, 965, 865, 9, 150]. Recall [152]. Recall-Oriented [152]. Reciprocities [588]. Recognition [32, 50, 879, 942, 337, 38, 635, 169, 53, 941]. RECOMB [963]. Recombination [682]. Recommendation [856]. Recommendations [372]. Reconfiguration [388]. Reconfigurations [71]. Reconsidering [924]. Reconstruction [302, 809, 821]. Record [35, 545]. Recordings [521]. Recovery [250]. Rectangular [410]. Recurrent [301]. Recursive [306, 311, 318, 436]. Rediscovering [618]. Reduced [619, 814, 163]. Reducing [781, 81]. Reduction [225, 422, 361]. Redundancy [760]. redundant [330]. Reflections [190]. Reflective [648]. Region [900, 662, 933]. Region-Building [900]. Registers [569]. Registration [246, 52, 254, 630]. Regression [364, 305, 304, 310, 820, 315, 340]. Regularity [904]. Regulatory [809]. Rehabilitation [28]. Reimagining [404]. Reinforcement [533, 906]. Relation [865]. Relational [184]. Relations [327, 143, 562]. Relationship [600]. Relaxations [474, 475]. Relay [491]. Release [167]. Relevance [144]. Reliability [402]. Reliable [480, 497, 500]. Reliant [391]. Remarks [184, 675]. Remedy [756]. Remote [39]. Renaming [117, 118]. Rényi [109, 730, 186]. Reoptimization [112]. Reoptimizing [477]. Repair [537, 206]. Repeat [231]. Replicated [389]. Report [467]. Representation [658, 150, 907]. Representations [486]. Representative [15]. Republic [991, 989, 990]. Reputation [781, 793]. Reputation-Aware [793]. Request [876]. Rerouting [772]. Research [992, 618, 693, 681, 684, 687, 579, 25, 986, 575, 994, 573, 695, 22, 963]. Resilience [751, 750]. Resilient [754, 542, 393]. Resistance [178]. Resisting [450]. Resolution [416, 743, 879, 507]. Resource [75, 886, 784, 590, 870, 874]. Resource-Restricted [75]. Resources [332]. Response [689]. Restricted [75]. Restriction [661, 191]. Restrictions [334]. Resulting [610]. Results [473, 29, 31]. Retrieval [584, 50, 861, 907]. Retrieving [336]. Reveal [511]. Revenue [782, 870]. Reverse [220]. Reversed [614]. Reversible [409, 136]. Review [12]. Revised [954, 955, 962, 958, 960, 964, 959, 953]. Revision [272]. Revisited [830, 921, 167, 492]. Revisiting [120]. Revocable [173]. Revocation [288, 947]. Revoked [173]. Rewriting [133]. Rich [57, 713]. Ride [928]. Right [162]. Rigs [628]. Ring [458]. Risk [66]. RNA [470, 236, 238, 244]. RNA-seq [238, 236]. RNBB [500]. Road [479]. Roadmap [573]. Robinson [817]. Robot [33, 531, 257]. Robot-Environment [33]. Robotic [468]. Robotics [346, 276]. Robots [347, 721]. Robust [394, 473, 236, 53]. Robustness [758]. Rod [619]. Rogue [814]. Rooted [817, 156]. Rotating [732]. Rough [102, 338, 956, 514]. Round [163]. Round-Reduced [163].

Route [291, 928, 791]. Routing [496, 478, 927, 759, 758, 369]. Row [641]. RPL [537]. RSA [453, 454, 171, 455]. Rule [363, 775, 100, 150, 896]. Rule-Based [100, 150]. Rulers [564]. Rules [269, 275, 330, 339]. Run [93, 79]. Run-Time [93, 79]. Runtime [580, 71].

s [536]. SaaS [876, 948]. SaaS-Driven [948]. Safe [65]. Safety [80, 33, 984, 655, 589, 213]. Sagittal [621]. Saliency [61]. SAM [215]. Same [460]. Sample [334]. Sanctioning [95]. Santorini [977]. SAR [47]. SAT [435, 204, 211, 675]. Satisfaction [827]. Satisfiability [212, 677]. Satisfy [738]. Scalability [514, 389]. Scalable [895, 757]. Scale [433, 484, 851, 549, 489, 48, 155]. Scaling [235, 755]. Scan [454]. Scandinavian [725]. Scenario [856, 448]. Scenarios [59, 75, 290, 503]. SCGMM [59]. Scheduled [556]. Scheduling [107, 671, 915, 401, 421, 924, 666, 292, 672, 492, 928, 670]. Schemas [199, 197]. Scheme [495, 789, 831, 114, 173, 945, 884]. Schemes [673, 184, 171, 787, 947]. Schindler [461]. Science [693, 681, 684, 687, 986, 695, 706]. Sciences [702]. Scientific [705]. Scoped [433]. Screen [599]. SEA [976]. Seamless [400]. Search [470, 584, 917, 352, 248, 344, 239, 925, 823, 930, 891, 861, 214, 584]. Searchable [946]. Searches [816]. Searching [906]. Second [191, 155]. Second-Order [191]. Secondary [801]. Secret [947]. Secure [420, 940, 179, 444, 778, 777, 168, 975, 761]. Securing [939]. Security [998, 452, 581, 459, 392, 938, 283, 167, 960, 530, 979, 284, 180, 172, 776]. Segment [844]. Segmentation [59, 751]. Segments [58]. Selected [954, 955, 962, 958, 960, 964, 959, 953]. Selecting [5]. Selection [866, 329, 308, 333, 48, 313]. Selective [51, 148]. Self [543, 148, 215, 391, 442, 313, 238, 30, 319, 28]. Self-adaptive [215]. Self-Correcting [313]. Self-efficacy [28]. Self-healing [148]. Self-Initiative [442]. Self-management [30]. Self-Organising [319]. Self-organized [543]. Self-Reliant [391]. Self-selective [148]. Self-training [238]. Semantic [832, 57, 521, 191, 526, 892, 336, 994, 848]. Semantically [585]. Semantics [140, 273, 855, 419, 146, 885]. Semi [57]. Semi-automatic [57]. Semidefinite [225]. Semiosis [373]. Sense [618, 894]. Sensing [639, 37, 291, 503]. Sensing-Before-Transmit [503]. Sensitive [749, 201, 897]. Sensor [496, 110, 732, 882, 395, 499, 39, 497]. Sensors [543, 732]. Sentence [325]. Sentences [152]. Sentiment [749]. Seoul [960]. separable [51]. Separation [48]. September [954, 964]. seq [236, 232, 238]. Sequence [470, 119, 570, 904, 563, 239]. Sequence-Structure [470]. Sequences [566, 109, 915, 133, 568, 561, 556, 571, 553, 560, 552, 565, 980]. Sequencing [232, 241]. Sequent [832]. Sequential [539]. Sequentially [494]. Serene [391]. Serializability [214]. Series [305, 306, 304, 315, 545]. Series-Type [305, 306, 304]. Serious [655]. Serum [36]. Server [873]. Service [40, 396, 83, 519, 406, 582, 72, 696, 684, 395, 878, 405, 794, 23]. Service-Oriented [83, 519, 878]. Services [754, 24, 73, 755, 877, 979, 389, 585]. Session [391]. Sessions [601]. Set [263, 913, 118, 327, 273, 101, 8]. SETA [980]. SETN [996]. Sets [570, 811, 838, 102, 824, 956, 514, 552]. Shape [56, 49, 50, 854]. Shapes [597]. Shared [696, 767]. Sharing [782, 747, 932, 906, 947]. Shift [569, 631]. Shoes [608]. Shop [915]. Short [749, 946, 231, 315]. Shortest [717, 478]. Shots [57]. SIDE [969, 509, 391, 450, 447, 162, 975, 463]. Side-Channel [447, 975, 463]. Sieving [123]. Sign [761]. Signal [498, 239]. Signaling [789]. Signature [173, 171, 175]. Signcryption [174]. Significance [820]. Silhouettes [49]. Similarities [512, 17]. Similarity [323, 100, 823, 8]. Simple [265, 329, 367, 928, 929]. Simplex [312]. Simplex-Method [312]. Simplification [13]. Simulating [275]. Simulation [346, 227, 908]. Simulations [290]. Simultaneous [55, 212, 947]. Single [736]. Single-Digit [736]. Singletons [392]. SIP [755]. Sites [818]. Situated [93]. Situation [399, 935]. Situation-Awareness [399]. Size [725, 844]. Size-Constrained [844]. Skein [462]. Skeletons [843]. Skill [252]. Skin [609]. SLA [793]. Slead [387]. Sleep [41, 38]. Slicing [387]. Sliding [482]. Slums [651]. Small [216, 410, 709]. Smallest [725]. Smart [588, 24, 952, 800, 879, 587, 690, 23, 449]. Smartphone [631, 641]. Smartphones [583]. Smooth [370]. SmoothCache [765]. SMT [414, 214]. SMT-Based [414, 214]. Snake [640]. SOA [877]. Social [583, 748, 746, 710, 576, 693, 747, 637, 707, 852, 427, 30, 428, 404, 649]. Societies [95, 576]. Socio [403]. Socio-technical [403]. Soft [967, 968]. Software [954, 187, 888, 962, 381, 201, 971, 656, 428, 164, 590]. Software-Based [428]. Solution [819]. Solutions [24, 36, 581, 952, 282, 688, 284]. Solvability [830]. Solve [435]. Solver [838, 922]. Solvers [922, 211, 925, 857]. Solving [912, 836, 122, 736, 929, 358]. SOM [893]. SOM-Based [893]. Some [184, 112, 675]. Something [404]. Sophisticated [752]. Sort [902]. Sorting [734, 902]. Soundness [413]. Sourcing [696]. Southeast [933]. Space [470, 114, 442, 409, 441]. Space-Based [442]. Space-Efficient [114]. Space-Time [441]. Spain [963]. Spam [539]. Spanners [113]. SPARQL [855]. Sparse [812, 480]. Special [460]. Specialize [798]. Specific [326, 926]. Specification [68, 217]. Specifications [864]. Specifics [324]. Spectra [823, 316, 394]. Speed [482, 635]. Speeding [303]. Spike [316]. Spike-Flow [316]. Spin [546]. Spin-Off [546]. Spinal [255]. Spiral [205]. SPIT [539]. Splice [238]. Splicing [244]. Splitting [265, 188]. Spreading [17, 552]. Springs [609]. SPRT [539]. SQL [835, 189]. SQL-Like [189]. Squares [311]. SSH [540]. SSHCure [540]. Stability [515, 809, 722]. Stability-Based [809]. Stable [378]. Stack [133]. Standard [179]. Standardization [35, 575]. Standards [579]. State [187, 83, 134, 756, 192, 278]. Statecharts [368]. Statecharts-Based [368]. Stateful [755]. Statelets [427]. Statements [897]. States [392]. Static [420, 611, 81]. stationary [305, 310]. Statistical [904, 186, 820, 771]. Statistics [310, 61]. Status [467]. Steady [387, 756]. Stealthy [775]. Steps [435]. Stereo [313]. Stereo-Images [313]. Stimulation [624, 614, 633, 617]. Stimulus [622]. Stockholm [973, 972, 974]. Storage [874]. Strategic [722]. Strategies [652, 803, 468, 359, 361]. Strategy [344, 209, 449]. Stream [328, 939, 451, 58]. Streaming [495, 888, 769, 764, 765, 534, 763]. Streams [330, 331, 332]. Strengthened [477]. Stretch [609]. Strictly [659]. String [482]. stroke [29]. Stroll [400]. Strong [271, 305, 306, 484, 167, 318, 175]. Structator [470]. Structural [650, 233, 711]. Structure [470, 225, 746, 327, 521, 250, 450, 151, 51, 243, 154, 153]. Structure-Guided [243]. Structured [251]. Structures [244]. Study [401, 751, 537, 150, 27, 337, 633, 656, 31, 402, 933, 802]. Styles [69]. Sub [118]. Sub-consensus [118]. Subgraph [112, 18]. Subgraphs [724, 122]. Subject [187]. Subject-Oriented [187]. Sublinear [673]. Subnetwork [483]. Subobject [430]. Subproblems [736]. Subscribe [397, 390]. Subscriptions [397]. Succinct [486, 659]. Sudoku [736]. Suitability [692]. Sum [660, 844]. Sum-Type [660]. Summaries [904, 218]. SUMO [290]. Supertrees [816]. Support [34, 24, 406, 650, 336, 315, 153, 691]. Supporting [608, 534, 586]. Supportive [638]. Supra [629]. Supra-Threshold [629]. Surface [251]. Surfers [738]. Surgery [248, 257]. Surveillance [286]. Survey [196]. Surveys [302]. Swarm [346, 531, 350, 352, 902, 355, 969]. Swarms [768]. Sweden [984, 973, 972, 974]. Switches [230]. SWRL [896]. SWRL2COOL [896]. Symbol [904]. Symbolic [221, 413, 418]. Symposia [969]. Symposium [954, 992, 957, 958, 976, 993]. Synchronization [429, 433]. Synchronous [421]. Syntactic [151]. Synthesis [67, 208, 209]. Synthetic [239, 177]. System [285, 40, 679, 300, 836, 955, 888, 650, 838, 939, 818, 540, 943, 291, 937, 705, 382, 64, 219, 287, 286, 153, 257, 689, 874]. Systematic [575, 449]. Systems [322, 67, 397, 443, 914, 193, 812, 374, 433, 274, 693, 438, 368, 973, 781, 972, 622, 432, 294, 494, 683, 373, 686, 694, 387, 877, 713, 986, 541, 65, 207, 389, 691, 375, 653, 695, 22, 372].

Tabled [837]. Tables [752, 564]. Tactile [602, 607, 614]. Tag [448]. Tailoring [645]. Taipei [955]. Taiwan [955]. Takamatsu [959]. Talk [643]. TAMC [987]. Taming [64]. Tampere [982, 983]. Tandem [231, 823]. Tapping [610]. Targets [621]. Task [400, 866, 28, 768]. Task-Based [768]. Tasks [107, 118, 909, 43]. Taxa [814]. Taylor [911]. TC [991, 989, 990]. TCP [884]. Team [697]. Teamwork [87]. Technical [695, 403]. Technique [684, 422]. Techniques [997, 204, 973, 255, 929, 81, 12]. Technological [934, 707, 205]. Technologies [24, 282, 528, 646, 966]. Technology [642, 984, 651, 639, 303, 656, 647, 649]. Teenagers [642]. Tele [28]. Tele-Rehabilitation [28]. Telemanipulation [601]. Telematics [952]. Telerehabilitation [29]. TELIX [867]. Temperature [671]. Temperature-Aware [671]. Template [345, 384, 682]. Temporal [71, 928]. Tenants [876]. Term [7, 803]. Termination [843]. Terms [839]. Terrorism [933]. Test [222, 223, 247, 539, 488, 131]. Test-Bed [488]. Testing [223, 221, 962, 72, 134, 135, 415, 422]. Text [749, 943, 6, 944]. Text-Based [943]. Textual [323]. Textural [48]. Texture [56, 55, 54, 53]. Textures [595]. Thalheim [961]. Their [141, 980, 569, 692]. Theorem [117, 410]. Theoretic [798, 147]. Theoretical [961, 957, 646]. Theories [421, 996]. Theory [68, 848, 272, 273, 182, 101, 702, 525, 986, 893, 379, 691, 987]. Therapeutic [243]. There [267]. Thins [725]. Third [975]. Those [620]. Three [116, 240]. Three-Dimensional [116, 240]. Threshold [629, 945]. Throughput [241]. Time [93, 544, 673, 834, 718, 193, 36, 123, 421, 306, 438, 304, 674, 816, 191, 542, 614, 716, 74, 311, 355, 60, 79, 713, 315, 357, 317, 130, 441, 545]. Time-Dependent [193]. Time-Reversed [614]. Time-Varying [306, 304, 311, 317]. Timed [424, 72, 167, 414]. Timed-Release [167]. Timers [412]. Tissue [250]. Toh [724]. Tolerance [463]. Tolerant [80]. Tomography [715]. Tool [470, 34, 702, 685]. Toolkit [940]. Tools [951, 813, 368, 20, 591]. Top [725, 858]. Top- [858]. Topologies [97]. Topology [821]. Toronto [964]. Torque [598]. Touch [599, 600, 603]. Trace [583, 130]. Trace-Based [130]. Traces [133, 164]. Trackers [767]. Tracking [355, 61]. Tractable [728, 158]. Tradeoffs [614, 780]. Traffic [538, 788, 771, 759, 767]. Train [293, 728, 295]. Train-to-Wayside [293]. Trainers [252]. Training [303, 257, 238]. TrAins [296]. Trajectory [314]. Trajectory-Based [314]. Trans [748]. Trans-Social [748]. Transaction [157]. Transactional [128, 430]. Transactions [971, 970, 956]. Transcatheter [254]. Transcript [810, 236]. Transcription [809]. Transcripts [824]. TransDPOR [422]. Transfer [604]. Transfers [928]. Transform [51, 557]. Transformation [420, 69, 589, 196, 896, 569]. Transformations [70]. Transforming [199]. Transition [548]. Transitions [534]. Translating [264]. Translation [152]. Transmit [503]. Transmitting [602]. Transparency [598]. Transportation [292, 294]. Trapdoor [946]. Tree [921, 486, 669, 158, 897]. Tree-Width [921]. Trees [115, 811, 364, 816, 817, 156, 825]. Trial [153]. Triconnected [724]. Trie [806]. Trie-based [806]. Triggered [421]. Trivium [162]. True [458]. TrueSight [238]. Trustworthy [582, 393]. TSP [477]. Tuning [895]. Turing [701, 700]. Turning [640]. Turns [663]. Tussle [574]. TV [58]. Tweaked [161]. Twitter [857, 851, 935]. Two [915, 124, 46, 662]. Two-Class [46]. Two-Dimensional [124]. Two-Machine [915]. TX [992]. Type [305, 306, 304, 310, 660, 318, 317, 665]. Typed [264, 829]. Types [228, 444].

Ubiquitous [403, 694]. Ulam [730]. Ultrasound [251]. UMBEL [861]. Unbalanced [51]. Unbounded [566]. Uncertainty [473]. Understand [585]. Understanding [521, 953]. undertaking [37]. Unidirectional [168]. Unified [584, 691]. Uniform [304, 554]. Unity [832]. Universal [400, 940, 125]. Unless [674]. Unmanned [295]. Unordered [156, 158]. Unrooted [816, 817]. Unsupervised [47, 862, 899]. Update [190]. Upper [125, 677]. Urban [588, 587]. USA [992, 958, 986]. Use [519, 29, 229, 643].

User [645, 742, 27, 586, 22]. User-Assisted [742]. User-Centred [22]. Users [596, 173, 494, 643]. Using [583, 596, 597, 598, 746, 502, 363, 326, 246, 866, 274, 732, 424, 275, 939, 769, 860, 232, 311, 295, 747, 539, 627, 714, 47, 51, 432, 413, 537, 312, 355, 684, 238, 255, 877, 488, 241, 30, 460, 632, 864, 371, 81, 861, 162, 547, 790, 859, 136, 53, 545]. Utilitarian [95]. Utilizing [628].

Vagueness [103]. Validation [374, 893, 695]. Validity [341]. Value [830, 687, 853, 795]. valued [101, 522]. Values [460]. Valve [254]. VANET [290]. VANETs [288]. Vapnik [333]. Variable [226, 913, 556, 358]. Variable-Length [226]. Variables [266, 834, 711]. Variants [348]. Variation [228]. Variations [444, 158]. Various [348, 524, 800]. Varying [306, 304, 311, 317]. Vector [717, 336, 315]. Vegas [986]. Vehicle [927, 286, 369]. Vehicles [292, 966]. Vehicular [285, 289, 180]. Venice [988]. Verifiable [174]. Verification [875, 423, 71, 424, 962, 205, 713, 135, 207, 828]. Verified [235]. Verifier [175]. Verifying [80]. versus [134, 649]. Vertex [471, 710]. Vertex-Pursuit [710]. Vertical [504]. VHOs [790]. via [226, 188, 418, 906, 767]. Vibrating [599]. Vibration [610]. Vibrations [629]. Vibrotactile [602, 623, 626, 629, 633]. Video [903, 246, 764, 766, 534, 763]. Video-Games [903]. View [194, 885]. Viewing [766]. Viewpoint [213]. Views [835, 190]. VII [955, 970]. Vilnius [966]. Violations [214]. Viral [890]. Virtual [754, 880, 888, 605, 255, 757, 488, 697]. Virtualization [489, 873]. Virtualized [591]. Visibility [347]. Visible [52]. Vision [609, 251, 625]. Visit [651]. Visits [315]. Visual [323, 600, 581, 619, 661, 248, 630, 43]. Visual-Haptic [630]. Visualisation [248]. Visualization [938, 253, 255, 802]. Vladimir [965, 262, 261]. VM [756]. VNEMX [488]. VoIP [539]. Volume [253]. Voluntariness [647]. Voronoi [661]. VR [257]. vs [213, 214]. Vulcan [724]. Vulnerability [752, 170].

Waitfreedom [840]. Walk [941]. Walking [640]. Walsh [557]. Waterloo [980]. Wave [196]. Wavelength [563]. Wavelet [51, 302]. Wayside [293]. Weak [337, 317]. Weaker [118]. Weaknesses [171]. Wearable [39]. Web [509, 850, 994, 869, 323, 848, 509, 406, 738, 855, 526, 336, 852, 877, 861]. Web/LD [848]. webinos [398]. Website [315]. Weight [112, 559, 477, 556, 617]. Weighted [476, 235, 157, 660]. Weights [334]. WG [973, 972, 979]. Where [576]. Which [798]. whilst [37]. Who [142]. Whole [627]. Whole-Hand [627]. WI [955]. WI-IAT [955]. Wide [688]. Widespread [244]. Width [921, 675]. Width-Parameterized [675]. WiFi [498]. Wiki [692]. Wikis [692]. Willingness [794]. Wind [632]. Windows [888, 482]. Wired [977]. Wired/Wireless [977]. Wireless [496, 417, 491, 501, 395, 788, 499, 571, 286, 784, 497, 799, 884, 977]. Wirelessly [490]. Wise [78]. within [532]. without [453, 347, 209]. Word [894, 569]. Word-Oriented [569]. Words [371]. Work [692]. Workflow [35]. Working [909]. Workload [606, 207]. Workshop [998, 964, 149, 953, 975, 155, 966]. Workshops [991, 955, 959]. World [476, 448, 490, 361]. Worst [677]. Worst-Case [677]. Wrapper [329]. Wrong [835]. WUENIC [150]. WWIC [977]. WWW [326].

X [256]. X-ray [256]. XACML [76]. Xen [488]. XV [956].

Yield [620]. yo [144]. Youth [262].

Zakopane [967, 968, 969]. Zapping [494]. Zero [170, 552]. ZigZag [405]. Zone [570, 552]. Zonotopes [718].

References

Berthold:2012:TBK

[1] Michael R. Berthold. Towards bisociative knowledge discovery. Lecture Notes in Computer Science, 7250:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-31830-6_1.

Dubitzky:2012:TCI

[2] Werner Dubitzky, Tobias Kötter, Oliver Schmidt, and Michael R. Berthold. Towards creative information exploration based on Koestler’s concept of bisociation. Lecture Notes in Computer Science, 7250:11– 32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/978-
3-642-31830-6_2.

Kotter:2012:INB

[3] Tobias Kötter and Michael R. Berthold. From information networks to bisociative information networks. Lecture Notes in Computer Science, 7250:33–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/content/pdf/10.1007/
978-3-642-31830-6_3.

Borgelt:2012:NCO

[4] Christian Borgelt. Network creation: Overview. Lecture Notes in Computer Science, 7250:51–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-31830-6_4.

Segond:2012:SLB

[5] Marc Segond and Christian Borgelt. Selecting the links in BisoNets generated from document collections. Lecture Notes in Computer Science, 7250:54–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/content/pdf/10.1007/
978-3-642-31830-6_5.

Jursic:2012:BCI

[6] Matjaž Juršič, Borut Sluban, Bojan Cestnik, Miha Grčar, and Nada Lavrač. Bridging concept identification for constructing information networks from text documents. Lecture Notes in Computer Science, 7250: 66–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/978-
3-642-31830-6_6.

Hynonen:2012:DNT

[7] Teemu Hynönen, Sébastien Mahler, and Hannu Toivonen. Discovery of novel term associations in a document collection. Lecture Notes in Computer Science, 7250:91–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-31830-6_7.

Segond:2012:CSB

[8] Marc Segond and Christian Borgelt. Cover similarity based item set mining. Lecture Notes in Computer Science, 7250:104–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/content/pdf/10.1007/978-3-642-31830-6_8.

Kimmig:2012:PLR

[9] Angelika Kimmig, Esther Galbrun, Hannu Toivonen, and Luc De Raedt. Patterns and logic for reasoning with networks. Lecture Notes in Computer Science, 7250:122–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/content/pdf/
10.1007/978-3-642-31830-6_9.

Toivonen:2012:NAO

[10] Hannu Toivonen. Network analysis: Overview. Lecture Notes in Computer Science, 7250:144–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/content/pdf/
10.1007/978-3-642-31830-6_10.

Dries:2012:BQL

[11] Anton Dries, Siegfried Nijssen, and Luc De Raedt. BiQL: a query language for analyzing information networks. Lecture Notes in Computer Science, 7250:147–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-31830-6_11.

Zhou:2012:RBA

[12] Fang Zhou, Sébastien Mahler, and Hannu Toivonen. Review of BisoNet abstraction techniques. Lecture Notes in Computer Science, 7250:166– 178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/978-
3-642-31830-6_12.

Zhou:2012:SNE

[13] Fang Zhou, Sébastien Mahler, and Hannu Toivonen. Simplification of networks by edge pruning. Lecture Notes in Computer Science, 7250: 179–198, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/content/pdf/10.1007/
978-3-642-31830-6_13.

Toivonen:2012:NCN

[14] Hannu Toivonen, Fang Zhou, Aleksi Hartikainen, and Atte Hinkka. Network compression by node and edge mergers. Lecture Notes in Computer Science, 7250:199–217, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/content/pdf/
10.1007/978-3-642-31830-6_14.

Langohr:2012:FRN

[15] Laura Langohr and Hannu Toivonen. Finding representative nodes in probabilistic graphs. Lecture Notes in Computer Science, 7250:218–229, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/978-
3-642-31830-6_15.

Kotter:2012:MCD

[16] Tobias Kötter and Michael R. Berthold. (missing) concept discovery in heterogeneous information networks. Lecture Notes in Computer Science, 7250:230–245, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-31830-6_16.

Thiel:2012:NSS

[17] Kilian Thiel and Michael R. Berthold. Node similarities from spreading activation. Lecture Notes in Computer Science, 7250:246–262, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/978-
3-642-31830-6_17.

Nagel:2012:TDS

[18] Uwe Nagel, Kilian Thiel, Tobias Kötter, Dawid Piatek, and Michael R. Berthold. Towards discovery of subgraph bisociations. Lecture Notes in Computer Science, 7250:263–284, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-31830-6_18.

Nurnberger:2012:EO

[19] Andreas Nürnberger. Exploration: Overview. Lecture Notes in Computer Science, 7250:285–286, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/content/pdf/
10.1007/978-3-642-31830-6_19.

Gossen:2012:DEB

[20] Tatiana Gossen, Marcus Nitsche, Stefan Haun, and Andreas Nürnberger. Data exploration for bisociative knowledge discovery: a brief overview of tools and evaluation methods. Lecture Notes in Computer Science, 7250: 287–300, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/content/pdf/10.1007/
978-3-642-31830-6_20.

Anonymous:2012:FMa

[21] Anonymous. Front matter. Lecture Notes in Computer Science, 7250: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31830-6/1.

Woolrych:2012:CUC

[22] Ryan Woolrych and Andrew Sixsmith. Challenges of user-centred research in the development of ambient assisted living systems. Lecture Notes in Computer Science, 7251:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30779-9_1/.

Yang:2012:FSM

[23] Hen-I Yang, Ryan Babbitt, Johnny Wong, and Carl K. Chang. A framework for service morphing and heterogeneous service discovery in smart environments. Lecture Notes in Computer Science, 7251:9–17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30779-9_2/.

Barrientos:2012:IAA

[24] James Barrientos, Jeffrey Soar, and Ying Su. Impact analysis of assessment, consultation and education services to support the adoption of smart home technologies, innovations for chronic disease prevention and solutions for independent living. Lecture Notes in Computer Science, 7251:18–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30779-9_3/.

Parent:2012:EFF

[25] Anne-Sophie Parent, Nena Georgantzi, and Ilenia Gheno. Ensuring a fruitful future to innovation and research: Practical guidance for the involvement of older people in research. Lecture Notes in Computer Science, 7251:26–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30779-9_4/.

Bonfiglio:2012:FCC

[26] Silvio Bonfiglio. Fostering a continuum of care. Lecture Notes in Computer Science, 7251:35–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30779-9_5/.

Leijdekkers:2012:UAM

[27] Peter Leijdekkers and Valerie Gay. User adoption of mobile apps for chronic disease management: a case study based on myFitnessCompanion(R). Lecture Notes in Computer Science, 7251:42–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30779-9_6/.

Sanford:2012:EHT

[28] Jon A. Sanford, Patricia C. Griffiths, and Helen Hoenig. Effects of in-home tele-rehabilitation on task self-efficacy in mobility impaired adults. Lecture Notes in Computer Science, 7251:50–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30779-9_7/.

Corriveau:2012:UTP

[29] Hélène Corriveau, Michel Tousignant, Sylvie Gosselin, and Patrick Boissy. The use of telerehabilitation to provide an exercise program to improve balance in a post-stroke population: Preliminary results. Lecture Notes in Computer Science, 7251:58–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30779-9_8/.

Mukhtar:2012:PHS

[30] Hamid Mukhtar, Arshad Ali, Sungyoung Lee, and Djamel Belaïd. Personalized healthcare self-management using social persuasion. Lecture Notes in Computer Science, 7251:66–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30779-9_9/.

Puddu:2012:COS

[31] Paolo Emilio Puddu, John M. Morgan, Concetta Torromeo, Nick Curzen, and Michele Schiariti. A clinical observational study in the CHIRON project: Rationale and expected results. Lecture Notes in Computer Science, 7251:74–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30779-9_10/.

Chen:2012:DLH

[32] Chao Chen, Daqing Zhang, Lin Sun, Mossaab Hariz, and Yang Yuan. Does location help daily activity recognition? Lecture Notes in Computer Science, 7251:83–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30779-9_11/.

Andriatrimoson:2012:REC

[33] Antonio Andriatrimoson, Nadia Abchiche, Simon Galerne, and Etienne Colle. A robot-environment cooperation architecture for the safety of elderly people at home. Lecture Notes in Computer Science, 7251: 91–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30779-9_12/.

Barca:2012:MES

[34] Carlos Cavero Barca, Juan Mario Rodríguez, Rosana Valle Soriano, and Alberto Rugnone. Medical expert support tool (MEST): a person-centric approach for healthcare management. Lecture Notes in Computer Science, 7251:99–106, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30779-9_13/.

Gazzarata:2012:IHC

[35] Roberta Gazzarata, Fabio Vergari, Jan-Marc Verlinden, Francesco Morandi, and Simone Naso. The integration of e-health into the clinical workflow — electronic health record and standardization efforts. Lecture Notes in Computer Science, 7251:107–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30779-9_14/.

Corsi:2012:ISH

[36] Cristiana Corsi, Johan De Bie, David Mortara, and Stefano Severi. Innovative solutions in health monitoring at home: The real-time assessment of serum potassium concentration from ECG. Lecture Notes in Computer Science, 7251:116–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30779-9_15/.

Cleland:2012:ACF

[37] Ian Cleland, Chris Nugent, Dewar Finlay, William Burns, and Jennifer Bougourd. Assessment of custom fitted heart rate sensing garments whilst undertaking everyday activities. Lecture Notes in Computer Science, 7251: 124–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30779-9_16/.

Ni:2012:MMN

[38] Hongbo Ni, Bessam Abdulrazak, Daqing Zhang, Shu Wu, Xingshe Zhou, and Kejian Miao. Multi-modal non-intrusive sleep pattern recognition in elder assistive environment. Lecture Notes in Computer Science, 7251: 132–139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30779-9_17/.

Solar:2012:NIW

[39] Héctor Solar, Erik Fernández, Gennaro Tartarisco, Giovanni Pioggia, and Božidara Cvetković. A non invasive, wearable sensor platform for multi-parametric remote monitoring in CHF patients. Lecture Notes in Computer Science, 7251:140–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30779-9_18/.

Allegre:2012:NIM

[40] Willy Allègre, Thomas Burger, Pascal Berruet, and Jean-Yves Antoine. A non-intrusive monitoring system for ambient assisted living service delivery. Lecture Notes in Computer Science, 7251:148–156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30779-9_19/.

McDowell:2012:PSA

[41] Andrew McDowell, Mark Donnelly, Chris Nugent, and Michael McGrath. Passive sleep actigraphy: Evaluating a non-contact method of monitoring sleep. Lecture Notes in Computer Science, 7251:157–164, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30779-9_20/.

Anonymous:2012:FMb

[42] Anonymous. Front matter. Lecture Notes in Computer Science, 7251: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30779-9/1.

Perner:2012:LOV

[43] Petra Perner. Learning an ontology for visual tasks. Lecture Notes in Computer Science, 7252:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32436-9_1/.

Colantonio:2012:OAI

[44] Sara Colantonio and Massimo Martinelli. Ontology and algorithms integration for image analysis. Lecture Notes in Computer Science, 7252: 17–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32436-9_2/.

Malandrakis:2012:EAL

[45] Nikos Malandrakis and Alexandros Potamianos. EmotiWord: Affective lexicon creation with application to interaction and multimedia data. Lecture Notes in Computer Science, 7252:30–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32436-9_3/.

Ruiz:2012:BAL

[46] Pablo Ruiz, Javier Mateos, and Rafael Molina. A Bayesian active learning framework for a two-class classification problem. Lecture Notes in Computer Science, 7252:42–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32436-9_4/.

Kayabol:2012:UCS

[47] Koray Kayabol and Vladimir A. Krylov. Unsupervised classification of SAR images using hierarchical agglomeration and EM. Lecture Notes in Computer Science, 7252:54–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32436-9_5/.

Sziranyi:2012:GTC

[48] Tamás Szirányi and Dániel Szolgay. Geometrical and textural component separation with adaptive scale selection. Lecture Notes in Computer Science, 7252:66–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32436-9_6/.

Kim:2012:BSS

[49] Donghoon Kim and Rozenn Dahyot. Bayesian shape from silhouettes. Lecture Notes in Computer Science, 7252:78–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32436-9_7/.

Kovacs:2012:SRR

[50] Levente Kovács. Shape retrieval and recognition on mobile devices. Lecture Notes in Computer Science, 7252:90–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32436-9_8/.

Keskin:2012:DSF

[51] Furkan Keskin and A. Enis Çetin. Directionally selective fractional wavelet transform using a 2-D non-separable unbalanced lifting structure. Lecture Notes in Computer Science, 7252:102–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32436-9_9/.

Han:2012:VII

[52] Jungong Han, Eric Pauwels, and Paul de Zeeuw. Visible and infrared image registration employing line-based geometric analysis. Lecture Notes in Computer Science, 7252:114–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32436-9_10/.

Vacha:2012:TRU

[53] Pavel Vácha and Michal Haindl. Texture recognition using robust Markovian features. Lecture Notes in Computer Science, 7252:126–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32436-9_11/.

Haindl:2012:PTE

[54] Michal Haindl and Vojtěch Havlíček. A plausible texture enlargement and editing compound Markovian model. Lecture Notes in Computer Science, 7252:138–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32436-9_12/.

Haindl:2012:BTF

[55] Michal Haindl and Michal Havlíček. Bidirectional texture function simultaneous autoregressive model. Lecture Notes in Computer Science, 7252: 149–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32436-9_13/.

Filip:2012:AHG

[56] Jiří Filip, Pavel Vácha, and Michal Haindl. Analysis of human gaze interactions with texture and shape. Lecture Notes in Computer Science, 7252: 160–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32436-9_14/.

Carcel:2012:RIA

[57] Elisabet Carcel and Manuel Martos. Rich Internet application for semiautomatic annotation of semantic shots on keyframes. Lecture Notes in Computer Science, 7252:172–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32436-9_15/.

Martienne:2012:LTS

[58] Emmanuelle Martienne and Vincent Claveau. Labeling TV stream segments with conditional random fields. Lecture Notes in Computer Science, 7252:183–194, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32436-9_16/.

Gallego:2012:FOS

[59] Jaime Gallego, Montse Pardàs, and Montse Solano. Foreground objects segmentation for moving camera scenarios based on SCGMM. Lecture Notes in Computer Science, 7252:195–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32436-9_17/.

Magrini:2012:RTI

[60] Massimo Magrini and Davide Moroni. Real time image analysis for infomobility. Lecture Notes in Computer Science, 7252:207–218, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32436-9_18/.

Szalai:2012:TSF

[61] Szilard Szalai and Tamás Szirányi. Tracking the saliency features in images based on human observation statistics. Lecture Notes in Computer Science, 7252:219–233, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32436-9_19/.

Anonymous:2012:BMa

[62] Anonymous. Back matter. Lecture Notes in Computer Science, 7252: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-32436-9/1.

Anonymous:2012:FMc

[63] Anonymous. Front matter. Lecture Notes in Computer Science, 7252: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32436-9/1.

Meseguer:2012:TDS

[64] José Meseguer. Taming distributed system complexity through formal patterns. Lecture Notes in Computer Science, 7253:1–2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-35743-5_1.

Rushby:2012:CSS

[65] John Rushby. Composing safe systems. Lecture Notes in Computer Science, 7253:3–11, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-35743-5_2/.

Braendeland:2012:DMC

[66] Gyrd Brændeland, Atle Refsdal, and Ketil Stølen. A denotational model for component-based risk analysis. Lecture Notes in Computer Science, 7253:12–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35743-5_3/.

Aminof:2012:SHS

[67] Benjamin Aminof, Fabio Mogavero, and Aniello Murano. Synthesis of hierarchical systems. Lecture Notes in Computer Science, 7253:42–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35743-5_4/.

Bauer:2012:MST

[68] Sebastian S. Bauer and Kim Guldstrand Larsen. A modal specification theory for components with data. Lecture Notes in Computer Science, 7253:61–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35743-5_5/.

Bruni:2012:EPM

[69] Roberto Bruni and Alberto Lluch Lafuente. Evaluating the performance of model transformation styles in Maude. Lecture Notes in Computer Science, 7253:79–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35743-5_6/.

Li:2012:ITO

[70] Dan Li, Xiaoshan Li, Zhiming Liu, and Volker Stolz. Interactive transformations from object-oriented models to component-based models. Lecture Notes in Computer Science, 7253:97–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35743-5_7/.

Dormoy:2012:RVT

[71] Julien Dormoy, Olga Kouchnarenko, and Arnaud Lanoix. Runtime verification of temporal patterns for dynamic reconfigurations of components. Lecture Notes in Computer Science, 7253:115–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35743-5_8/.

Escobedo:2012:TCT

[72] Jose Pablo Escobedo and Christophe Gaston. Timed conformance testing for orchestrated service discovery. Lecture Notes in Computer Science, 7253:133–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-35743-5_9/.

Gossler:2012:RCS

[73] Gregor Gössler and Gwen Salaün. Realizability of choreographies for services interacting asynchronously. Lecture Notes in Computer Science, 7253:151–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-35743-5_10/.

Jaghoori:2012:NRT

[74] Mohammad Mahdi Jaghoori and Ólafur Hlynsson. Networks of real-time actors. Lecture Notes in Computer Science, 7253:168–186, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35743-5_11/.

Johnsen:2012:FMO

[75] Einar Broch Johnsen and Rudolf Schlatte. A formal model of object mobility in resource-restricted deployment scenarios. Lecture Notes in Computer Science, 7253:187–204, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35743-5_12/.

Ramli:2012:LX

[76] Carroline Dewi Puspa Kencana Ramli. The logic of XACML. Lecture Notes in Computer Science, 7253:205–222, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35743-5_13/.

Kezadri:2012:PAB

[77] Mounira Kezadri, Benoît Combemale, and Marc Pantel. A proof assistant based formalization of MDE components. Lecture Notes in Computer Science, 7253:223–240, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35743-5_14/.

Limet:2012:CIW

[78] Sébastien Limet, Sophie Robert, and Ahmed Turki. Controlling an iteration-wise coherence in dataflow. Lecture Notes in Computer Science, 7253:241–258, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35743-5_15/.

Martin:2012:LFL

[79] José Antonio Martín and Antonio Brogi. Learning from failures: a lightweight approach to run-time behavioural adaptation. Lecture Notes in Computer Science, 7253:259–277, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35743-5_16/.

Ameur-Boulifa:2012:VSF

[80] Rabéa Ameur-Boulifa, Raluca Halalai, and Ludovic Henrio. Verifying safety of fault-tolerant distributed components. Lecture Notes in Computer Science, 7253:278–295, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35743-5_17/.

Sabouri:2012:RMC

[81] Hamideh Sabouri and Ramtin Khosravi. Reducing the model checking cost of product lines using static analysis techniques. Lecture Notes in Computer Science, 7253:296–312, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35743-5_18/.

Sanchez:2012:BMA

[82] Alejandro Sanchez and Luís Soares. Barbosa. Bigraphical modelling of architectural patterns. Lecture Notes in Computer Science, 7253: 313–330, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35743-5_19/.

Brugali:2012:CEH

[83] Davide Brugali, Luca Gherardi, and Elvinia Riccobene. Coordinated execution of heterogeneous service-oriented components by abstract state machines. Lecture Notes in Computer Science, 7253:331–349, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35743-5_20/.

Anonymous:2012:BMb

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-35743-5/1.

Anonymous:2012:FMd

[85] Anonymous. Front matter. Lecture Notes in Computer Science, 7253: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-35743-5/1.

Lurgi:2012:MAC

[86] Miguel Lurgi and David Robertson. Multi-agent coordination through mutualistic interactions. Lecture Notes in Computer Science, 7254:1– 20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35545-5_1/.

Harbers:2012:EHA

[87] Maaike Harbers, Jeffrey M. Bradshaw, Matthew Johnson, and Paul Feltovich. Explanation in human-agent teamwork. Lecture Notes in Computer Science, 7254:21–37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35545-5_2/.

Keogh:2012:ACD

[88] Kathleen Keogh and Liz Sonenberg. Adaptive coordination in distributed and dynamic agent organizations. Lecture Notes in Computer Science, 7254:38–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35545-5_3/.

Jiang:2012:ABI

[89] Jie Jiang, Virginia Dignum, and Yao-Hua Tan. An agent-based interorganizational collaboration framework: OperA+. Lecture Notes in Computer Science, 7254:58–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35545-5_4/.

Tampitsikas:2012:MMF

[90] Charalampos Tampitsikas, Stefano Bromuri, and Michael Ignaz Schumacher. MANET: a model for first-class electronic institutions. Lecture Notes in Computer Science, 7254:75–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35545-5_5/.

Panagiotidi:2012:TPN

[91] Sofia Panagiotidi and Javier Vázquez-Salceda. Towards practical normative agents: a framework and an implementation for norm-aware planning. Lecture Notes in Computer Science, 7254:93–109, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35545-5_6/.

Letia:2012:TJN

[92] Ioan Alfred Letia and Anca Goron. Towards justifying norm compliance. Lecture Notes in Computer Science, 7254:110–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35545-5_7/.

Balke:2012:NRT

[93] Tina Balke, Marina De Vos, and Julian Padget. Normative run-time reasoning for institutionally-situated BDI agents. Lecture Notes in Computer Science, 7254:129–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35545-5_8/.

Cranefield:2012:MMI

[94] Stephen Cranefield, Michael Winikoff, and Wamberto Vasconcelos. Modelling and monitoring interdependent expectations. Lecture Notes in Computer Science, 7254:149–166, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35545-5_9/.

Balke:2012:OSP

[95] Tina Balke and Daniel Villatoro. Operationalization of the sanctioning process in utilitarian artificial societies. Lecture Notes in Computer Science, 7254:167–185, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35545-5_10/.

Mahmoud:2012:OON

[96] Samhar Mahmoud, Nathan Griffiths, Jeroen Keppens, and Michael Luck. Overcoming omniscience for norm emergence in Axelrod’s metanorm model. Lecture Notes in Computer Science, 7254:186–202, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35545-5_11/.

Mahmoud:2012:ENN

[97] Samhar Mahmoud, Nathan Griffiths, Jeroen Keppens, and Michael Luck. Establishing norms for network topologies. Lecture Notes in Computer Science, 7254:203–220, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35545-5_12/.

Anonymous:2012:BMc

[98] Anonymous. Back matter. Lecture Notes in Computer Science, 7254: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-35545-5/1.

Anonymous:2012:FMe

[99] Anonymous. Front matter. Lecture Notes in Computer Science, 7254: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-35545-5/1.

Janusz:2012:DRB

[100] Andrzej Janusz. Dynamic rule-based similarity model for DNA microarray data. Lecture Notes in Computer Science, 7255:1–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31903-7_1/.

El-Monsef:2012:MVA

[101] M. E. Abd El-Monsef, H. M. Abu-Donia, and E. A. Marei. Multi-valued approach to near set theory. Lecture Notes in Computer Science, 7255: 26–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31903-7_2/.

Henry:2012:PIR

[102] Christopher J. Henry. Perceptual indiscernibility, rough sets, descriptively near sets, and image analysis. Lecture Notes in Computer Science, 7255: 41–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31903-7_3/.

Mani:2012:DCM

[103] A. Mani. Dialectics of counting and the mathematics of vagueness. Lecture Notes in Computer Science, 7255:122–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31903-7_4/.

Anonymous:2012:BMd

[104] Anonymous. Back matter. Lecture Notes in Computer Science, 7255: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-31903-7/1.

Anonymous:2012:FMf

[105] Anonymous. Front matter. Lecture Notes in Computer Science, 7255: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31903-7/1.

Ahn:2012:GCK

[106] Hee-Kap Ahn, Sang Won Bae, and Otfried Cheong. A generalization of the convex Kakeya problem. Lecture Notes in Computer Science, 7256:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29344-3_1/.

Angel:2012:LCS

[107] Eric Angel, Evripidis Bampis, and Vincent Chau. Low complexity scheduling algorithm minimizing the energy for tasks with agreeable deadlines. Lecture Notes in Computer Science, 7256:13–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29344-3_2/.

Arkin:2012:BCP

[108] Esther M. Arkin and José Miguel Díaz-Báñez. Bichromatic 2-center of pairs of points. Lecture Notes in Computer Science, 7256:25–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29344-3_3/.

Arvind:2012:ERS

[109] Vikraman Arvind and Partha Mukhopadhyay. Erdős–Rényi sequences and deterministic construction of expanding Cayley graphs. Lecture Notes in Computer Science, 7256:37–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29344-3_4/.

Barbosa:2012:BAR

[110] Rafael da Ponte Barbosa and Yoshiko Wakabayashi. A better approximation ratio and an IP formulation for a sensor cover problem. Lecture Notes in Computer Science, 7256:49–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29344-3_5/.

Bockenhauer:2012:ACK

[111] Hans-Joachim Böckenhauer and Dennis Komm. On the advice complexity of the knapsack problem. Lecture Notes in Computer Science, 7256:61–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29344-3_6/.

Boria:2012:RSM

[112] Nicolas Boria, Jérôme Monnot, and Vangelis Th. Paschos. Reoptimization of some maximum weight induced hereditary subgraph problems. Lecture Notes in Computer Science, 7256:73–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29344-3_7/.

Bose:2012:PCB

[113] Prosenjit Bose, Rolf Fagerberg, and André van Renssen. On plane constrained bounded-degree spanners. Lecture Notes in Computer Science, 7256:85–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29344-3_8/.

Brody:2012:SEA

[114] Joshua Brody, Hongyu Liang, and Xiaoming Sun. Space-efficient approximation scheme for circular Earth mover distance. Lecture Notes in Computer Science, 7256:97–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29344-3_9/.

Busic:2012:DCI

[115] Ana Bušić, Nazim Fatès, and Jean Mairesse. Density classification on infinite lattices and trees. Lecture Notes in Computer Science, 7256: 109–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29344-3_10/.

Cardinal:2012:CPH

[116] Jean Cardinal and Matias Korman. Coloring planar homothets and threedimensional hypergraphs. Lecture Notes in Computer Science, 7256: 121–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29344-3_11/.

Castaneda:2012:ETA

[117] Armando Castañeda, Maurice Herlihy, and Sergio Rajsbaum. An equivariance theorem with applications to renaming. Lecture Notes in Computer Science, 7256:133–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29344-3_12/.

Castaneda:2012:RWT

[118] Armando Castañeda, Damien Imbs, and Sergio Rajsbaum. Renaming is weaker than set agreement but for perfect renaming: a map of sub-consensus tasks. Lecture Notes in Computer Science, 7256:145–156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29344-3_13/.

Cesaratto:2012:PRK

[119] Eda Cesaratto and Brigitte Vallée. Pseudorandomness of a random Kronecker sequence. Lecture Notes in Computer Science, 7256:157–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29344-3_14/.

Cole:2012:RCM

[120] Richard Cole and Vijaya Ramachandran. Revisiting the cache miss analysis of multithreaded algorithms. Lecture Notes in Computer Science, 7256: 172–183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29344-3_15/.

Crowston:2012:PCM

[121] Robert Crowston, Gregory Gutin, and Mark Jones. Parameterized complexity of MaxSat above average. Lecture Notes in Computer Science, 7256:184–194, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29344-3_16/.

Cygan:2012:SDC

[122] Marek Cygan, Marcin Pilipczuk, and Michał Pilipczuk. Solving the 2disjoint connected subgraphs problem faster than 2^{n}. Lecture Notes in Computer Science, 7256:195–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29344-3_17/.

Dadush:2012:TSA

[123] Daniel Dadush. A O(1/ϵ^{2})^{n}-time sieving algorithm for approximate integer programming. Lecture Notes in Computer Science, 7256:207–218, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29344-3_18/.

Davoodi:2012:TDR

[124] Pooya Davoodi, Michiel Smid, and Freek van Walderveen. Twodimensional range diameter queries. Lecture Notes in Computer Science, 7256:219–230, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29344-3_19/.

Dellamonica:2012:IUB

[125] Domingos Dellamonica Jr. and Yoshiharu Kohayakawa. An improved upper bound on the density of universal random graphs. Lecture Notes in Computer Science, 7256:231–242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29344-3_20/.

Anonymous:2012:FMg

[126] Anonymous. Front matter. Lecture Notes in Computer Science, 7256: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29344-3/1.

Middelkoop:2012:FIA

[127] Arie Middelkoop and Alexander B. Elyasov. Functional instrumentation of ActionScript programs with Asil. Lecture Notes in Computer Science, 7257:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34407-7_1/.

Amsden:2012:FTE

[128] Edward Amsden and Matthew Fluet. Fairness for transactional events. Lecture Notes in Computer Science, 7257:17–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-34407-7_2/.

Maier:2012:IHL

[129] Patrick Maier and Phil Trinder. Implementing a high-level distributedmemory parallel Haskell in Haskell. Lecture Notes in Computer Science, 7257:35–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34407-7_3/.

Schilling:2012:CTB

[130] Thomas Schilling. Challenges for a trace-based just-in-time compiler for Haskell. Lecture Notes in Computer Science, 7257:51–68, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34407-7_4/.

Reich:2012:LGC

[131] Jason S. Reich and Matthew Naylor. Lazy generation of canonical test programs. Lecture Notes in Computer Science, 7257:69–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34407-7_5/.

Persson:2012:GMC

[132] Anders Persson and Emil Axelsson. Generic monadic constructs for embedded languages. Lecture Notes in Computer Science, 7257:85–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34407-7_6/.

Chang:2012:STL

[133] Stephen Chang, Eli Barzilay, and John Clements. From stack traces to lazy rewriting sequences. Lecture Notes in Computer Science, 7257: 100–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34407-7_7/.

Koopman:2012:MBT

[134] Pieter Koopman and Peter Achten. Model based testing with logical properties versus state machines. Lecture Notes in Computer Science, 7257:116–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34407-7_8/.

Page:2012:PBT

[135] Rex Page. Property-based testing and verification: a catalog of classroom examples. Lecture Notes in Computer Science, 7257:134–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34407-7_9/.

Thomsen:2012:DOR

[136] Michael Kirkedal Thomsen. Describing and optimising reversible logic using a functional language. Lecture Notes in Computer Science, 7257: 148–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34407-7_10/.

Megacz:2012:HDG

[137] Adam Megacz. Hardware design with generalized arrows. Lecture Notes in Computer Science, 7257:164–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34407-7_11/.

Anonymous:2012:BMe

[138] Anonymous. Back matter. Lecture Notes in Computer Science, 7257: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-34407-7/1.

Anonymous:2012:FMh

[139] Anonymous. Front matter. Lecture Notes in Computer Science, 7257: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34407-7/1.

Butler:2012:LEN

[140] Alastair Butler. Logic and engineering of natural language semantics (LENLS) 8. Lecture Notes in Computer Science, 7258:1–2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-32090-3_1.

Asher:2012:BQT

[141] Nicholas Asher and Jason Quinley. Begging questions, their answers and basic cooperativity. Lecture Notes in Computer Science, 7258:3– 12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32090-3_2/.

Hernandes:2012:PWL

[142] Mauricio Hernandes. Players who are learning how to play: a Haskell implementation of awareness dynamics. Lecture Notes in Computer Science, 7258:13–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32090-3_3/.

Colinet:2012:ENF

[143] Margot Colinet and Grégoire Winterstein. Emphatic NPI/FCI and adversative discourse relations, a probabilistic approach. Lecture Notes in Computer Science, 7258:27–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32090-3_4/.

Oshima:2012:JPY

[144] David Y. Oshima. The Japanese particle yo in declaratives: Relevance, priority, and blaming. Lecture Notes in Computer Science, 7258: 40–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32090-3_5/.

Hayashishita:2012:CNE

[145] J.-R. Hayashishita and Daisuke Bekki. Conjoined nominal expressions in Japanese. Lecture Notes in Computer Science, 7258:54–67, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32090-3_6/.

Unger:2012:DSM

[146] Christina Unger. Dynamic semantics as monadic computation. Lecture Notes in Computer Science, 7258:68–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32090-3_7/.

Suzuki:2012:MTF

[147] Satoru Suzuki. Measurement-theoretic foundations of gradable-predicate logic. Lecture Notes in Computer Science, 7258:82–95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32090-3_8/.

Butler:2012:TSS

[148] Alastair Butler and Kei Yoshimoto. Towards a self-selective and selfhealing evaluation. Lecture Notes in Computer Science, 7258:96–109, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32090-3_9/.

Ota:2012:FIW

[149] Shozo Ota, Ken Satoh, and Makoto Nakamura. The Fifth International Workshop on Juris-Informatics (JURISIN 2011). Lecture Notes in Computer Science, 7258:110–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-32090-3_10.

Kowalski:2012:WCS

[150] Robert Kowalski and Anthony Burton. WUENIC — a case study in rulebased knowledge representation and reasoning. Lecture Notes in Computer Science, 7258:112–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32090-3_11/.

Igari:2012:DSA

[151] Hirokazu Igari, Akira Shimazu, and Koichiro Ochimizu. Document structure analysis with syntactic model and parsers: Application to legal judgments. Lecture Notes in Computer Science, 7258:126–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32090-3_12/.

Ogawa:2012:ROE

[152] Yasuhiro Ogawa, Masaki Mori, and Katsuhiko Toyama. Recall-oriented evaluation metrics for consistent translation of Japanese legal sentences. Lecture Notes in Computer Science, 7258:141–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-32090-3_13/.

Sato:2012:DPS

[153] Takanori Sato, Shogo Okada, and Katsumi Nitta. Deliberation process support system for citizen judge trial based on structure of factors. Lecture Notes in Computer Science, 7258:155–169, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32090-3_14/.

Sano:2012:IIP

[154] Katsuhiko Sano, Shingo Hagiwara, and Satoshi Tojo. An intuitionistic investigation of prerequisite-effect structure. Lecture Notes in Computer Science, 7258:170–183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32090-3_15/.

Tsuda:2012:SWA

[155] Koji Tsuda and Shin ichi Minato. Second workshop on algorithms for large-scale information processing in knowledge discovery (ALSIP). Lecture Notes in Computer Science, 7258:184–185, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/accesspage/chapter/10.1007/978-3-642-32090-3_16.

Higuchi:2012:ACE

[156] Shoichi Higuchi, Tomohiro Kan, Yoshiyuki Yamamoto, and Kouichi Hirata. An A* algorithm for computing edit distance between rooted labeled unordered trees. Lecture Notes in Computer Science, 7258: 186–196, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32090-3_17/.

Kameda:2012:MCW

[157] Yuichi Kameda and Akihiro Yamamoto. Mining closed weighted itemsets for numerical transaction databases. Lecture Notes in Computer Science, 7258:197–210, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32090-3_18/.

Yamamoto:2012:CTV

[158] Yoshiyuki Yamamoto, Kouichi Hirata, and Tetsuji Kuboyama. On computing tractable variations of unordered tree edit distance with network algorithms. Lecture Notes in Computer Science, 7258:211–223, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32090-3_19/.

Bono:2012:MMI

[159] Mayumi Bono and Nobuhiro Furuyama. Multimodality in multispace interaction (MiMI). Lecture Notes in Computer Science, 7258:224–225, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-32090-3_20.

Anonymous:2012:FMi

[160] Anonymous. Front matter. Lecture Notes in Computer Science, 7258: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32090-3/1.

Sasaki:2012:IIA

[161] Yu Sasaki and Kazumaro Aoki. Improved integral analysis on tweaked Lesamnta. Lecture Notes in Computer Science, 7259:1–17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31912-9_1/.

Schilling:2012:ATU

[162] Thorsten Ernst Schilling and Håvard Raddum. Analysis of Trivium using compressed right hand side equations. Lecture Notes in Computer Science, 7259:18–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31912-9_2/.

Mendel:2012:CRR

[163] Florian Mendel, Tomislav Nad, and Martin Schläffer. Cryptanalysis of round-reduced HAS-160. Lecture Notes in Computer Science, 7259: 33–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31912-9_3/.

Strobel:2012:EME

[164] Daehyun Strobel and Christof Paar. An efficient method for eliminating random delays in power traces of embedded software. Lecture Notes in Computer Science, 7259:48–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31912-9_4/.

Zhang:2012:ELC

[165] Hailong Zhang, Yongbin Zhou, and Dengguo Feng. An efficient leakage characterization method for profiled power analysis attacks. Lecture Notes in Computer Science, 7259:61–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31912-9_5/.

Lee:2012:CEP

[166] Hyung Tae Lee, HongTae Kim, Yoo-Jin Baek, and Jung Hee Cheon. Correcting errors in private keys obtained from Cold boot attacks. Lecture Notes in Computer Science, 7259:74–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31912-9_6/.

Kikuchi:2012:SSN

[167] Ryo Kikuchi, Atsushi Fujioka, Yoshiaki Okamoto, and Taiichi Saito. Strong security notions for timed-release public-key encryption revisited. Lecture Notes in Computer Science, 7259:88–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31912-9_7/.

Luo:2012:FSU

[168] Song Luo, Qingni Shen, and Zhong Chen. Fully secure unidirectional identity-based proxy re-encryption. Lecture Notes in Computer Science, 7259:109–126, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31912-9_8/.

Qiao:2012:DPP

[169] Yong Qiao, Yuexiang Yang, Jie He, Bo Liu, and Yingzhi Zeng. Detecting parasite P2P botnet in eMule-like networks through quasiperiodicity recognition. Lecture Notes in Computer Science, 7259:127– 139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31912-9_9/.

Chen:2012:ADL

[170] Kai Chen, Yifeng Lian, and Yingjun Zhang. AutoDunt: Dynamic latent dependence analysis for detection of zero day vulnerability. Lecture Notes in Computer Science, 7259:140–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31912-9_10/.

Kramer:2012:WCR

[171] Juliane Krämer, Dmitry Nedospasov, and Jean-Pierre Seifert. Weaknesses in current RSA signature schemes. Lecture Notes in Computer Science, 7259:155–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31912-9_11/.

Yang:2012:BPN

[172] Shuguo Yang, Yongbin Zhou, Jiye Liu, and Danyang Chen. Back propagation neural network based leakage characterization for practical security analysis of cryptographic implementations. Lecture Notes in Computer Science, 7259:169–185, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31912-9_12/.

Emura:2012:RGS

[173] Keita Emura, Atsuko Miyaji, and Kazumasa Omote. A revocable group signature scheme with the property of hiding the number of revoked users. Lecture Notes in Computer Science, 7259:186–203, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31912-9_13/.

ElAimani:2012:GCV

[174] Laila El Aimani. Generic constructions for verifiable signcryption. Lecture Notes in Computer Science, 7259:204–218, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31912-9_14/.

Tian:2012:NDS

[175] Haibo Tian, Xiaofeng Chen, Zhengtao Jiang, and Yusong Du. Nondelegatable strong designated verifier signature on elliptic curves. Lecture Notes in Computer Science, 7259:219–234, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31912-9_15/.

Jeong:2012:IKP

[176] Kyung Chul Jeong, Dong Hoon Lee, and Daewan Han. An improved known plaintext attack on PKZIP encryption algorithm. Lecture Notes in Computer Science, 7259:235–247, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31912-9_16/.

Lu:2012:SLA

[177] Yi Lu, Serge Vaudenay, Willi Meier, Liping Ding, and Jianchun Jiang. Synthetic linear analysis: Improved attacks on CubeHash and Rabbit. Lecture Notes in Computer Science, 7259:248–260, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31912-9_17/.

Du:2012:RBF

[178] Yusong Du, Fangguo Zhang, and Meicheng Liu. On the resistance of Boolean functions against fast algebraic attacks. Lecture Notes in Computer Science, 7259:261–274, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31912-9_18/.

Chen:2012:CSI

[179] Yu Chen, Liqun Chen, and Zongyang Zhang. CCA secure IB-KEM from the computational bilinear Diffie–Hellman assumption in the standard model. Lecture Notes in Computer Science, 7259:275–301, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31912-9_19/.

Wolf:2012:DIE

[180] Marko Wolf and Timo Gendrullis. Design, implementation, and evaluation of a vehicular hardware security module. Lecture Notes in Computer Science, 7259:302–318, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31912-9_20/.

Anonymous:2012:FMj

[181] Anonymous. Front matter. Lecture Notes in Computer Science, 7259: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31912-9/1.

Dusterhoft:2012:DTM

[182] Antje Düsterhöft and Meike Klettke. Dedication to a theory of modelling. Lecture Notes in Computer Science, 7260:1–6, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-28279-9_1/.

Paredaens:2012:WAC

[183] Jan Paredaens. What about constraints in RDF? Lecture Notes in Computer Science, 7260:7–18, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28279-9_2/.

Biskup:2012:SRR

[184] Joachim Biskup. Some remarks on relational database schemes having few minimal keys. Lecture Notes in Computer Science, 7260:19– 28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28279-9_3/.

Katona:2012:RDC

[185] Gyula O. H. Katona. Random databases with correlated data. Lecture Notes in Computer Science, 7260:29–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28279-9_4/.

Kallberg:2012:SIR

[186] David Källberg and Nikolaj Leonenko. Statistical inference for Rényi entropy functionals. Lecture Notes in Computer Science, 7260:36–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28279-9_5/.

Borger:2012:SOA

[187] Egon Börger. The subject-oriented approach to software design and the abstract state machines method. Lecture Notes in Computer Science, 7260:52–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28279-9_6/.

Makowsky:2012:BAS

[188] Johann A. Makowsky and Elena V. Ravve. BCNF via attribute splitting. Lecture Notes in Computer Science, 7260:73–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28279-9_7/.

Ferrarotti:2012:FFN

[189] Flavio Ferrarotti and Sven Hartmann. Foundations for a fourth normal form over SQL-like databases. Lecture Notes in Computer Science, 7260: 85–100, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28279-9_8/.

Hegner:2012:IUR

[190] Stephen J. Hegner. Independent update reflections on interdependent database views. Lecture Notes in Computer Science, 7260:101–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28279-9_9/.

Grosso:2012:SRS

[191] Alejandro L. Grosso and José M. Turull Torres. SO^{F}: a semantic restriction over second-order logic and its polynomial-time hierarchy. Lecture Notes in Computer Science, 7260:116–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28279-9_10/.

Wang:2012:ASM

[192] Qing Wang. Abstract state machines for data-parallel computing. Lecture Notes in Computer Science, 7260:136–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28279-9_11/.

Clyde:2012:OLF

[193] Stephen W. Clyde and David W. Embley. OSM-logic: a fact-oriented, time-dependent formalization of object-oriented systems modeling. Lecture Notes in Computer Science, 7260:151–172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28279-9_12/.

Murthy:2012:CDE

[194] Sudarshan Murthy and David Maier. Cloaking data to ease view creation, query expression, and query execution. Lecture Notes in Computer Science, 7260:173–189, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28279-9_13/.

Solvberg:2012:MCD

[195] Arne Sølvberg. On models of concepts and data. Lecture Notes in Computer Science, 7260:190–196, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28279-9_14/.

Kappel:2012:MTE

[196] Gerti Kappel and Philip Langer. Model transformation by-example: a survey of the first wave. Lecture Notes in Computer Science, 7260: 197–215, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28279-9_15/.

Villegas:2012:CIA

[197] Antonio Villegas and Antoni Olivé. On computing the importance of associations in large conceptual schemas. Lecture Notes in Computer Science, 7260:216–230, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28279-9_16/.

Pastor:2012:CMH

[198] Oscar Pastor and Juan Carlos Casamayor. Conceptual modeling of human genome: Integration challenges. Lecture Notes in Computer Science, 7260: 231–250, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28279-9_17/.

Ma:2012:TGE

[199] Hui Ma. Transforming geometrically enhanced conceptual model schemas to GML. Lecture Notes in Computer Science, 7260:251–267, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28279-9_18/.

Duzi:2012:ELH

[200] Marie Duží. Extensional logic of hyperintensions. Lecture Notes in Computer Science, 7260:268–290, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28279-9_19/.

Jaakkola:2012:CSA

[201] Hannu Jaakkola. Culture sensitive aspects in software engineering. Lecture Notes in Computer Science, 7260:291–315, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28279-9_20/.

Anonymous:2012:BMf

[202] Anonymous. Back matter. Lecture Notes in Computer Science, 7260: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-28279-9/1.

Anonymous:2012:FMk

[203] Anonymous. Front matter. Lecture Notes in Computer Science, 7260: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28279-9/1.

Biere:2012:PIT

[204] Armin Biere. Preprocessing and inprocessing techniques in SAT. Lecture Notes in Computer Science, 7261:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-34188-5_1.

Kranen:2012:PFV

[205] Kathryn Kranen. Pioneering the future of verification: a spiral of technological and business innovation. Lecture Notes in Computer Science, 7261:2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/accesspage/chapter/
10.1007/978-3-642-34188-5_2.

Liblit:2012:ADR

[206] Ben Liblit. Automated detection and repair of concurrency bugs. Lecture Notes in Computer Science, 7261:3, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-34188-5_3.

Schubert:2012:VCW

[207] Klaus-Dieter Schubert. Verification challenges of workload optimized hardware systems. Lecture Notes in Computer Science, 7261:4, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-34188-5_4.

Kupferman:2012:SC

[208] Orna Kupferman, Dorsa Sadigh, and Sanjit A. Seshia. Synthesis with clairvoyance. Lecture Notes in Computer Science, 7261:5–19, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34188-5_5/.

Schlaipfer:2012:GRS

[209] Matthias Schlaipfer, Georg Hofferek, and Roderick Bloem. Generalized reactivity(1) synthesis without a monolithic strategy. Lecture Notes in Computer Science, 7261:20–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34188-5_6/.

Bu:2012:IGD

[210] Lei Bu, Yang Yang, and Xuandong Li. IIS-guided DFS for efficient bounded reachability analysis of linear hybrid automata. Lecture Notes in Computer Science, 7261:35–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34188-5_7/.

Heule:2012:CCG

[211] Marijn J. H. Heule, Oliver Kullmann, Siert Wieringa, and Armin Biere. Cube and conquer: Guiding CDCL SAT solvers by lookaheads. Lecture Notes in Computer Science, 7261:50–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34188-5_8/.

Khasidashvili:2012:ISS

[212] Zurab Khasidashvili and Alexander Nadel. Implicative simultaneous satisfiability and applications. Lecture Notes in Computer Science, 7261: 66–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34188-5_9/.

Krishna:2012:LVS

[213] B. A. Krishna, Jonathan Michelson, Vigyan Singhal, and Alok Jain. Liveness vs safety — a practical viewpoint. Lecture Notes in Computer Science, 7261:80–94, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34188-5_10/.

Sinha:2012:PSV

[214] Arnab Sinha, Sharad Malik, Chao Wang, and Aarti Gupta. Predicting serializability violations: SMT-based search vs. DPOR-based search. Lecture Notes in Computer Science, 7261:95–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34188-5_11/.

Chen:2012:SSA

[215] Qichang Chen, Liqiang Wang, and Zijiang Yang. SAM: Self-adaptive dynamic analysis for multithreaded programs. Lecture Notes in Computer Science, 7261:115–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34188-5_12/.

Huth:2012:CSP

[216] Michael Huth, Jim Huan-Pu Kuo, and Nir Piterman. Concurrent small progress measures. Lecture Notes in Computer Science, 7261:130–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34188-5_13/.

Johnson:2012:SQA

[217] Kenneth Johnson, Simon Reed, and Radu Calinescu. Specification and quantitative analysis of probabilistic cloud deployment patterns. Lecture Notes in Computer Science, 7261:145–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34188-5_14/.

Sery:2012:IBF

[218] Ondrej Sery, Grigory Fedyukovich, and Natasha Sharygina. Interpolationbased function summaries in bounded model checking. Lecture Notes in Computer Science, 7261:160–175, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34188-5_15/.

Ostrand:2012:CFL

[219] Thomas J. Ostrand and Elaine J. Weyuker. Can file level characteristics help identify system level fault-proneness? Lecture Notes in Computer Science, 7261:176–189, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34188-5_16/.

Birnbaum:2012:RCA

[220] Ariel Birnbaum, Laurent Fournier, Steve Mittermaier, and Avi Ziv. Reverse coverage analysis. Lecture Notes in Computer Science, 7261: 190–202, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34188-5_17/.

Collingbourne:2012:STO

[221] Peter Collingbourne, Cristian Cadar, and Paul H. J. Kelly. Symbolic testing of OpenCL code. Lecture Notes in Computer Science, 7261: 203–218, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34188-5_18/.

Adir:2012:DTD

[222] Allon Adir, Ronen Levy, and Tamer Salman. Dynamic test data generation for data intensive applications. Lecture Notes in Computer Science, 7261:219–233, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34188-5_19/.

Aharony:2012:IFP

[223] Merav Aharony, Emanuel Gofman, Elena Guralnik, and Anatoly Koyfman. Injecting floating-point testing knowledge into test generators. Lecture Notes in Computer Science, 7261:234–241, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-34188-5_20/.

Anonymous:2012:FMl

[224] Anonymous. Front matter. Lecture Notes in Computer Science, 7261: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34188-5/1.

Alipanahi:2012:PSS

[225] Babak Alipanahi, Nathan Krislock, and Ali Ghodsi. Protein structure by semidefinite facial reduction. Lecture Notes in Computer Science, 7262:1–11, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29627-7_1/.

Bercovici:2012:AIC

[226] Sivan Bercovici and Jesse M. Rodriguez. Ancestry inference in complex admixtures via variable-length Markov chain linkage models. Lecture Notes in Computer Science, 7262:12–28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29627-7_2/.

Canzar:2012:CGP

[227] Stefan Canzar, Mohammed El-Kebir, and René Pool. Charge group partitioning in biomolecular simulation. Lecture Notes in Computer Science, 7262:29–43, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29627-7_3/.

Bravo:2012:IMV

[228] Hector Corrada Bravo. Increased methylation variation in epigenetic domains across cancer types. Lecture Notes in Computer Science, 7262:44, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-29627-7_4.

DeBlasio:2012:EAM

[229] Dan F. DeBlasio and Travis J. Wheeler. Estimating the accuracy of multiple alignments and its use in parameter advising. Lecture Notes in Computer Science, 7262:45–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29627-7_5/.

Francis:2012:QDC

[230] Matthew R. Francis and Elana J. Fertig. Quantifying the dynamics of coupled networks of switches and oscillators. Lecture Notes in Computer Science, 7262:60–61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-29627-7_6.

Gymrek:2012:LST

[231] Melissa Gymrek, David Golan, and Saharon Rosset. lobSTR: a short tandem repeat profiler for personal genomes. Lecture Notes in Computer Science, 7262:62–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-29627-7_7.

He:2012:HSO

[232] Dan He, Buhm Han, and Eleazar Eskin. Hap-seq: An optimal algorithm for haplotype phasing with imputation using sequencing data. Lecture Notes in Computer Science, 7262:64–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29627-7_8/.

He:2012:BBB

[233] Lu He, Fabio Vandin, and Gopal Pandurangan. Ballast: a ball-based algorithm for structural motifs. Lecture Notes in Computer Science, 7262: 79–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29627-7_9/.

Holloway:2012:EGO

[234] Patrick Holloway and Krister Swenson. Evolution of genome organization by duplication and loss: An alignment approach. Lecture Notes in Computer Science, 7262:94–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29627-7_10/.

Holtby:2012:LLM

[235] Daniel Holtby, Shuai Cheng Li, and Ming Li. LoopWeaver — loop modeling by the weighted scaling of verified proteins. Lecture Notes in Computer Science, 7262:113–126, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29627-7_11/.

Huang:2012:RMT

[236] Yan Huang, Yin Hu, and Corbin D. Jones. A robust method for transcript quantification with RNA-seq data. Lecture Notes in Computer Science, 7262:127–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29627-7_12/.

Kinsella:2012:MBF

[237] Marcus Kinsella and Vineet Bafna. Modeling the breakage-fusion-bridge mechanism: Combinatorics and cancer genomics. Lecture Notes in Computer Science, 7262:148–162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29627-7_13/.

Li:2012:TST

[238] Yang Li, Hong-Mei Li, and Paul Burns. TrueSight: Self-training algorithm for splice junction detection using RNA-seq. Lecture Notes in Computer Science, 7262:163–164, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-29627-7_14.

Lin:2012:SSD

[239] Yaw-Ling Lin, Charles Ward, and Steven Skiena. Synthetic sequence design for signal location search. Lecture Notes in Computer Science, 7262: 165–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29627-7_15/.

Marti-Renom:2012:TDA

[240] Marc A. Marti-Renom. The three-dimensional architecture of a bacterial genome and its alteration by genetic perturbation. Lecture Notes in Computer Science, 7262:180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-29627-7_16.

McPherson:2012:DCG

[241] Andrew McPherson and Chunxiao Wu. Discovery of complex genomic rearrangements in cancer using high-throughput sequencing. Lecture Notes in Computer Science, 7262:181–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-29627-7_17.

Murali:2012:NBP

[242] T. M. Murali, Matthew D. Dyer, and David Badger. Network-based prediction and analysis of HIV dependency factors. Lecture Notes in Computer Science, 7262:183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-29627-7_18.

Parker:2012:SGD

[243] Andrew S. Parker and Karl E. Griswold. Structure-guided deimmunization of therapeutic proteins. Lecture Notes in Computer Science, 7262: 184–198, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29627-7_19/.

Pervouchine:2012:EWA

[244] Dmitri Pervouchine and Ekaterina Khrameeva. Evidence for widespread association of mammalian splicing and conserved long-range RNA structures. Lecture Notes in Computer Science, 7262:199, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-29627-7_20.

Anonymous:2012:FMm

[245] Anonymous. Front matter. Lecture Notes in Computer Science, 7262: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29627-7/1.

Chang:2012:SRP

[246] Ping-Lin Chang, Dongbin Chen, and Daniel Cohen. 2D/3D registration of a preoperative model with endoscopic video using colour-consistency. Lecture Notes in Computer Science, 7264:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32630-1_1/.

Drouin:2012:RTD

[247] Simon Drouin and Marta Kersten-Oertel. A realistic test and development environment for mixed reality in neurosurgery. Lecture Notes in Computer Science, 7264:13–23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32630-1_2/.

Fujii:2012:VSB

[248] Kenko Fujii, Johannes Totz, and Guang-Zhong Yang. Visual search behaviour and analysis of augmented visualisation for minimally invasive surgery. Lecture Notes in Computer Science, 7264:24–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32630-1_3/.

Gavaghan:2012:ARI

[249] Kate Alicia Gavaghan and Sylvain Anderegg. Augmented reality image overlay projection for image guided open liver ablation of metastatic liver cancer. Lecture Notes in Computer Science, 7264:36–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32630-1_4/.

Giannarou:2012:TDR

[250] Stamatia Giannarou and Guang-Zhong Yang. Tissue deformation recovery with Gaussian mixture model based structure from motion. Lecture Notes in Computer Science, 7264:47–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32630-1_5/.

Horvath:2012:TUP

[251] Samantha Horvath, John Galeotti, and Bo Wang. Towards an ultrasound probe with vision: Structured light to determine surface orientation. Lecture Notes in Computer Science, 7264:58–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32630-1_6/.

Jayender:2012:MMC

[252] Jagadeesan Jayender and Inbar Spofford. Markov modeling of colonoscopy gestures to develop skill trainers. Lecture Notes in Computer Science, 7264:65–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32630-1_7/.

Kainz:2012:VVC

[253] Bernhard Kainz and Rupert H. Portugaller. Volume visualization in the clinical practice. Lecture Notes in Computer Science, 7264:74–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32630-1_8/.

Lang:2012:CUR

[254] Pencilla Lang and Michael W. Chu. CT-US registration for guidance of transcatheter aortic valve implantation. Lecture Notes in Computer Science, 7264:85–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32630-1_9/.

Linte:2012:EPI

[255] Cristian A. Linte and Kurt E. Augustine. Enhanced planning of interventions for spinal deformity correction using virtual modeling and visualization techniques. Lecture Notes in Computer Science, 7264: 93–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32630-1_10/.

Metz:2012:ACC

[256] Coert Metz, Michiel Schaap, and Stefan Klein. Alignment of 4D coronary CTA with monoplane X-ray angiography. Lecture Notes in Computer Science, 7264:106–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32630-1_11/.

Suzuki:2012:VTS

[257] Naoki Suzuki, Asaki Hattori, and Satoshi Ieiri. VR training system for endoscopic surgery robot. Lecture Notes in Computer Science, 7264: 117–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32630-1_12/.

Wu:2012:BPA

[258] Jue Wu, Kathryn Davis, and Allan Azarion. Brain parcellation aids in electrode localization in epileptic patients. Lecture Notes in Computer Science, 7264:130–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32630-1_13/.

Anonymous:2012:BMg

[259] Anonymous. Back matter. Lecture Notes in Computer Science, 7264: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-32630-1/1.

Anonymous:2012:FMn

[260] Anonymous. Front matter. Lecture Notes in Computer Science, 7264: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32630-1/1.

Minker:2012:VLH

[261] Jack Minker. To Vladimir Lifschitz on his 65th birthday. Lecture Notes in Computer Science, 7265:1–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30743-0_1/.

Kartha:2012:VLY

[262] Neelakantan Kartha, Esra Erdem, Joohyung Lee, and Paolo Ferraris. Vladimir Lifschitz — a youth at 65. Lecture Notes in Computer Science, 7265:14–23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30743-0_2/.

Balduccini:2012:CAE

[263] Marcello Balduccini. A “conservative” approach to extending answer set programming with non-herbrand functions. Lecture Notes in Computer Science, 7265:24–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30743-0_3/.

Baral:2012:ILC

[264] Chitta Baral, Marcos Alvarez Gonzalez, and Aaron Gottesman. The inverse lambda calculus algorithm for typed first order logic lambda calculus and its application to translating English to FOL. Lecture Notes in Computer Science, 7265:40–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30743-0_4/.

Baumann:2012:PSS

[265] Ringo Baumann, Gerhard Brewka, Wolfgang Dvořák, and Stefan Woltran. Parameterized splitting: a simple modification-based approach. Lecture Notes in Computer Science, 7265:57–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30743-0_5/.

Bjorner:2012:PIL

[266] Nikolaj Bjørner, Guido de Caso, and Yuri Gurevich. From primal infon logic with individual variables to datalog. Lecture Notes in Computer Science, 7265:72–86, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30743-0_6/.

Bochman:2012:HTA

[267] Alexander Bochman. Here and there among logics for logic programming. Lecture Notes in Computer Science, 7265:87–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30743-0_7/.

Cabalar:2012:CLP

[268] Pedro Cabalar. Causal logic programming. Lecture Notes in Computer Science, 7265:102–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30743-0_8/.

Cali:2012:IER

[269] Andrea Calì, Georg Gottlob, Giorgio Orsi, and Andreas Pieris. On the interaction of existential rules and equality constraints in ontology querying. Lecture Notes in Computer Science, 7265:117–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30743-0_9/.

Chen:2012:EAL

[270] Xiaoping Chen, Guoqiang Jin, and Fangkai Yang. Extending action language C+ by formalizing composite actions. Lecture Notes in Computer Science, 7265:134–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30743-0_10/.

Costantini:2012:SER

[271] Stefania Costantini, Andrea Formisano, and David Pearce. Strong equivalence of RASP programs. Lecture Notes in Computer Science, 7265: 149–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30743-0_11/.

Delgrande:2012:CBR

[272] James Delgrande. Considerations on belief revision in an action theory. Lecture Notes in Computer Science, 7265:164–177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30743-0_12/.

Denecker:2012:AFT

[273] Marc Denecker, Maurice Bruynooghe, and Joost Vennekens. Approximation fixpoint theory and the semantics of logic and answers set programs. Lecture Notes in Computer Science, 7265:178–194, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30743-0_13/.

Dix:2012:PPU

[274] Jürgen Dix, Wolfgang Faber, and V. S. Subrahmanian. Privacy preservation using multi-context systems and default logic. Lecture Notes in Computer Science, 7265:195–210, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30743-0_14/.

Eiter:2012:SPR

[275] Thomas Eiter, Cristina Feier, and Michael Fink. Simulating production rules using ACTHEX. Lecture Notes in Computer Science, 7265: 211–228, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30743-0_15/.

Erdem:2012:AAL

[276] Esra Erdem and Volkan Patoglu. Applications of action languages in cognitive robotics. Lecture Notes in Computer Science, 7265:229–246, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30743-0_16/.

Faber:2012:IGD

[277] Wolfgang Faber, Nicola Leone, and Simona Perri. The intelligent grounder of DLV. Lecture Notes in Computer Science, 7265:247–264, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30743-0_17/.

delCerro:2012:BSL

[278] Luis Fariñas del Cerro, David Pearce, and Agustín Valverde. Bi-state logic. Lecture Notes in Computer Science, 7265:265–278, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30743-0_18/.

Gabbay:2012:EAL

[279] Dov M. Gabbay. An equational approach to logic programming. Lecture Notes in Computer Science, 7265:279–295, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30743-0_19/.

Gebser:2012:GEA

[280] Martin Gebser, Roland Kaufmann, and Torsten Schaub. Gearing up for effective ASP planning. Lecture Notes in Computer Science, 7265: 296–310, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30743-0_20/.

Anonymous:2012:FMo

[281] Anonymous. Front matter. Lecture Notes in Computer Science, 7265: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30743-0/1.

Fasbender:2012:CNT

[282] Andreas Fasbender, Martin Gerdes, and Sascha Smets. Cellular networking technologies in ITS solutions: Opportunities and challenges. Lecture Notes in Computer Science, 7266:1–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29667-3_1/.

Idrees:2012:EES

[283] Muhammad Sabir Idrees and Yves Roudier. Effective and efficient security policy engines for automotive on-board networks. Lecture Notes in Computer Science, 7266:14–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29667-3_2/.

Sikora:2012:SSH

[284] Axel Sikora. Security solutions for highly dynamic Car2X networks in the KoFAS initiative. Lecture Notes in Computer Science, 7266: 27–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29667-3_3/.

Alazawi:2012:IIC

[285] Zubaida Alazawi, Mohmmad B. Abdljabar, Saleh Altowaijri, and Anna Maria Vegni. ICDMS: An intelligent cloud based disaster management system for vehicular networks. Lecture Notes in Computer Science, 7266:40–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29667-3_4/.

Rozanowski:2012:WDV

[286] Krzysztof Różanowski, Zbigniew Piotrowski, Tadeusz Sondej, and Krzysztof Sawicki. Wireless driver and vehicle surveillance system based on IEEE 802.11 networks. Lecture Notes in Computer Science, 7266: 57–67, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29667-3_5/.

Rozanowski:2012:ACM

[287] Krzysztof Różanowski, Tadeusz Sondej, Zbigniew Piotrowski, and Krzysztof Sawicki. Architecture of car measurement system for driver monitoring. Lecture Notes in Computer Science, 7266:68–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29667-3_6/.

Ganan:2012:TRD

[288] Carlos Gañán, Jose L. Muñoz, Oscar Esparza, Jorge Mata-Díaz, and Juanjo Alins. Toward revocation data handling efficiency in VANETs. Lecture Notes in Computer Science, 7266:80–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29667-3_7/.

Shoaib:2012:CBD

[289] Muhammad Shoaib, Wang-Cheol Song, and Keun Hyung Kim. Cluster based data aggregation in vehicular adhoc network. Lecture Notes in Computer Science, 7266:91–102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29667-3_8/.

Kaisser:2012:SVS

[290] Florent Kaisser, Christophe Gransart, and Marion Berbineau. Simulations of VANET scenarios with OPNET and SUMO. Lecture Notes in Computer Science, 7266:103–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29667-3_9/.

Jagyasi:2012:HPS

[291] Bhushan G. Jagyasi, Vikrant Kumar, and Arun Pande. Human participatory sensing in fixed route bus information system. Lecture Notes in Computer Science, 7266:113–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29667-3_10/.

Kowalski:2012:SVT

[292] Dariusz Kowalski, Zeev Nutov, and Michael Segal. Scheduling of vehicles in transportation networks. Lecture Notes in Computer Science, 7266: 124–136, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29667-3_11/.

Bergs:2012:DPT

[293] Johan Bergs, Erwin Van de Velde, Daan Pareit, Dries Naudts, and Milos Rovcanin. Design and prototype of a train-to-wayside communication architecture. Lecture Notes in Computer Science, 7266:137–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29667-3_12/.

Kumar:2012:CAD

[294] Navin Kumar and Nidhu Kumari. Conceptual architectural design of Indian railway intelligent transportation systems. Lecture Notes in Computer Science, 7266:151–162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29667-3_13/.

Jha:2012:AAG

[295] Abhinav Jha, Amit Kumar Agrawal, and Chirabrata Bhaumik. Automatic alert generation from train to the people at unmanned level crossings using principles of IoT. Lecture Notes in Computer Science, 7266: 163–173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29667-3_14/.

Noriega-Vivas:2012:MNA

[296] Patricia Noriega-Vivas, Celeste Campo, Carlos Garcia-Rubio, and Alicia Rodriguez-Carrion. MOFETA: a network architecture based on MObile FEmtocells to Enhance Cellular Connectivity on TrAins. Lecture Notes in Computer Science, 7266:174–185, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29667-3_15/.

Anonymous:2012:BMh

[297] Anonymous. Back matter. Lecture Notes in Computer Science, 7266: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-29667-3/1.

Anonymous:2012:FMp

[298] Anonymous. Front matter. Lecture Notes in Computer Science, 7266: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29667-3/1.

Anonymous:2012:FMq

[299] Anonymous. Front matter. Lecture Notes in Computer Science, 7267: 1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29347-4/1/1.

Bartecki:2012:NNB

[300] Krzysztof Bartecki. Neural network-based PCA: An application to approximation of a distributed parameter system. Lecture Notes in Computer Science, 7267:3–11, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29347-4_1/.

Bilski:2012:PRR

[301] Jarosław Bilski and Jacek Smolag. Parallel realisation of the recurrent multi layer perceptron learning. Lecture Notes in Computer Science, 7267:12–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29347-4_2/.

Capizzi:2012:IHN

[302] Giacomo Capizzi, Christian Napoli, and Lucio Paternò. An innovative hybrid neuro-wavelet method for reconstruction of missing data in astronomical photometric surveys. Lecture Notes in Computer Science, 7267:21–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29347-4_3/.

Chevitarese:2012:STN

[303] Daniel Salles Chevitarese, Dilza Szwarcman, and Marley Vellasco. Speeding up the training of neural networks with CUDA technology. Lecture Notes in Computer Science, 7267:30–38, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29347-4_4/.

Er:2012:UCO

[304] Meng Joo Er and Piotr Duda. On the uniform convergence of the orthogonal series-type kernel regression neural networks in a time-varying environment. Lecture Notes in Computer Science, 7267:39–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29347-4_5/.

Duda:2012:SCO

[305] Piotr Duda, Yoichi Hayashi, and Maciej Jaworski. On the strong convergence of the orthogonal series-type kernel regression neural networks in a non-stationary environment. Lecture Notes in Computer Science, 7267:47–54, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29347-4_6/.

Duda:2012:SCR

[306] Piotr Duda and Marcin Korytkowski. On the strong convergence of the recursive orthogonal series-type kernel probabilistic neural networks handling time-varying noise. Lecture Notes in Computer Science, 7267: 55–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29347-4_7/.

Fiksak:2012:INN

[307] Bogumił Fiksak and Maciej Krawczak. Incidental neural networks as nomograms generators. Lecture Notes in Computer Science, 7267:63– 71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29347-4_8/.

Halawa:2012:SAF

[308] Krzysztof Halawa. Selection of activation functions in the last hidden layer of the multilayer perceptron. Lecture Notes in Computer Science, 7267:72–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29347-4_9/.

Horzyk:2012:IFA

[309] Adrian Horzyk. Information freedom and associative artificial intelligence. Lecture Notes in Computer Science, 7267:81–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29347-4_10/.

Jaworski:2012:APT

[310] Maciej Jaworski, Meng Joo Er, and Lena Pietruczuk. On the application of the Parzen-type kernel regression neural network and order statistics for learning in a non-stationary environment. Lecture Notes in Computer Science, 7267:90–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29347-4_11/.

Jaworski:2012:LTV

[311] Maciej Jaworski and Marcin Gabryel. On learning in a time-varying environment by using a probabilistic neural network and the recursive least squares method. Lecture Notes in Computer Science, 7267: 99–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29347-4_12/.

Kryzhanovskiy:2012:BPL

[312] Vladimir Kryzhanovskiy, Irina Zhelavskaya, and Jakov Karandashev. Binary perceptron learning algorithm using simplex-method. Lecture Notes in Computer Science, 7267:111–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29347-4_13/.

Laskowski:2012:OAS

[313] Łukasz Laskowski. Objects auto-selection from stereo-images realised by self-correcting neural network. Lecture Notes in Computer Science, 7267: 119–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29347-4_14/.

Lawrynczuk:2012:LTB

[314] Maciej Ławryńczuk. On-line trajectory-based linearisation of neural models for a computationally efficient predictive control algorithm. Lecture Notes in Computer Science, 7267:126–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29347-4_15/.

Olej:2012:STS

[315] Vladimir Olej and Jana Filipova. Short time series of Website visits prediction by RBF neural networks and support vector machine regression. Lecture Notes in Computer Science, 7267:135–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29347-4_16/.

Piersa:2012:SSF

[316] Jarosław Piersa and Tomasz Schreiber. Spectra of the spike-flow graphs in geometrically embedded neural networks. Lecture Notes in Computer Science, 7267:143–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29347-4_17/.

Pietruczuk:2012:WCP

[317] Lena Pietruczuk and Meng Joo Er. Weak convergence of the Parzen-type probabilistic neural network handling time-varying noise. Lecture Notes in Computer Science, 7267:152–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29347-4_18/.

Pietruczuk:2012:SCR

[318] Lena Pietruczuk and Yoichi Hayashi. Strong convergence of the recursive Parzen-type probabilistic neural network handling nonstationary noise. Lecture Notes in Computer Science, 7267:160–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29347-4_19/.

Plonski:2012:IPS

[319] Piotr Płoński and Krzysztof Zaremba. Improving performance of selforganising maps with distance metric learning method. Lecture Notes in Computer Science, 7267:169–177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29347-4_20/.

Anonymous:2012:FMr

[320] Anonymous. Front matter. Lecture Notes in Computer Science, 7267: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29347-4/1.

Anonymous:2012:FMs

[321] Anonymous. Front matter. Lecture Notes in Computer Science, 7268: 1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29350-4/1/1.

Abgaz:2012:DAO

[322] Yalemisew M. Abgaz, Muhammad Javed, and Claus Pahl. Dependency analysis in ontology-driven content-based systems. Lecture Notes in Computer Science, 7268:3–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29350-4_1/.

Bartik:2012:MWP

[323] Vladimír Bartík. Measuring Web page similarity based on textual and visual properties. Lecture Notes in Computer Science, 7268:13– 21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29350-4_2/.

Brodowski:2012:NSH

[324] Stanisław Brodowski and Andrzej Bielecki. New specifics for a hierarchial estimator meta-algorithm. Lecture Notes in Computer Science, 7268:22–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29350-4_3/.

Ceglarek:2012:FPD

[325] Dariusz Ceglarek and Konstanty Haniewicz. Fast plagiarism detection by sentence hashing. Lecture Notes in Computer Science, 7268: 30–37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29350-4_4/.

Chang:2012:EDS

[326] Harry Chang. Enriching domain-specific language models using domain independent WWW N-gram corpus. Lecture Notes in Computer Science, 7268:38–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29350-4_5/.

Codara:2012:SIR

[327] Pietro Codara. On the structure of indiscernibility relations compatible with a partially ordered set. Lecture Notes in Computer Science, 7268:47–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29350-4_6/.

Duda:2012:PPA

[328] Piotr Duda, Maciej Jaworski, and Lena Pietruczuk. On pre-processing algorithms for data stream. Lecture Notes in Computer Science, 7268: 56–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29350-4_7/.

Grochowski:2012:SII

[329] Marek Grochowski. Simple incremental instance selection wrapper for classification. Lecture Notes in Computer Science, 7268:64–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29350-4_8/.

Gupta:2012:MMN

[330] Anamika Gupta, Naveen Kumar, and Vasudha Bhatnagar. Mining of multiobjective non-redundant association rules in data streams. Lecture Notes in Computer Science, 7268:73–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29350-4_9/.

Jaworski:2012:FCD

[331] Maciej Jaworski, Piotr Duda, and Lena Pietruczuk. On fuzzy clustering of data streams with concept drift. Lecture Notes in Computer Science, 7268:82–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29350-4_10/.

Jaworski:2012:ROF

[332] Maciej Jaworski, Lena Pietruczuk, and Piotr Duda. On resources optimization in fuzzy clustering of data streams. Lecture Notes in Computer Science, 7268:92–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29350-4_11/.

Klesk:2012:CCS

[333] Przemysław Klesk. A comparison of complexity selection approaches for polynomials based on: Vapnik–Chervonenkis dimension, Rademacher complexity and covering numbers. Lecture Notes in Computer Science, 7268:100–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29350-4_12/.

Korzen:2012:SCL

[334] Marcin Korzeń and Przemysław Klesk. Sample complexity of linear learning machines with different restrictions over weights. Lecture Notes in Computer Science, 7268:111–119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29350-4_13/.

Krawczak:2012:CAB

[335] Maciej Krawczak and Grażyna Szkatuła. A clustering algorithm based on distinguishability for nominal attributes. Lecture Notes in Computer Science, 7268:120–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29350-4_14/.

Ladyzynski:2012:RIC

[336] Piotr Ładyżyński and Przemysław Grzegorzewski. Retrieving informative content from Web pages with conditional learning of support vector machines and semantic analysis. Lecture Notes in Computer Science, 7268: 128–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29350-4_15/.

Maciejewski:2012:ERW

[337] Henryk Maciejewski, Ewa Walkowicz, Olgierd Unold, and Paweł Skrobanek. Enhancing recognition of a weak class — comparative study based on biological population data mining. Lecture Notes in Computer Science, 7268:136–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29350-4_16/.

Michalak:2012:FRB

[338] Marcin Michalak. Foundations of rough biclustering. Lecture Notes in Computer Science, 7268:144–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29350-4_17/.

Michalak:2012:OOR

[339] Marcin Michalak, Marek Sikora, and Patryk Ziarnik. ORG- oblique rules generator. Lecture Notes in Computer Science, 7268:152–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29350-4_18/.

Plucinski:2012:MML

[340] Marcin Pluciński. Mini-models — local regression models for the function approximation learning. Lecture Notes in Computer Science, 7268: 160–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29350-4_19/.

Starczewski:2012:CVI

[341] Artur Starczewski. A cluster validity index for hard clustering. Lecture Notes in Computer Science, 7268:168–174, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29350-4_20/.

Anonymous:2012:FMt

[342] Anonymous. Front matter. Lecture Notes in Computer Science, 7268: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29350-4/1.

Anonymous:2012:FMu

[343] Anonymous. Front matter. Lecture Notes in Computer Science, 7269: 1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29353-5/1/1.

Kountche:2012:PAA

[344] Djibrilla Amadou Kountché, Nicolas Monmarché, and Mohamed Slimane. The Pachycondyla Apicalis ants search strategy for data clustering problems. Lecture Notes in Computer Science, 7269:3–11, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29353-5_1/.

Arabas:2012:PMP

[345] Jarosław Arabas, Ogier Maitre, and Pierre Collet. PARADE: a massively parallel differential evolution template for EASEA. Lecture Notes in Computer Science, 7269:12–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29353-5_2/.

Bai:2012:DCS

[346] Li Bai and David Feltell. A 3D discrete-continuum swarm intelligence simulation on GPU for swarm robotics and biomedical applications. Lecture Notes in Computer Science, 7269:21–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29353-5_3/.

Bolla:2012:GFR

[347] Kálmán Bolla, Tamás Kovacs, and Gábor Fazekas. Gathering of fat robots with limited visibility and without global navigation. Lecture Notes in Computer Science, 7269:30–38, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29353-5_4/.

Bujok:2012:PMM

[348] Petr Bujok and Josef Tvrdík. Parallel migration model employing various adaptive variants of differential evolution. Lecture Notes in Computer Science, 7269:39–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29353-5_5/.

DePrisco:2012:DEA

[349] Roberto De Prisco, Gianluca Zaccagnino, and Rocco Zaccagnino. A differential evolution algorithm assisted by ANFIS for music fingering. Lecture Notes in Computer Science, 7269:48–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29353-5_6/.

ElDor:2012:HDE

[350] Abbas El Dor, Maurice Clerc, and Patrick Siarry. Hybridization of differential evolution and particle swarm optimization in a new algorithm: DEPSO-2S. Lecture Notes in Computer Science, 7269:57–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29353-5_7/.

Fister:2012:HAB

[351] Iztok Fister Jr., Iztok Fister, and Janez Brest. A hybrid artificial bee colony algorithm for graph 3-coloring. Lecture Notes in Computer Science, 7269:66–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29353-5_8/.

Fix:2012:MCS

[352] Jeremy Fix and Matthieu Geist. Monte–Carlo swarm policy search. Lecture Notes in Computer Science, 7269:75–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29353-5_9/.

Iacca:2012:CBF

[353] Giovanni Iacca, Ferrante Neri, and Ernesto Mininno. Compact bacterial foraging optimization. Lecture Notes in Computer Science, 7269: 84–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29353-5_10/.

Koh:2012:CMA

[354] Andrew Koh. A coevolutionary MiniMax algorithm for the detection of Nash equilibrium. Lecture Notes in Computer Science, 7269:93–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29353-5_11/.

Krzeszowski:2012:RTT

[355] Tomasz Krzeszowski, Bogdan Kwolek, Boguslaw Rymut, and Konrad Wojciechowski. Real-time tracking of full-body motion using parallel particle swarm optimization with a pool of best particles. Lecture Notes in Computer Science, 7269:102–109, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29353-5_12/.

Opara:2012:DMM

[356] Karol Opara and Jarosław Arabas. Decomposition and metaoptimization of mutation operator in differential evolution. Lecture Notes in Computer Science, 7269:110–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29353-5_13/.

Paplinski:2012:CAC

[357] Janusz Papliński. Continuous ant colony optimization for identification of time delays in the linear plant. Lecture Notes in Computer Science, 7269: 119–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29353-5_14/.

Tasgetiren:2012:VIG

[358] M. Fatih Tasgetiren, Quan-Ke Pan, P. N. Suganthan, and Ozge Buyukdagli. A variable iterated greedy algorithm with differential evolution for solving no-idle flowshops. Lecture Notes in Computer Science, 7269:128–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29353-5_15/.

Tvrdik:2012:DEC

[359] Josef Tvrdík and Ivan Křivý. Differential evolution with competing strategies applied to partitional clustering. Lecture Notes in Computer Science, 7269:136–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29353-5_16/.

Weber:2012:CBC

[360] Matthieu Weber and Ferrante Neri. Contiguous binomial crossover in differential evolution. Lecture Notes in Computer Science, 7269:145– 153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29353-5_17/.

Zamuda:2012:PRD

[361] Aleš Zamuda and Janez Brest. Population reduction differential evolution with multiple mutation strategies in real world industry challenges. Lecture Notes in Computer Science, 7269:154–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29353-5_18/.

Anonymous:2012:FMv

[362] Anonymous. Front matter. Lecture Notes in Computer Science, 7269: 163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/content/pdf/bfm:978-
3-642-29353-5/2/1.

Cermak:2012:GOF

[363] Petr Cermak and Michal Mura. Genetic optimization of fuzzy rule based MAS using cognitive analysis. Lecture Notes in Computer Science, 7269: 165–173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29353-5_19/.

Czajkowski:2012:DMA

[364] Marcin Czajkowski and Marek Kretowski. Does memetic approach improve global induction of regression and model trees? Lecture Notes in Computer Science, 7269:174–181, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29353-5_20/.

Anonymous:2012:FMw

[365] Anonymous. Front matter. Lecture Notes in Computer Science, 7269: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29353-5/1.

Burato:2012:PRA

[366] Elisa Burato and Matteo Cristani. The process of reaching agreement in meaning negotiation. Lecture Notes in Computer Science, 7270:1– 42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32066-8_1/.

Jascanu:2012:FEC

[367] Veronica Jascanu and Nicolae Jascanu. Formalizing emotional Ecommerce agents for a simple negotiation protocol. Lecture Notes in Computer Science, 7270:43–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32066-8_2/.

Fortino:2012:EMA

[368] Giancarlo Fortino and Francesco Rango. Engineering multi-agent systems through statecharts-based JADE agents and tools. Lecture Notes in Computer Science, 7270:61–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32066-8_3/.

Zargayouna:2012:FOM

[369] Mahdi Zargayouna. Fleet organization models for online vehicle routing problems. Lecture Notes in Computer Science, 7270:82–102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32066-8_4/.

Adanhounme:2012:NSF

[370] Villèvo Adanhounmè and Théophile K. Dagba. Neural smooth function approximation and prediction with adaptive learning rate. Lecture Notes in Computer Science, 7270:103–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32066-8_5/.

OShea:2012:MCA

[371] James O’Shea, Zuhair Bandar, and Keeley Crockett. A multi-classifier approach to dialogue act classification using function words. Lecture Notes in Computer Science, 7270:119–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32066-8_6/.

Zakrzewska:2012:BGR

[372] Danuta Zakrzewska. Building group recommendations in E-learning systems. Lecture Notes in Computer Science, 7270:144–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32066-8_7/.

Lorkiewicz:2012:ISM

[373] Wojciech Lorkiewicz and Radoslaw Katarzyniak. Individual semiosis in multi-agent systems. Lecture Notes in Computer Science, 7270:164– 197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32066-8_8/.

DiBitonto:2012:EMA

[374] Pierpaolo Di Bitonto and Maria Laterza. Evaluation of multi-agent systems: Proposal and validation of a metric plan. Lecture Notes in Computer Science, 7270:198–221, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32066-8_9/.

Was:2012:EMT

[375] Jarosław Was. Egress modeling through cellular automata based multiagent systems. Lecture Notes in Computer Science, 7270:222–235, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32066-8_10/.

Anonymous:2012:BMi

[376] Anonymous. Back matter. Lecture Notes in Computer Science, 7270: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-32066-8/1.

Anonymous:2012:FMx

[377] Anonymous. Front matter. Lecture Notes in Computer Science, 7270: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32066-8/1.

VanLanduyt:2012:DDD

[378] Dimitri Van Landuyt, Steven Op de beeck, Eddy Truyen, and Wouter Joosen. Domain-driven discovery of stable abstractions for pointcut interfaces. Lecture Notes in Computer Science, 7271:1–52, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35551-6_1/.

Tigli:2012:AAT

[379] Jean-Yves Tigli, Stéphane Lavirotte, Gaëtan Rey, Nicolas Ferry, and Vincent Hourdin. Aspect of assembly: From theory to performance. Lecture Notes in Computer Science, 7271:53–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35551-6_2/.

Ansaloni:2012:DAO

[380] Danilo Ansaloni, Walter Binder, Philippe Moret, and Alex Villazón. Dynamic aspect-oriented programming in Java: The HotWave experience. Lecture Notes in Computer Science, 7271:92–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35551-6_3/.

Holzer:2012:ACE

[381] Adrian Holzer, Lukasz Ziarek, K. R. Jayaram, and Patrick Eugster. Abstracting context in event-based software. Lecture Notes in Computer Science, 7271:123–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35551-6_4/.

Lohmann:2012:AAD

[382] Daniel Lohmann, Olaf Spinczyk, Wanja Hofer, and Wolfgang Schröder-Preikschat. The aspect-aware design and implementation of the CiAO operating-system family. Lecture Notes in Computer Science, 7271: 168–215, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35551-6_5/.

Bergmans:2012:FCC

[383] Lodewijk Bergmans, Wilke Havinga, and Mehmet Aksit. First-class compositions. Lecture Notes in Computer Science, 7271:216–267, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35551-6_6/.

Axelsen:2012:CDP

[384] Eyvind W. Axelsen, Fredrik Sørensen, Stein Krogdahl, and Birger Møller-Pedersen. Challenges in the design of the package template mechanism. Lecture Notes in Computer Science, 7271:268–305, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35551-6_7/.

Anonymous:2012:BMj

[385] Anonymous. Back matter. Lecture Notes in Computer Science, 7271: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-35551-6/1.

Anonymous:2012:FMy

[386] Anonymous. Front matter. Lecture Notes in Computer Science, 7271: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-35551-6/1.

Maia:2012:SLM

[387] Francisco Maia and Miguel Matos. Slead: Low-memory, steady distributed systems slicing. Lecture Notes in Computer Science, 7272: 1–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30823-9_1/.

Fouquet:2012:DRP

[388] François Fouquet and Erwan Daubert. Dissemination of reconfiguration policies on mesh networks. Lecture Notes in Computer Science, 7272: 16–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30823-9_2/.

Stamatakis:2012:SRM

[389] Dimokritos Stamatakis and Nikos Tsikoudis. Scalability of replicated metadata services in distributed file systems. Lecture Notes in Computer Science, 7272:31–44, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30823-9_3/.

Rahimian:2012:LAP

[390] Fatemeh Rahimian and Thinh Le Nguyen Huu. Locality-awareness in a peer-to-peer publish/subscribe network. Lecture Notes in Computer Science, 7272:45–58, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30823-9_4/.

DeRyck:2012:SSR

[391] Philippe De Ryck and Nick Nikiforakis. Serene: Self-reliant client-side protection against session fixation. Lecture Notes in Computer Science, 7272:59–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30823-9_5/.

Gunawan:2012:BSC

[392] Linda Ariani Gunawan. Behavioral singletons to consistently handle global states of security patterns. Lecture Notes in Computer Science, 7272:73–86, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30823-9_6/.

Kreutz:2012:TRE

[393] Diego Kreutz and António Casimiro. A trustworthy and resilient event broker for monitoring cloud infrastructures. Lecture Notes in Computer Science, 7272:87–95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30823-9_7/.

Borges:2012:SRE

[394] Miguel Borges, Paulo Jesus, and Carlos Baquero. Spectra: Robust estimation of distribution functions in networks. Lecture Notes in Computer Science, 7272:96–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30823-9_8/.

Maerien:2012:FFA

[395] Jef Maerien and Pieter Agten. FAMoS: a flexible active monitoring service for wireless sensor networks. Lecture Notes in Computer Science, 7272: 104–117, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30823-9_9/.

Benchi:2012:MSO

[396] Abdulkader Benchi and Frédéric Guidec. A message service for opportunistic computing in disconnected MANETs. Lecture Notes in Computer Science, 7272:118–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30823-9_10/.

Bainomugisha:2012:FDS

[397] Engineer Bainomugisha and Koosha Paridel. Flexub: Dynamic subscriptions for publish/subscribe systems in MANETs. Lecture Notes in Computer Science, 7272:132–139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30823-9_11/.

Lyle:2012:DDW

[398] John Lyle, Shamal Faily, and Ivan Fléchais. On the design and development of webinos: a distributed mobile application middleware. Lecture Notes in Computer Science, 7272:140–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30823-9_12/.

Thomson:2012:MPS

[399] Graham Thomson and Sotirios Terzis. A middleware for pervasive situation-awareness. Lecture Notes in Computer Science, 7272:148–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30823-9_13/.

Azab:2012:SUF

[400] Abdulrahman Azab and Hein Meling. Stroll: a universal filesystem-based interface for seamless task deployment in grid computing. Lecture Notes in Computer Science, 7272:162–176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30823-9_14/.

Degerlund:2012:SCI

[401] Fredrik Degerlund. Scheduling of compute-intensive code generated from event-B models: An empirical efficiency study. Lecture Notes in Computer Science, 7272:177–184, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30823-9_15/.

Veeraragavan:2012:RMA

[402] Narasimha Raghavan Veeraragavan. Reliability modeling and analysis of modern distributed interactive multimedia applications: a case study of a distributed opera performance. Lecture Notes in Computer Science, 7272: 185–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30823-9_16/.

Comes:2012:DST

[403] Diana Elena Comes and Christoph Evers. Designing socio-technical applications for ubiquitous computing. Lecture Notes in Computer Science, 7272:194–201, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30823-9_17/.

Voras:2012:SON

[404] Ivan Voras, Marin Orlić, and Mario Žagar. Something old is new again: Reimagining the oldest social networking platform. Lecture Notes in Computer Science, 7272:202–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30823-9_18/.

Rodrigues:2012:ZMS

[405] Preston Rodrigues and Yérom-David Bromberg. ZigZag: a middleware for service discovery in future Internet. Lecture Notes in Computer Science, 7272:208–221, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30823-9_19/.

Comes:2012:BBF

[406] Diana Elena Comes and Harun Baraki. BPRules and the BPR-framework: Comprehensive support for managing QoS in Web service compositions. Lecture Notes in Computer Science, 7272:222–235, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30823-9_20/.

Anonymous:2012:BMk

[407] Anonymous. Back matter. Lecture Notes in Computer Science, 7272: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-30823-9/1.

Anonymous:2012:FMz

[408] Anonymous. Front matter. Lecture Notes in Computer Science, 7272: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30823-9/1.

Lienhardt:2012:RAM

[409] Michael Lienhardt and Ivan Lanese. A reversible abstract machine and its space overhead. Lecture Notes in Computer Science, 7273:1– 17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30793-5_1/.

Johnson:2012:SMT

[410] Taylor T. Johnson and Sayan Mitra. A small model theorem for rectangular hybrid automata networks. Lecture Notes in Computer Science, 7273:18–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30793-5_2/.

Albert:2012:AMH

[411] Elvira Albert and Antonio E. Flores-Montoya. Analysis of may-happen-in-parallel in concurrent objects. Lecture Notes in Computer Science, 7273:35–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30793-5_3/.

Aman:2012:BEM

[412] Bogdan Aman, Gabriel Ciobanu, and Maciej Koutny. Behavioural equivalences over migrating processes with timers. Lecture Notes in Computer Science, 7273:52–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30793-5_4/.

Klai:2012:CSB

[413] Kais Klai and Jörg Desel. Checking soundness of business processes compositionally using symbolic observation graphs. Lecture Notes in Computer Science, 7273:67–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30793-5_5/.

Kindermann:2012:BLC

[414] Roland Kindermann and Tommi Junttila. Beyond lassos: Complete SMTbased bounded model checking for timed automata. Lecture Notes in Computer Science, 7273:84–100, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30793-5_6/.

Prabhakar:2012:CTB

[415] Pavithra Prabhakar and Mahesh Viswanathan. Conformance testing of Boolean programs with multiple faults. Lecture Notes in Computer Science, 7273:101–117, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30793-5_7/.

Bensalem:2012:KBD

[416] Saddek Bensalem, Marius Bozga, and Jean Quilbeuf. Knowledge-based distributed conflict resolution for multiparty interactions and priorities. Lecture Notes in Computer Science, 7273:118–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30793-5_8/.

Cerone:2012:MPW

[417] Andrea Cerone and Matthew Hennessy. Modelling probabilistic wireless networks. Lecture Notes in Computer Science, 7273:135–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30793-5_9/.

Milushev:2012:NSE

[418] Dimiter Milushev, Wim Beck, and Dave Clarke. Noninterference via symbolic execution. Lecture Notes in Computer Science, 7273:152–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30793-5_10/.

Hernandez:2012:DDA

[419] David Romero Hernández and David de Frutos Escrig. Defining distances for all process semantics. Lecture Notes in Computer Science, 7273: 169–185, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30793-5_11/.

Barthe:2012:SME

[420] Gilles Barthe and Juan Manuel Crespo. Secure multi-execution through static program transformation. Lecture Notes in Computer Science, 7273: 186–202, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30793-5_12/.

Delahaye:2012:SIT

[421] Benoît Delahaye and Uli Fahrenberg. Synchronous interface theories and time triggered scheduling. Lecture Notes in Computer Science, 7273: 203–218, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30793-5_13/.

Tasharofi:2012:TND

[422] Samira Tasharofi and Rajesh K. Karmani. TransDPOR: a novel dynamic partial-order reduction technique for testing actor programs. Lecture Notes in Computer Science, 7273:219–234, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30793-5_14/.

Delzanno:2012:VAH

[423] Giorgio Delzanno and Arnaud Sangnier. Verification of ad hoc networks with node and communication failures. Lecture Notes in Computer Science, 7273:235–250, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30793-5_15/.

Earle:2012:VTE

[424] Clara Benac Earle and Lars-Åke Fredlund. Verification of timed Erlang programs using McErlang. Lecture Notes in Computer Science, 7273: 251–267, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30793-5_16/.

Anonymous:2012:BMl

[425] Anonymous. Back matter. Lecture Notes in Computer Science, 7273: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-30793-5/1.

Anonymous:2012:FMba

[426] Anonymous. Front matter. Lecture Notes in Computer Science, 7273: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30793-5/1.

Liptchinsky:2012:SCS

[427] Vitaliy Liptchinsky, Roman Khazankin, and Hong-Linh Truong. Statelets: Coordination of social collaboration processes. Lecture Notes in Computer Science, 7274:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30829-1_1/.

Rossi:2012:SSB

[428] Davide Rossi. A social software-based coordination platform. Lecture Notes in Computer Science, 7274:17–28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30829-1_2/.

AlvaresdeOliveira:2012:SMA

[429] Frederico Alvares de Oliveira Jr. and Remi Sharrock. Synchronization of multiple autonomic control loops: Application to cloud computing. Lecture Notes in Computer Science, 7274:29–43, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30829-1_3/.

vanDooren:2012:STM

[430] Marko van Dooren and Dave Clarke. Subobject transactional memory. Lecture Notes in Computer Science, 7274:44–58, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-30829-1_4/.

Clarke:2012:PCC

[431] Dave Clarke and José Proença. Partial connector colouring. Lecture Notes in Computer Science, 7274:59–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30829-1_5/.

Khosravi:2012:UCA

[432] Ramtin Khosravi and Hamideh Sabouri. Using coordinated actors to model families of distributed systems. Lecture Notes in Computer Science, 7274:74–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30829-1_6/.

Dinges:2012:SSC

[433] Peter Dinges and Gul Agha. Scoped synchronization constraints for large scale actor systems. Lecture Notes in Computer Science, 7274: 89–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30829-1_7/.

Bruni:2012:FOD

[434] Roberto Bruni, Carla Ferreira, and Anne Kersten Kauer. First-order dynamic logic for compensable processes. Lecture Notes in Computer Science, 7274:104–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30829-1_8/.

Aman:2012:CPM

[435] Bogdan Aman and Gabriel Ciobanu. Coordinating parallel mobile ambients to solve SAT problem in polynomial number of steps. Lecture Notes in Computer Science, 7274:122–136, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30829-1_9/.

Terepeta:2012:RAC

[436] Michał Terepeta, Hanne Riis Nielson, and Flemming Nielson. Recursive advice for coordination. Lecture Notes in Computer Science, 7274: 137–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30829-1_10/.

Massink:2012:FAF

[437] Mieke Massink and Diego Latella. Fluid analysis of foraging ants. Lecture Notes in Computer Science, 7274:152–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30829-1_11/.

Dziwok:2012:RTC

[438] Stefan Dziwok, Christian Heinzemann, and Matthias Tichy. Real-time coordination patterns for advanced mechatronic systems. Lecture Notes in Computer Science, 7274:166–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30829-1_12/.

Philips:2012:GOM

[439] Eline Philips and Jorge Vallejos. Group orchestration in a mobile environment. Lecture Notes in Computer Science, 7274:181–195, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30829-1_13/.

Petre:2012:NCP

[440] Luigia Petre, Petter Sandvik, and Kaisa Sere. Node coordination in peer-to-peer networks. Lecture Notes in Computer Science, 7274:196– 211, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30829-1_14/.

Viroli:2012:LST

[441] Mirko Viroli, Danilo Pianini, and Jacob Beal. Linda in space-time: An adaptive coordination model for mobile ad-hoc environments. Lecture Notes in Computer Science, 7274:212–229, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30829-1_15/.

Kuhn:2012:SBG

[442] Eva Kühn, Alexander Marek, and Thomas Scheller. A space-based generic pattern for self-initiative load clustering agents. Lecture Notes in Computer Science, 7274:230–244, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30829-1_16/.

Bartoletti:2012:RCD

[443] Massimo Bartoletti, Emilio Tuosto, and Roberto Zunino. On the realizability of contracts in dishonest systems. Lecture Notes in Computer Science, 7274:245–260, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30829-1_17/.

Degano:2012:TCS

[444] Pierpaolo Degano and Gian-Luigi Ferrari. Types for coordinating secure behavioural variations. Lecture Notes in Computer Science, 7274: 261–276, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30829-1_18/.

Anonymous:2012:BMm

[445] Anonymous. Back matter. Lecture Notes in Computer Science, 7274: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-30829-1/1.

Anonymous:2012:FMbb

[446] Anonymous. Front matter. Lecture Notes in Computer Science, 7274: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30829-1/1.

Hutter:2012:EDS

[447] Michael Hutter, Mario Kirschbaum, Thomas Plos, Jörn-Marc Schmidt, and Stefan Mangard. Exploiting the difference of side-channel leakages. Lecture Notes in Computer Science, 7275:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-29912-4_1/.

Korak:2012:AAE

[448] Thomas Korak, Thomas Plos, and Michael Hutter. Attacking an AESenabled NFC tag: Implications from design to a real-world scenario. Lecture Notes in Computer Science, 7275:17–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29912-4_2/.

Wagner:2012:APS

[449] Mathias Wagner. 700+ attacks published on Smart Cards: The need for a systematic counter strategy. Lecture Notes in Computer Science, 7275:33–38, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29912-4_3/.

He:2012:IEI

[450] Wei He, Eduardo de la Torre, and Teresa Riesgo. An interleaved EPEimmune PA-DPL structure for resisting concentrated EM side channel attacks on FPGA implementation. Lecture Notes in Computer Science, 7275:39–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29912-4_4/.

Mansouri:2012:ACA

[451] Shohreh Sharif Mansouri and Elena Dubrova. An architectural countermeasure against power analysis attacks for FSR-based stream ciphers. Lecture Notes in Computer Science, 7275:54–68, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29912-4_5/.

Coron:2012:CSP

[452] Jean-Sébastien Coron, Christophe Giraud, Emmanuel Prouff, and Soline Renner. Conversion of security proofs from one leakage model to another: a new issue. Lecture Notes in Computer Science, 7275: 69–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29912-4_6/.

Bauer:2012:AEB

[453] Sven Bauer. Attacking exponent blinding in RSA without CRT. Lecture Notes in Computer Science, 7275:82–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29912-4_7/.

DaRolt:2012:NSA

[454] Jean Da Rolt, Amitabh Das, Giorgio Di Natale, Marie-Lise Flottes, and Bruno Rouzeyre. A new scan attack on RSA in presence of industrial countermeasures. Lecture Notes in Computer Science, 7275: 89–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29912-4_8/.

Vuillaume:2012:RKG

[455] Camille Vuillaume, Takashi Endo, and Paul Wooderson. RSA key generation: New attacks. Lecture Notes in Computer Science, 7275: 105–119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29912-4_9/.

Jovanovic:2012:FAL

[456] Philipp Jovanovic, Martin Kreuzer, and Ilia Polian. A fault attack on the LED Block cipher. Lecture Notes in Computer Science, 7275: 120–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29912-4_10/.

Zhao:2012:DFA

[457] Liang Zhao, Takashi Nishide, and Kouichi Sakurai. Differential fault analysis of full LBlock. Lecture Notes in Computer Science, 7275: 135–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29912-4_11/.

Bayon:2012:CEA

[458] Pierre Bayon, Lilian Bossuet, Alain Aubert, Viktor Fischer, and François Poucheret. Contactless electromagnetic active attack on ring oscillator based true random number generator. Lecture Notes in Computer Science, 7275:151–166, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29912-4_12/.

Fischer:2012:CLS

[459] Viktor Fischer. A closer look at security in random number generators design. Lecture Notes in Computer Science, 7275:167–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29912-4_13/.

Murdica:2012:SVP

[460] Cédric Murdica, Sylvain Guilley, Jean-Luc Danger, Philippe Hoogvorst, and David Naccache. Same values power analysis using special points on elliptic curves. Lecture Notes in Computer Science, 7275:183–198, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29912-4_14/.

Kruger:2012:SIA

[461] Alexander Krüger. The Schindler–Itoh-attack in case of partial information leakage. Lecture Notes in Computer Science, 7275:199–214, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29912-4_15/.

Zohner:2012:BAS

[462] Michael Zohner, Michael Kasper, and Marc Stöttinger. Butterfly-attack on Skein’s modular addition. Lecture Notes in Computer Science, 7275: 215–230, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29912-4_16/.

Zhao:2012:MEA

[463] Xinjie Zhao, Fan Zhang, Shize Guo, Tao Wang, Zhijie Shi, Huiying Liu, and Keke Ji. MDASCA: An enhanced algebraic side-channel attack for error tolerance and new leakage model exploitation. Lecture Notes in Computer Science, 7275:231–248, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29912-4_17/.

Heuser:2012:IMH

[464] Annelie Heuser and Michael Zohner. Intelligent machine homicide. Lecture Notes in Computer Science, 7275:249–264, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29912-4_18/.

Anonymous:2012:BMn

[465] Anonymous. Back matter. Lecture Notes in Computer Science, 7275: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-29912-4/1.

Anonymous:2012:FMbc

[466] Anonymous. Front matter. Lecture Notes in Computer Science, 7275: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29912-4/1.

Lubbecke:2012:ADB

[467] Marco E. Lübbecke. Automatic decomposition and branch-and-price— a status report. Lecture Notes in Computer Science, 7276:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30850-5_1/.

Kempkes:2012:CLS

[468] Barbara Kempkes and Friedhelm Meyer auf der Heide. Continuous local strategies for robotic formation problems. Lecture Notes in Computer Science, 7276:9–17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30850-5_2/.

Osipov:2012:EGP

[469] Vitaly Osipov, Peter Sanders, and Christian Schulz. Engineering graph partitioning algorithms. Lecture Notes in Computer Science, 7276:18– 26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30850-5_3/.

Albrecht:2012:SEM

[470] Benjamin Albrecht and Volker Heun. Space efficient modifications to structator— a fast index-based search tool for RNA sequence-structure patterns. Lecture Notes in Computer Science, 7276:27–38, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30850-5_4/.

Angel:2012:ICH

[471] Eric Angel, Romain Campigotto, and Christian Laforest. Implementation and comparison of heuristics for the vertex cover problem on huge graphs. Lecture Notes in Computer Science, 7276:39–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30850-5_5/.

Berger:2012:HAN

[472] Annabell Berger and Matthias Müller-Hannemann. How to attack the NP-Complete dag realization problem in practice. Lecture Notes in Computer Science, 7276:51–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30850-5_6/.

Busing:2012:NRA

[473] Christina Büsing and Fabio D’Andreagiovanni. New results about multiband uncertainty in robust optimization. Lecture Notes in Computer Science, 7276:63–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30850-5_7/.

Cafieri:2012:CRP

[474] Sonia Cafieri, Pierre Hansen, Lucas Létocart, Leo Liberti, and Frédéric Messine. Compact relaxations for polynomial programming problems. Lecture Notes in Computer Science, 7276:75–86, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30850-5_8/.

Costa:2012:RMC

[475] Alberto Costa and Leo Liberti. Relaxations of multilinear convex envelopes: Dual is better than primal. Lecture Notes in Computer Science, 7276:87–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30850-5_9/.

Crescenzi:2012:CDR

[476] Pierluigi Crescenzi, Roberto Grossi, Leonardo Lanzi, and Andrea Marino. On computing the diameter of real-world directed (weighted) graphs. Lecture Notes in Computer Science, 7276:99–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30850-5_10/.

DAndrea:2012:RSM

[477] Annalisa D’Andrea and Guido Proietti. Reoptimizing the strengthened metric TSP on multiple edge weight modifications. Lecture Notes in Computer Science, 7276:111–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30850-5_11/.

DAngelo:2012:ENL

[478] Gianlorenzo D’Angelo, Mattia D’Emidio, Daniele Frigioni, and Vinicio Maurizio. Engineering a new loop-free shortest paths routing algorithm. Lecture Notes in Computer Science, 7276:123–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30850-5_12/.

DAngelo:2012:FDM

[479] Gianlorenzo D’Angelo, Mattia D’Emidio, Daniele Frigioni, and Camillo Vitale. Fully dynamic maintenance of arc-flags in road networks. Lecture Notes in Computer Science, 7276:135–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30850-5_13/.

Dietzfelbinger:2012:MRG

[480] Martin Dietzfelbinger, Hendrik Peilke, and Michael Rink. A more reliable greedy heuristic for maximum matchings in sparse random graphs. Lecture Notes in Computer Science, 7276:148–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30850-5_14/.

Elmasry:2012:BMD

[481] Amr Elmasry, Jyrki Katajainen, and Max Stenmark. Branch mispredictions don’t affect Mergesort. Lecture Notes in Computer Science, 7276: 160–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30850-5_15/.

Faro:2012:MSW

[482] Simone Faro and Thierry Lecroq. A multiple sliding Windows approach to speed up string matching algorithms. Lecture Notes in Computer Science, 7276:172–183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30850-5_16/.

Fertin:2012:ASM

[483] Guillaume Fertin, Hafedh Mohamed Babou, and Irena Rusu. Algorithms for subnetwork mining in heterogeneous networks. Lecture Notes in Computer Science, 7276:184–194, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30850-5_17/.

Firmani:2012:CSA

[484] Donatella Firmani, Giuseppe F. Italiano, Luigi Laura, and Alessio Orlandi. Computing strong articulation points and strong bridges in large scale graphs. Lecture Notes in Computer Science, 7276:195–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30850-5_18/.

Georgiadis:2012:ADM

[485] Giorgos Georgiadis and Marina Papatriantafilou. Adaptive distributed bmatching in overlays with preferences. Lecture Notes in Computer Science, 7276:208–223, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30850-5_19/.

Joannou:2012:DST

[486] Stelios Joannou and Rajeev Raman. Dynamizing succinct tree representations. Lecture Notes in Computer Science, 7276:224–235, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30850-5_20/.

Anonymous:2012:FMbd

[487] Anonymous. Front matter. Lecture Notes in Computer Science, 7276: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30850-5/1.

Masti:2012:VVN

[488] Sarang Bharadwaj Masti, Siva P. Meenakshi, and Serugudi V. Raghavan. VNEMX: Virtual network embedding test-bed using MPLS and Xen. Lecture Notes in Computer Science, 7277:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-30630-3_1/.

Papadimitriou:2012:TLS

[489] Panagiotis Papadimitriou, Ines Houidi, Wajdi Louati, Djamal Zeghlache, and Christoph Werle. Towards large-scale network virtualization. Lecture Notes in Computer Science, 7277:13–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30630-3_2/.

Lakafosis:2012:PWI

[490] Vasileios Lakafosis, Sreenivas Addagatla, Christian Belady, and Suyash Sinha. Prometheus: a wirelessly interconnected, pico-datacenter framework for the developing world. Lecture Notes in Computer Science, 7277: 26–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30630-3_3/.

Galinina:2012:PAC

[491] Olga Galinina, Sergey Andreev, and Yevgeni Koucheryavy. Performance analysis of client relay cloud in wireless cellular networks. Lecture Notes in Computer Science, 7277:40–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30630-3_4/.

Liaskos:2012:PSC

[492] Christos Liaskos, Andreas Xeros, Georgios I. Papadimitriou, and Marios Lestas. Periodic scheduling with costs revisited. Lecture Notes in Computer Science, 7277:52–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30630-3_5/.

Liaskos:2012:ML

[493] Christos Liaskos, Ageliki Tsioliaridou, and Georgios I. Papadimitriou. More for less. Lecture Notes in Computer Science, 7277:64–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30630-3_6/.

Lai:2012:MIC

[494] Junyu Lai and Bernd E. Wolfinger. A method to improve the channel availability of IPTV systems with users zapping channels sequentially. Lecture Notes in Computer Science, 7277:76–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30630-3_7/.

Ababneh:2012:ABA

[495] Nedal Ababneh, Nicholas Timmons, and Jim Morrison. An adaptive bandwidth allocation scheme for data streaming over body area networks. Lecture Notes in Computer Science, 7277:90–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30630-3_8/.

Ahvar:2012:ENE

[496] Ehsan Ahvar, René Serral-Gracià, Eva Marín-Tordera, and Xavier Masip-Bruin. EQR: a new energy-aware query-based routing protocol for wireless sensor networks. Lecture Notes in Computer Science, 7277: 102–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30630-3_9/.

Wagenknecht:2012:PER

[497] Gerald Wagenknecht, Markus Anwander, and Torsten Braun. Performance evaluation of reliable overlay multicast in wireless sensor networks. Lecture Notes in Computer Science, 7277:114–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30630-3_10/.

Dimitrova:2012:ECB

[498] Desislava C. Dimitrova, Islam Alyafawi, and Torsten Braun. Experimental comparison of Bluetooth and WiFi signal propagation for indoor localisation. Lecture Notes in Computer Science, 7277:126–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30630-3_11/.

Mondal:2012:LPM

[499] Kaushik Mondal, Partha Sarathi Mandal, and Bhabani P. Sinha. Localization in presence of multipath effect in wireless sensor networks. Lecture Notes in Computer Science, 7277:138–149, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30630-3_12/.

Yousef:2012:RRH

[500] Ausama Yousef, Samer Rishe, Andreas Mitschele-Thiel, and Abdalkarim Awad. RNBB: a reliable hybrid broadcasting algorithm for ad-hoc networks. Lecture Notes in Computer Science, 7277:150–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30630-3_13/.

Hosseinabadi:2012:EOO

[501] Ghazale Hosseinabadi and Nitin H. Vaidya. Exploiting opportunistic overhearing to improve performance of mutual exclusion in wireless ad hoc networks. Lecture Notes in Computer Science, 7277:162–173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30630-3_14/.

Cadger:2012:MLP

[502] Fraser Cadger, Kevin Curran, Jose Santos, and Sandra Moffett. MANET location prediction using machine learning algorithms. Lecture Notes in Computer Science, 7277:174–185, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30630-3_15/.

Marinho:2012:CSB

[503] José Marinho and Edmundo Monteiro. Cooperative sensing-before-transmit in ad-hoc multi-hop cognitive radio scenarios. Lecture Notes in Computer Science, 7277:186–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30630-3_16/.

Oredope:2012:EVH

[504] Adetola Oredope, Guilherme Frassetto, and Barry Evans. An evaluation of vertical handovers in LTE networks. Lecture Notes in Computer Science, 7277:198–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30630-3_17/.

Vondra:2012:CCB

[505] Michal Vondra and Zdenek Becvar. Connection cost based handover decision for offloading macrocells by femtocells. Lecture Notes in Computer Science, 7277:208–219, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30630-3_18/.

Hsu:2012:DLA

[506] Chung-Hsien Hsu. Direct link assignment approach for IEEE 802.16 networks. Lecture Notes in Computer Science, 7277:220–231, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30630-3_19/.

Yu:2012:DCR

[507] Dongxiao Yu, Qiang-Sheng Hua, Weiguo Dai, Yuexuan Wang, and Francis C. M. Lau. Dynamic contention resolution in multiple-access channels. Lecture Notes in Computer Science, 7277:232–243, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30630-3_20/.

Anonymous:2012:FMbe

[508] Anonymous. Front matter. Lecture Notes in Computer Science, 7277: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30630-3/1.

Chen:2012:DWE

[509] Hsinchun Chen. Dark web: Exploring and mining the dark side of the Web. Lecture Notes in Computer Science, 7278:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-29892-9_1.

DeRaedt:2012:DMM

[510] Luc De Raedt. Declarative modeling for machine learning and data mining. Lecture Notes in Computer Science, 7278:2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/accesspage/chapter/10.1007/978-3-642-29892-9_2.

Elzinga:2012:CCR

[511] Paul Elzinga. Can concepts reveal criminals? Lecture Notes in Computer Science, 7278:3, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-29892-9_3.

Goethals:2012:CSI

[512] Bart Goethals. Cartification: From similarities to itemset frequencies. Lecture Notes in Computer Science, 7278:4, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/accesspage/chapter/10.1007/978-3-642-29892-9_4.

Peters:2012:PCA

[513] Ir. Edward Peters. Processes are concepts, aren’t they? Lecture Notes in Computer Science, 7278:5, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-29892-9_5.

Slezak:2012:RSF

[514] Dominik Ślezak. Rough sets and FCA — scalability challenges. Lecture Notes in Computer Science, 7278:6, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-29892-9_6.

Babin:2012:ACS

[515] Mikhail A. Babin and Sergei O. Kuznetsov. Approximating concept stability. Lecture Notes in Computer Science, 7278:7–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29892-9_7/.

Bartl:2012:LAC

[516] Eduard Bartl and Michal Krupka. Logical analysis of concept lattices by factorization. Lecture Notes in Computer Science, 7278:16– 27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29892-9_8/.

Belohlavek:2012:BLC

[517] Radim Belohlavek and Martin Trnecka. Basic level of concepts in formal concept analysis. Lecture Notes in Computer Science, 7278:28– 44, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29892-9_9/.

Berry:2012:PTL

[518] Anne Berry and Alain Sigayret. A peep through the looking glass: Articulation points in lattices. Lecture Notes in Computer Science, 7278: 45–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29892-9_10/.

Chollet:2012:PUF

[519] Stéphanie Chollet, Vincent Lestideau, Yoann Maurel, Etienne Gandrille, and Philippe Lalanda. Practical use of formal concept analysis in service-oriented computing. Lecture Notes in Computer Science, 7278: 61–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29892-9_11/.

Doerfel:2012:PAF

[520] Stephan Doerfel, Robert Jäschke, and Gerd Stumme. Publication analysis of the formal concept analysis community. Lecture Notes in Computer Science, 7278:77–95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29892-9_12/.

Endres:2012:USS

[521] Dominik Endres, Ruth Adam, Martin A. Giese, and Uta Noppeney. Understanding the semantic structure of human fMRI brain recordings with formal concept analysis. Lecture Notes in Computer Science, 7278: 96–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29892-9_13/.

Ferre:2012:CCM

[522] Sébastien Ferré, Pierre Allard, and Olivier Ridoux. Cubes of concepts: Multi-dimensional exploration of multi-valued contexts. Lecture Notes in Computer Science, 7278:112–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29892-9_14/.

Ganter:2012:OFA

[523] Bernhard Ganter and Cynthia Vera Glodeanu. Ordinal factor analysis. Lecture Notes in Computer Science, 7278:128–139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29892-9_15/.

Kaiser:2012:MAF

[524] Tim B. Kaiser and Stefan E. Schmidt. A macroscopic approach to FCA and its various fuzzifications. Lecture Notes in Computer Science, 7278: 140–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29892-9_16/.

Kerkhoff:2012:CBC

[525] Sebastian Kerkhoff. A connection between clone theory and FCA provided by duality theory. Lecture Notes in Computer Science, 7278:

148–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29892-9_17/.

Kirchberg:2012:FCD

[526] Markus Kirchberg, Erwin Leonardi, Yu Shyang Tan, Sebastian Link, and Ryan K. L. Ko. Formal concept discovery in semantic Web data. Lecture Notes in Computer Science, 7278:164–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29892-9_18/.

Krupka:2012:CLI

[527] Michal Krupka and Jan Lastovicka. Concept lattices of incomplete data. Lecture Notes in Computer Science, 7278:180–194, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29892-9_19/.

Macko:2012:FCA

[528] Juraj Macko. Formal concept analysis as a framework for business intelligence technologies. Lecture Notes in Computer Science, 7278:195– 210, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29892-9_20/.

Anonymous:2012:FMbf

[529] Anonymous. Front matter. Lecture Notes in Computer Science, 7278: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29892-9/1.

Leita:2012:CCI

[530] Corrado Leita. Challenges in critical infrastructure security. Lecture Notes in Computer Science, 7279:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-30633-4_1.

Bouten:2012:ODD

[531] Niels Bouten, Anna Hristoskova, Femke Ongenae, Jelle Nelis, and Filip De Turck. Ontology-driven dynamic discovery and distributed coordination of a robot swarm. Lecture Notes in Computer Science, 7279: 2–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30633-4_2/.

Vancea:2012:CDC

[532] Andrei Vancea, Guilherme Sperb Machado, Laurent d’Orazio, and Burkhard Stiller. Cooperative database caching within cloud environments. Lecture Notes in Computer Science, 7279:14–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30633-4_3/.

Georgoulas:2012:FRL

[533] Stylianos Georgoulas, Klaus Moessner, Alexis Mansour, and Menelaos Pissarides. A fuzzy reinforcement learning approach for pre-congestion notification based admission control. Lecture Notes in Computer Science, 7279:26–37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30633-4_4/.

Ruckert:2012:BGT

[534] Julius Rückert and David Hausheer. Bridging the gap: Towards an adaptive video streaming approach supporting transitions. Lecture Notes in Computer Science, 7279:38–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30633-4_5/.

Rautio:2012:MEL

[535] Teemu Rautio and Jukka Mäkelä. A multiaccess enabled load balancing in cognitive networks. Lecture Notes in Computer Science, 7279: 42–45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30633-4_6/.

Pus:2012:HAM

[536] Viktor Puš. Hardware acceleration for measurements in 100 Gb/s networks. Lecture Notes in Computer Science, 7279:46–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30633-4_7/.

Korte:2012:SRR

[537] Kevin Dominik Korte, Anuj Sehgal, and Jürgen Schönwälder. A study of the RPL repair process using ContikiRPL. Lecture Notes in Computer Science, 7279:50–61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30633-4_8/.

Krejci:2012:TMA

[538] Radek Krejčí, Pavel Čeleda, and Jakub Dobrovolný. Traffic measurement and analysis of building automation and control networks. Lecture Notes in Computer Science, 7279:62–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30633-4_9/.

Jung:2012:SSU

[539] Tobias Jung, Sylvain Martin, Damien Ernst, and Guy Leduc. SPRT for SPIT: Using the sequential probability ratio test for spam in VoIP prevention. Lecture Notes in Computer Science, 7279:74–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30633-4_10/.

Hellemons:2012:SFB

[540] Laurens Hellemons, Luuk Hendriks, Rick Hofstede, Anna Sperotto, Ramin Sadre, and Aiko Pras. SSHCure: a flow-based SSH intrusion detection system. Lecture Notes in Computer Science, 7279:86–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30633-4_11/.

Ries:2012:IDA

[541] Thorsten Ries, Radu State, and Thomas Engel. Instant degradation of anonymity in low-latency anonymisation systems. Lecture Notes in Computer Science, 7279:98–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30633-4_12/.

Hofstede:2012:RTR

[542] Rick Hofstede and Aiko Pras. Real-time and resilient intrusion detection: a flow-based approach. Lecture Notes in Computer Science, 7279: 109–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30633-4_13/.

Bartos:2012:DSO

[543] Karel Bartos and Martin Rehak. Distributed self-organized collaboration of autonomous IDS sensors. Lecture Notes in Computer Science, 7279: 113–117, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30633-4_14/.

Bartos:2012:NAD

[544] Václav Bartoš and Martin Žádník. Network anomaly detection: Comparison and real-time issues. Lecture Notes in Computer Science, 7279: 118–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30633-4_15/.

Wagner:2012:DAN

[545] Cynthia Wagner and Thomas Engel. Detecting anomalies in Netflow record time series by using a kernel function. Lecture Notes in Computer Science, 7279:122–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30633-4_16/.

Tobola:2012:HBI

[546] Jiří Tobola. How to build an IT spin-off company. Lecture Notes in Computer Science, 7279:126, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-30633-4_17.

Schmidt:2012:TBE

[547] Ricardo de O. Schmidt, Anna Sperotto, Ramin Sadre, and Aiko Pras. Towards bandwidth estimation using flow-level measurements. Lecture Notes in Computer Science, 7279:127–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30633-4_18/.

Bajpai:2012:FBI

[548] Vaibhav Bajpai, Nikolay Melnikov, Anuj Sehgal, and Jürgen Schönwälder. Flow-based identification of failures caused by IPv6 transition mechanisms. Lecture Notes in Computer Science, 7279:139–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30633-4_19/.

Marchal:2012:LSD

[549] Samuel Marchal and Thomas Engel. Large scale DNS analysis. Lecture Notes in Computer Science, 7279:151–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30633-4_20/.

Anonymous:2012:BMo

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-30633-4/1.

Anonymous:2012:FMbg

[551] Anonymous. Front matter. Lecture Notes in Computer Science, 7279: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30633-4/1.

Yang:2012:OPS

[552] Yang Yang, Guang Gong, and Xiaohu Tang. Odd perfect sequences and sets of spreading sequences with zero or low odd periodic correlation zone. Lecture Notes in Computer Science, 7280:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30615-0_1/.

Ozbudak:2012:NCA

[553] Ferruh Özbudak, Oğuz Yayla, and C. Cengiz Yıldırım. Nonexistence of certain almost p-ary perfect sequences. Lecture Notes in Computer Science, 7280:13–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30615-0_2/.

Tan:2012:NFD

[554] Yin Tan, Longjiang Qu, Chik How Tan, and Chao Li. New families of differentially 4-uniform permutations over f_{2^{2k}}. Lecture Notes in Computer Science, 7280:25–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30615-0_3/.

Flori:2012:DPH

[555] Jean-Pierre Flori and Sihem Mesnager. Dickson polynomials, hyperelliptic curves and hyper-bent functions. Lecture Notes in Computer Science, 7280:40–52, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30615-0_4/.

Lutz:2012:VWS

[556] Jonathan Lutz, Charles J. Colbourn, and Violet R. Syrotiuk. Variable weight sequences for adaptive scheduled access in MANETs. Lecture Notes in Computer Science, 7280:53–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30615-0_5/.

Klapper:2012:AWT

[557] Andrew Klapper. Arithmetic Walsh transform of quadratic Boolean functions. Lecture Notes in Computer Science, 7280:65–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30615-0_6/.

Sarkar:2012:CNB

[558] Sumanta Sarkar. Characterizing negabent Boolean functions over finite fields. Lecture Notes in Computer Science, 7280:77–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30615-0_7/.

Calik:2012:CWB

[559] Çağdaş Çalık and Ali Doğanaksoy. Computing the weight of a Boolean function from its algebraic normal form. Lecture Notes in Computer Science, 7280:89–100, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30615-0_8/.

Pirsic:2012:BFD

[560] Gottlieb Pirsic and Arne Winterhof. Boolean functions derived from pseudorandom binary sequences. Lecture Notes in Computer Science, 7280: 101–109, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30615-0_9/.

Golomb:2012:ISF

[561] Solomon W. Golomb. Infinite sequences with finite cross-correlation-II. Lecture Notes in Computer Science, 7280:110–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30615-0_10/.

Dorsey:2012:ICR

[562] Thomas J. Dorsey and Alfred W. Hales. Irreducible coefficient relations. Lecture Notes in Computer Science, 7280:117–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30615-0_11/.

Jedwab:2012:WIS

[563] Jonathan Jedwab and Jane Wodlinger. Wavelength isolation sequence pairs. Lecture Notes in Computer Science, 7280:126–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30615-0_12/.

Salagean:2012:ITF

[564] Ana Sălăgean, David Gardner, and Raphael Phan. Index tables of finite fields and modular Golomb rulers. Lecture Notes in Computer Science, 7280:136–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30615-0_13/.

Zhou:2012:AHC

[565] Zhengchun Zhou, Xiaohu Tang, Yang Yang, and Udaya Parampalli. On the aperiodic Hamming correlation of frequency-hopping sequences from norm functions. Lecture Notes in Computer Science, 7280:148– 158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30615-0_14/.

Acevedo:2012:PSU

[566] Santiago Barrera Acevedo and Thomas E. Hall. Perfect sequences of unbounded lengths over the basic quaternions. Lecture Notes in Computer Science, 7280:159–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30615-0_15/.

Vielhaber:2012:LCD

[567] Michael Vielhaber and Mónica del Pilar Canales Chacón. The linear complexity deviation of multisequences: Formulae for finite lengths and asymptotic distributions. Lecture Notes in Computer Science, 7280: 168–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30615-0_16/.

Chen:2012:LCB

[568] Zhixiong Chen and Domingo Gómez-Pérez. Linear complexity of binary sequences derived from polynomial quotients. Lecture Notes in Computer Science, 7280:181–189, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30615-0_17/.

UlHasan:2012:WOT

[569] Sartaj Ul Hasan, Daniel Panario, and Qiang Wang. Word-oriented transformation shift registers and their linear complexity. Lecture Notes in Computer Science, 7280:190–201, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30615-0_18/.

Chung:2012:LHZ

[570] Jin-Ho Chung and Kyeongcheol Yang. Low-hit-zone frequency-hopping sequence sets with new parameters. Lecture Notes in Computer Science, 7280:202–211, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30615-0_19/.

Moreno:2012:NOL

[571] Oscar Moreno and Andrew Tirkel. New optimal low correlation sequences for wireless communications. Lecture Notes in Computer Science, 7280: 212–223, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30615-0_20/.

Anonymous:2012:FMbh

[572] Anonymous. Front matter. Lecture Notes in Computer Science, 7280: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30615-0/1.

Wainwright:2012:IFR

[573] Nick Wainwright and Nick Papanikolaou. Introduction: The FIA research roadmap, priorities for future Internet research. Lecture Notes in Computer Science, 7281:1–5, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/content/pdf/
10.1007/978-3-642-30241-1_1.

Kostopoulos:2012:TAI

[574] Alexandros Kostopoulos, Ioanna Papafili, Costas Kalogiros, Tapio Levä, and Nan Zhang. A tussle analysis for information-centric networking architectures. Lecture Notes in Computer Science, 7281:6–17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/content/pdf/10.1007/978-3-642-30241-1_2.

Sales:2012:SAC

[575] Bernard Sales, Emmanuel Darmois, Dimitri Papadimitriou, and Didier Bourse. A systematic approach for closing the research to standardization gap. Lecture Notes in Computer Science, 7281:18–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/content/pdf/10.1007/978-3-642-30241-1_3.

Doolin:2012:SWP

[576] Kevin Doolin, Ioanna Roussaki, Mark Roddy, Nikos Kalatzis, and Elizabeth Papadopoulou. SOCIETIES: Where pervasive meets social. Lecture Notes in Computer Science, 7281:30–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-30241-1_4.

Oostveen:2012:CDL

[577] Anne-Marie Oostveen, Isis Hjorth, Brian Pickering, Michael Boniface, and Eric T. Meyer. Cross-disciplinary lessons for the future Internet. Lecture Notes in Computer Science, 7281:42–54, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-30241-1_5.

Papadimitriou:2012:DPF

[578] Dimitri Papadimitriou, Theodore Zahariadis, Pedro Martinez-Julia, and Ioanna Papafili. Design principles for the future Internet architecture. Lecture Notes in Computer Science, 7281:55–67, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-30241-1_6.

Papadimitriou:2012:IAR

[579] Dimitri Papadimitriou, Bernard Sales, Piet Demeester, and Theodore Zahariadis. From Internet architecture research to standards. Lecture Notes in Computer Science, 7281:68–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-30241-1_7.

BenHamida:2012:IDR

[580] Amira Ben Hamida, Fabio Kon, Gustavo Ansaldi Oliva, and Carlos Eduardo Moreira Dos Santos. An integrated development and runtime environment for the future Internet. Lecture Notes in Computer Science, 7281:81–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/content/pdf/10.1007/
978-3-642-30241-1_8.

Davey:2012:VAT

[581] James Davey, Florian Mansmann, Jörn Kohlhammer, and Daniel Keim. Visual analytics: Towards intelligent interactive Internet and security solutions. Lecture Notes in Computer Science, 7281:93–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/content/pdf/10.1007/978-3-642-30241-1_9.

DiCerbo:2012:TTS

[582] Francesco Di Cerbo, Michele Bezzi, Samuel Paul Kaluvuri, and Antonino Sabetta. Towards a trustworthy service marketplace for the future Internet. Lecture Notes in Computer Science, 7281:105–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/content/pdf/10.1007/978-3-642-30241-1_10.

Antoniou:2012:UFI

[583] Athanasios Antoniou, Evangelos Theodoridis, Ioannis Chatzigiannakis, and Georgios Mylonas. Using future Internet infrastructure and Smartphones for mobility trace acquisition and social interactions monitoring. Lecture Notes in Computer Science, 7281:117–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/content/pdf/10.1007/978-3-642-30241-1_11.

Axenopoulos:2012:SUF

[584] Apostolos Axenopoulos, Petros Daras, Sotiris Malassiotis, and Vincenzo Croce. I-SEARCH: a unified framework for multimodal search and retrieval. Lecture Notes in Computer Science, 7281:130–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/content/pdf/10.1007/978-3-642-30241-1_12.

deOliveiraSilva:2012:SES

[585] Flávio de Oliveira Silva, Alex Dias, Caio César Ferreira, and Eduardo De Souza Santos. Semantically enriched services to understand the need of entities. Lecture Notes in Computer Science, 7281:142–153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/content/pdf/10.1007/978-3-642-30241-1_13.

Swiatek:2012:SCC

[586] Paweł Światek, Krzysztof Juszczyszyn, Krzysztof Brzostowski, and Jarosław Drapała. Supporting content, context and user awareness in future Internet applications. Lecture Notes in Computer Science, 7281: 154–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/content/pdf/10.1007/
978-3-642-30241-1_14.

Srivastava:2012:TNA

[587] Lara Srivastava and Athena Vakali. Towards a narrative-aware design framework for smart urban environments. Lecture Notes in Computer Science, 7281:166–177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/content/pdf/
10.1007/978-3-642-30241-1_15.

Anthopoulos:2012:UPS

[588] Leonidas G. Anthopoulos and Athena Vakali. Urban planning and smart cities: Interrelations and reciprocities. Lecture Notes in Computer Science, 7281:178–189, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-30241-1_16.

Gimenez:2012:STF

[589] Roberto Gimenez, Diego Fuentes, Emilio Martin, Diego Gimenez, and Judith Pertejo. The safety transformation in the future Internet domain. Lecture Notes in Computer Science, 7281:190–200, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/content/pdf/10.1007/978-3-642-30241-1_17.

Tranoris:2012:FAS

[590] Christos Tranoris and Spyros Denazis. FSToolkit: Adopting software engineering practices for enabling definitions of federated resource infrastructures. Lecture Notes in Computer Science, 7281:201–212, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/content/pdf/10.1007/978-3-642-30241-1_18.

Lymberopoulos:2012:NTA

[591] Leonidas Lymberopoulos, Mary Grammatikou, Martin Potts, Paola Grosso, and Attila Fekete. NOVI tools and algorithms for federating virtualized infrastructures. Lecture Notes in Computer Science, 7281: 213–224, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/content/pdf/10.1007/
978-3-642-30241-1_19.

Tomkos:2012:NGF

[592] Ioannis Tomkos, Marianna Angelou, Ramón J. Durán Barroso, and Ignacio de Miguel. Next generation flexible and cognitive heterogeneous optical networks. Lecture Notes in Computer Science, 7281:225–236, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/978-
3-642-30241-1_20.

Anonymous:2012:BMp

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-30241-1/1.

Anonymous:2012:FMbi

[594] Anonymous. Front matter. Lecture Notes in Computer Science, 7281: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30241-1/1.

Argelaguet:2012:NAP

[595] Ferran Argelaguet, David Antonio Gómez Jáuregui, Maud Marchal, and Anatole Lécuyer. A novel approach for pseudo-haptic textures based on curvature information. Lecture Notes in Computer Science, 7282: 1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31401-8_1/.

Asque:2012:CNU

[596] Christopher T. Asque, Andy M. Day, and Stephen D. Laycock. Cursor navigation using haptics for motion-impaired computer users. Lecture Notes in Computer Science, 7282:13–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31401-8_2/.

Ban:2012:MIA

[597] Yuki Ban, Takashi Kajinami, Takuji Narumi, Tomohiro Tanikawa, and Michitaka Hirose. Modifying an identified angle of edged shapes using pseudo-haptic effects. Lecture Notes in Computer Science, 7282: 25–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31401-8_3/.

Baser:2012:TIH

[598] Ozgur Baser, E. Ilhan Konukseven, and Hakan Gurocak. Transparency improvement in haptic devices with a torque compensator using motor current. Lecture Notes in Computer Science, 7282:37–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31401-8_4/.

Baylan:2012:FEM

[599] Buket Baylan, Ugur Aridogan, and Cagatay Basdogan. Finite element modeling of a vibrating touch screen actuated by piezo patches for haptic feedback. Lecture Notes in Computer Science, 7282:47–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31401-8_5/.

Bellan:2012:EVE

[600] Valeria Bellan, Carlo Reverberi, and Alberto Gallace. Evidence for ‘visual enhancement of touch’ mediated by visual displays and its relationship with body ownership. Lecture Notes in Computer Science, 7282: 58–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31401-8_6/.

Brandi:2012:PAI

[601] Fernanda Brandi, Burak Cizmeci, and Eckehard Steinbach. On the perceptual artifacts introduced by packet losses on the forward channel of haptic telemanipulation sessions. Lecture Notes in Computer Science, 7282:67–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31401-8_7/.

Changeon:2012:TEV

[602] Gwénaël Changeon, Delphine Graeff, Margarita Anastassova, and José Lozada. Tactile emotions: a vibrotactile tactile gamepad for transmitting emotional messages to children with autism. Lecture Notes in Computer Science, 7282:79–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31401-8_8/.

VanDoorn:2012:CLC

[603] George H. Van Doorn, Vladimir Dubaj, Dianne B. Wuillemin, and Barry L. Richardson. Cognitive load can explain differences in active and passive touch. Lecture Notes in Computer Science, 7282:91–102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31401-8_9/.

Endo:2012:HBP

[604] Satoshi Endo, Geoff Pegman, Mark Burgin, Tarek Toumi, and Alan M. Wing. Haptics in between-person object transfer. Lecture Notes in Computer Science, 7282:103–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31401-8_10/.

Galiana:2012:NCA

[605] Ignacio Galiana, Jose Breñosa, Jorge Barrio, and Manuel Ferre. New control architecture based on PXI for a 3-finger haptic device applied to virtual manipulation. Lecture Notes in Computer Science, 7282: 112–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31401-8_11/.

George:2012:CBC

[606] Laurent George, Maud Marchal, Loeiz Glondu, and Anatole Lécuyer. Combining brain-computer interfaces and haptics: Detecting mental workload to adapt haptic assistance. Lecture Notes in Computer Science, 7282:124–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31401-8_12/.

Giachritsis:2012:DIT

[607] Christos Giachritsis, Gary Randall, and Samuel Roselier. Development of intuitive tactile navigational patterns. Lecture Notes in Computer Science, 7282:136–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31401-8_13/.

Greci:2012:FHD

[608] Luca Greci, Marco Sacco, Nicola Cau, and Flavia Buonanno. FootGlove: a haptic device supporting the customer in the choice of the best fitting shoes. Lecture Notes in Computer Science, 7282:148–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31401-8_14/.

Gurari:2012:DSV

[609] Netta Gurari, Jason Wheeler, Amy Shelton, and Allison M. Okamura. Discrimination of springs with vision, proprioception, and artificial skin stretch cues. Lecture Notes in Computer Science, 7282: 160–172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31401-8_15/.

Hachisu:2012:AMP

[610] Taku Hachisu, Michi Sato, Shogo Fukushima, and Hiroyuki Kajimoto. Augmentation of material property by modulating vibration resulting from tapping. Lecture Notes in Computer Science, 7282:173–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31401-8_16/.

Hamza-Lup:2012:FSK

[611] Felix G. Hamza-Lup and William H. Baird. Feel the static and kinetic friction. Lecture Notes in Computer Science, 7282:181–192, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31401-8_17/.

Hatzfeld:2012:MIC

[612] Christian Hatzfeld and Roland Werthschützky. Mechanical impedance as coupling parameter of force and deflection perception: Experimental evaluation. Lecture Notes in Computer Science, 7282:193–204, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31401-8_18/.

Hirota:2012:AED

[613] Koichi Hirota and Kazuyoshi Tagawa. Acquisition of elastically deformable object model based on measurement. Lecture Notes in Computer Science, 7282:205–217, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31401-8_19/.

Hudin:2012:TAT

[614] Charles Hudin, José Lozada, Michael Wiertlewski, and Vincent Hayward. Tradeoffs in the application of time-reversed acoustics to tactile stimulation. Lecture Notes in Computer Science, 7282:218–226, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31401-8_20/.

Anonymous:2012:FMbj

[615] Anonymous. Front matter. Lecture Notes in Computer Science, 7282: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31401-8/1.

Awed:2012:HFP

[616] Jalal Awed, Imad H. Elhajj, and Nadiya Slobodenyuk. Haptic force perception in bimanual manipulation. Lecture Notes in Computer Science, 7283:1–6, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31404-9_1/.

Tiest:2012:IEA

[617] Wouter M. Bergmann Tiest, Connie Lyklema, and Astrid M. L. Kappers. Investigating the effect of area of stimulation on cutaneous and proprioceptive weight perception. Lecture Notes in Computer Science, 7283:7–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31404-9_2/.

Coelho:2012:RHS

[618] Sandra Coelho and Miguel V. Correia. Rediscovering the haptic sense through crossroads of art and design research. Lecture Notes in Computer Science, 7283:13–18, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31404-9_3/.

Debats:2012:PHR

[619] Nienke B. Debats, Idsart Kingma, Peter J. Beek, and Jeroen B. J. Smeets. The precision of “haptic” rod length perception is reduced by lack of visual precision. Lecture Notes in Computer Science, 7283:19–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31404-9_4/.

VanDoorn:2012:CIY

[620] George H. Van Doorn, Barry L. Richardson, Mark A. Symmons, and Jacqui L. Howell. Cutaneous inputs yield judgments of line length that are equal to, or better than, those based on kinesthetic inputs. Lecture Notes in Computer Science, 7283:25–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31404-9_5/.

vanderGraaff:2012:CHM

[621] Marieke C. W. van der Graaff, Eli Brenner, and Jeroen B. J. Smeets. Is the curvature in hand movements to haptic targets in the mid sagittal plane caused by a misjudgment in direction? Lecture Notes in Computer Science, 7283:31–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31404-9_6/.

Gueniat:2012:HSD

[622] Florimond Guéniat, Yoren Gaffary, Luc Pastur, and Ammi Mehdi. Haptic stimulus for the discrimination between intrinsic properties of dynamic systems. Lecture Notes in Computer Science, 7283:37–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31404-9_7/.

Hatzfeld:2012:JND

[623] Christian Hatzfeld and Roland Werthschützky. Just noticeable differences of low-intensity vibrotactile forces at the fingertip. Lecture Notes in Computer Science, 7283:43–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31404-9_8/.

Gwak:2012:NSM

[624] Kiuk Gwak, Jun-Cheol Park, and Dae-Shik Kim. A novel stimulation method based on a neuromorphic mechanoreceptor model for haptic illusion. Lecture Notes in Computer Science, 7283:49–54, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31404-9_9/.

Howell:2012:MLV

[625] Jacqui L. Howell, Mark A. Symmons, and George H. Van Doorn. The misperception of length in vision, haptics and audition. Lecture Notes in Computer Science, 7283:55–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31404-9_10/.

Hwang:2012:EMG

[626] Inwook Hwang and Seungmoon Choi. Effect of mechanical ground on the vibrotactile perceived intensity of a handheld object. Lecture Notes in Computer Science, 7283:61–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31404-9_11/.

Kajimoto:2012:DCW

[627] Hiroyuki Kajimoto. Design of cylindrical whole-hand haptic interface using electrocutaneous display. Lecture Notes in Computer Science, 7283: 67–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31404-9_12/.

Keskinen:2012:UHF

[628] Tuuli Keskinen, Markku Turunen, Roope Raisamo, Grigori Evreinov, and Eemeli Haverinen. Utilizing haptic feedback in drill rigs. Lecture Notes in Computer Science, 7283:73–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31404-9_13/.

Kuroki:2012:DVF

[629] Scinob Kuroki, Junji Watanabe, and Shin’ya Nishida. Dissociation of vibrotactile frequency discrimination performances for supra-threshold and near-threshold vibrations. Lecture Notes in Computer Science, 7283: 79–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31404-9_14/.

Lee:2012:EIV

[630] Chang-Gyu Lee, Ian Oakley, and Jeha Ryu. Exploring the impact of visual-haptic registration accuracy in augmented reality. Lecture Notes in Computer Science, 7283:85–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31404-9_15/.

Lee:2012:SIS

[631] Jong uk Lee, Jeong-Mook Lim, Heesook Shin, and Ki-Uk Kyung. SHIFT: Interactive Smartphone bumper case. Lecture Notes in Computer Science, 7283:91–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31404-9_16/.

Nakano:2012:IWD

[632] Takuya Nakano, Shota Saji, and Yasuyuki Yanagida. Indicating wind direction using a fan-based wind display. Lecture Notes in Computer Science, 7283:97–102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31404-9_17/.

Okazaki:2012:VSC

[633] Ryuta Okazaki, Hiroyuki Kajimoto, and Vincent Hayward. Vibrotactile stimulation can affect auditory loudness: a pilot study. Lecture Notes in Computer Science, 7283:103–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31404-9_18/.

Olierook:2012:HRD

[634] Frederik Wilhelm Jozef Olierook and Mathias Funk. Hands reaching out of dreamland: a haptic peripheral alarm clock. Lecture Notes in Computer Science, 7283:109–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31404-9_19/.

Oshima:2012:HFM

[635] Kensuke Oshima and Shigeru Ichihara. How finger movement speed affects Braille pattern recognition. Lecture Notes in Computer Science, 7283: 115–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31404-9_20/.

Anonymous:2012:FMbk

[636] Anonymous. Front matter. Lecture Notes in Computer Science, 7283: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31404-9/1.

Kamal:2012:DDH

[637] Noreen Kamal and Sidney Fels. Determining the determinants of health behaviour change through an online social network. Lecture Notes in Computer Science, 7284:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31037-9_1/.

Burleson:2012:EIL

[638] Winslow Burleson, Naomi Newman, and Ryan Brotman. Empowering independent living for people with autism: Designing supportive, low-cost, interactive E-health environments. Lecture Notes in Computer Science, 7284:13–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31037-9_2/.

Chatterjee:2012:PSN

[639] Samir Chatterjee and Jongbok Byun. Persuasive sensing: a novel in-home monitoring technology to assist elderly adult diabetic patients. Lecture Notes in Computer Science, 7284:31–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31037-9_3/.

Chittaro:2012:TCS

[640] Luca Chittaro and Riccardo Sioni. Turning the classic snake mobile game into a location–based exergame that encourages walking. Lecture Notes in Computer Science, 7284:43–54, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31037-9_4/.

Zwinderman:2012:PRS

[641] Matthijs Jan Zwinderman and Azadeh Shirzad. Phone Row: a Smartphone game designed to persuade people to engage in moderate-intensity physical activity. Lecture Notes in Computer Science, 7284:55–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31037-9_5/.

Aagaard:2012:DPT

[642] Morten Aagaard and Peter Øhrstrøm. Developing persuasive technology for ASD challenged teenagers. Lecture Notes in Computer Science, 7284:67–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31037-9_6/.

Schumann:2012:TAH

[643] Sandy Schumann, Olivier Klein, and Karen Douglas. Talk to act: How Internet use empowers users to participate in collective actions offline. Lecture Notes in Computer Science, 7284:79–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31037-9_7/.

Midden:2012:IAI

[644] Cees Midden and Jaap Ham. The illusion of agency: The influence of the agency of an artificial agent on its persuasive power. Lecture Notes in Computer Science, 7284:90–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31037-9_8/.

Gamberini:2012:TFU

[645] Luciano Gamberini and Anna Spagnolli. Tailoring feedback to users’ actions in a persuasive game for household electricity conservation. Lecture Notes in Computer Science, 7284:100–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31037-9_9/.

Sundar:2012:MTT

[646] S. Shyam Sundar and Saraswathi Bellur. Motivational technologies: a theoretical framework for designing preventive health applications. Lecture Notes in Computer Science, 7284:112–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31037-9_10/.

Smids:2012:VPT

[647] Jilles Smids. The voluntariness of persuasive technology. Lecture Notes in Computer Science, 7284:123–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31037-9_11/.

Muller:2012:PRL

[648] Lars Müller and Verónica Rivera-Pelayo. Persuasion and reflective learning: Closing the feedback loop. Lecture Notes in Computer Science, 7284: 133–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31037-9_12/.

Waardenburg:2012:NSI

[649] Thijs Waardenburg and Robbert Winkel. Normative social influence in persuasive technology: Intensity versus effectiveness. Lecture Notes in Computer Science, 7284:145–156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31037-9_13/.

Drozd:2012:EPP

[650] Filip Drozd and Tuomas Lehto. Exploring perceived persuasiveness of a behavior change support system: a structural model. Lecture Notes in Computer Science, 7284:157–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31037-9_14/.

Bhatnagar:2012:BMP

[651] Nupur Bhatnagar and Abhishek Sinha. Biometric monitoring as a persuasive technology: Ensuring patients visit health centers in India’s slums. Lecture Notes in Computer Science, 7284:169–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31037-9_15/.

Ferebee:2012:NPM

[652] Susan Ferebee and James Davis. The neural persuasion model: Aligning neural readiness, perceived need, and intervention strategies. Lecture Notes in Computer Science, 7284:181–192, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31037-9_16/.

Wiafe:2012:APC

[653] Isaac Wiafe and Muna M. Alhammad. Analyzing the persuasion context of the persuasive systems design model with the 3D-RAB model. Lecture Notes in Computer Science, 7284:193–202, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31037-9_17/.

Orji:2012:TDD

[654] Rita Orji and Regan L. Mandryk. Towards a data-driven approach to intervention design: a predictive path model of healthy eating determinants. Lecture Notes in Computer Science, 7284:203–214, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31037-9_18/.

Chittaro:2012:PSA

[655] Luca Chittaro. Passengers’ safety in aircraft evacuations: Employing serious games to educate and persuade. Lecture Notes in Computer Science, 7284:215–226, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31037-9_19/.

Pribik:2012:TPT

[656] Ingo Pribik and Alexander Felfernig. Towards persuasive technology for software development environments: An empirical study. Lecture Notes in Computer Science, 7284:227–238, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31037-9_20/.

Anonymous:2012:FMbl

[657] Anonymous. Front matter. Lecture Notes in Computer Science, 7284: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31037-9/1.

He:2012:OBR

[658] Bryan He. Optimal binary representation of Mosaic floorplans and Baxter permutations. Lecture Notes in Computer Science, 7285:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29700-7_1/.

Wang:2012:SSC

[659] Jiun-Jie Wang and Xin He. Succinct strictly convex greedy drawing of 3-connected plane graphs. Lecture Notes in Computer Science, 7285: 13–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29700-7_2/.

Liu:2012:WIM

[660] Longcheng Liu, Yong Chen, Biao Wu, and Enyu Yao. Weighted inverse minimum cut problem under the sum-type Hamming distance. Lecture Notes in Computer Science, 7285:26–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29700-7_3/.

Fan:2012:VDV

[661] Chenglin Fan, Jun Luo, and Wencheng Wang. Voronoi diagram with visual restriction. Lecture Notes in Computer Science, 7285:36–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29700-7_4/.

Tan:2012:MMD

[662] Xuehou Tan and Bo Jiang. Minimization of the maximum distance between the two guards patrolling a polygonal region. Lecture Notes in Computer Science, 7285:47–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29700-7_5/.

Jiang:2012:CPM

[663] Minghui Jiang. On covering points with minimum turns. Lecture Notes in Computer Science, 7285:58–69, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29700-7_6/.

Hua:2012:EFP

[664] Xia Hua. On envy-free Pareto efficient pricing. Lecture Notes in Computer Science, 7285:70–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29700-7_7/.

Zhang:2012:OPM

[665] Yong Zhang and Francis Y. L. Chin. Online pricing for multi-type of items. Lecture Notes in Computer Science, 7285:82–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29700-7_8/.

Jiang:2012:ALN

[666] Yiwei Jiang, Zewei Weng, and Jueliang Hu. Algorithms with limited number of preemptions for scheduling on parallel machines. Lecture Notes in Computer Science, 7285:93–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29700-7_9/.

Chen:2012:CMN

[667] Danny Z. Chen, Xiaomin Liu, and Haitao Wang. Computing maximum non-crossing matching in convex bipartite graphs. Lecture Notes in Computer Science, 7285:105–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29700-7_10/.

Nishikawa:2012:ABC

[668] Kazuhide Nishikawa and Takao Nishizeki. Algorithms for bandwidth consecutive multicolorings of graphs. Lecture Notes in Computer Science, 7285:117–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29700-7_11/.

Song:2012:IDT

[669] Yu Song, Tian Liu, and Ke Xu. Independent domination on tree convex bipartite graphs. Lecture Notes in Computer Science, 7285:129– 138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29700-7_12/.

Ye:2012:LSP

[670] Deshi Ye and Lili Mei. On-line scheduling of parallel jobs in heterogeneous multiple clusters. Lecture Notes in Computer Science, 7285: 139–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29700-7_13/.

Bampis:2012:MTA

[671] Evripidis Bampis and Dimitrios Letsios. On multiprocessor temperatureaware scheduling problems. Lecture Notes in Computer Science, 7285: 149–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29700-7_14/.

Lan:2012:OMM

[672] Yan Lan, Xin Chen, Ning Ding, and György Dósa. Online minimum makespan scheduling with a buffer. Lecture Notes in Computer Science, 7285:161–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29700-7_15/.

Beigel:2012:DHS

[673] Richard Beigel and Bin Fu. A dense hierarchy of sublinear time approximation schemes for bin packing. Lecture Notes in Computer Science, 7285: 172–181, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29700-7_16/.

Fu:2012:MPI

[674] Bin Fu. Multivariate polynomial integration and differentiation are polynomial time inapproximable unless P=NP. Lecture Notes in Computer Science, 7285:182–191, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29700-7_17/.

Tang:2012:SRI

[675] Bangsheng Tang. Some remarks on the incompressibility of widthparameterized SAT instances. Lecture Notes in Computer Science, 7285: 192–198, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29700-7_18/.

Chen:2012:KPC

[676] Jianer Chen, Henning Fernau, and Peter Shaw. Kernels for packing and covering problems. Lecture Notes in Computer Science, 7285: 199–211, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29700-7_19/.

Zhou:2012:WCU

[677] Junping Zhou and Minghao Yin. The worst-case upper bound for exact 3-satisfiability with the number of clauses as the parameter. Lecture Notes in Computer Science, 7285:212–223, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29700-7_20/.

Anonymous:2012:FMbm

[678] Anonymous. Front matter. Lecture Notes in Computer Science, 7285: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29700-7/1.

Babik:2012:TCO

[679] Dmytro Babik, Lakshmi S. Iyer, and Eric W. Ford. Towards a comprehensive online peer assessment system. Lecture Notes in Computer Science, 7286:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29863-9_1/.

Hjalmarsson:2012:DDI

[680] Anders Hjalmarsson and Daniel Rudmark. Designing digital innovation contests. Lecture Notes in Computer Science, 7286:9–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29863-9_2/.

Iyer:2012:EAC

[681] Bala Iyer and George Wyner. Evaluating APIs: a call for design science research. Lecture Notes in Computer Science, 7286:28–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29863-9_3/.

Karunakaran:2012:DRP

[682] Arvind Karunakaran and Sandeep Purao. Designing for recombination: Process design through template combination. Lecture Notes in Computer Science, 7286:36–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29863-9_4/.

Lempinen:2012:DPI

[683] Heikki Lempinen, Matti Rossi, and Virpi Kristiina Tuunainen. Design principles for inter-organizational systems development — Case Hansel. Lecture Notes in Computer Science, 7286:52–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29863-9_5/.

Lessard:2012:UDS

[684] Lysanne Lessard and Eric Yu. Using design science research to develop a modeling technique for service design. Lecture Notes in Computer Science, 7286:66–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29863-9_6/.

VanLooy:2012:TDT

[685] Amy Van Looy, Manu De Backer, and Geert Poels. Towards a decision tool for choosing a business process maturity model. Lecture Notes in Computer Science, 7286:78–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29863-9_7/.

Lucas:2012:IDP

[686] Wendy Lucas and Tamara Babaian. Implementing design principles for collaborative ERP systems. Lecture Notes in Computer Science, 7286: 88–107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29863-9_8/.

Meyer:2012:ADS

[687] Martin Meyer, Markus Helfert, Brian Donnellan, and Jim Kenneally. Applying design science research for enterprise architecture business value assessments. Lecture Notes in Computer Science, 7286:108–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29863-9_9/.

Purao:2012:DLC

[688] Sandeep Purao, Narasimha Bolloju, and Chuan Hoo Tan. Designingin-the-Large: Combining local perspectives to generate enterprise-wide integration solutions. Lecture Notes in Computer Science, 7286:122– 138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29863-9_10/.

Valecha:2012:ERS

[689] Rohit Valecha, Raj Sharman, Raghav Rao, and Shambhu Upadhyaya. Emergency response system design: An examination of emergency communication messages. Lecture Notes in Computer Science, 7286:139– 146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29863-9_11/.

Varshney:2012:ASA

[690] Upkar Varshney. An approach for smart artifacts for mobile advertising. Lecture Notes in Computer Science, 7286:147–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29863-9_12/.

Voigt:2012:TUD

[691] Matthias Voigt, Björn Niehaves, and Jörg Becker. Towards a unified design theory for creativity support systems. Lecture Notes in Computer Science, 7286:152–173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29863-9_13/.

Yeo:2012:WMC

[692] M. Lisa Yeo and Ofer Arazy. What makes corporate Wikis work? Wiki affordances and their suitability for corporate knowledge work. Lecture Notes in Computer Science, 7286:174–190, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29863-9_14/.

Drechsler:2012:DSD

[693] Andreas Drechsler. Design science as design of social systems — implications for information systems research. Lecture Notes in Computer Science, 7286:191–205, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29863-9_15/.

Maass:2012:TDE

[694] Wolfgang Maass and Sabine Janzen. Towards design engineering of ubiquitous information systems. Lecture Notes in Computer Science, 7286: 206–219, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29863-9_16/.

Wieringa:2012:TAR

[695] Roel Wieringa and Ayşe Moralı. Technical action research as a validation method in information systems design science. Lecture Notes in Computer Science, 7286:220–238, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29863-9_17/.

Knol:2012:DES

[696] Arjan Knol, Henk Sol, and Johan van Wamelen. Decision enhancement for sourcing with shared service centres in the Dutch government. Lecture Notes in Computer Science, 7286:239–255, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29863-9_18/.

Pries-Heje:2012:DFV

[697] Jan Pries-Heje and Lene Pries-Heje. Designing a framework for virtual management and team building. Lecture Notes in Computer Science, 7286: 256–270, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29863-9_19/.

Silva:2012:IOD

[698] António Rito Silva and Michael Rosemann. Integrating organisational design with IT design. Lecture Notes in Computer Science, 7286:271– 286, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29863-9_20/.

Anonymous:2012:FMbn

[699] Anonymous. Front matter. Lecture Notes in Computer Science, 7286: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29863-9/1.

Hopcroft:2012:ITM

[700] John Hopcroft. On the impact of Turing machines. Lecture Notes in Computer Science, 7287:1–2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-29952-0_1.

Cooper:2012:TMM

[701] S. Barry Cooper. From Turing machine to morphogenesis: Forming and informing computation. Lecture Notes in Computer Science, 7287: 3–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29952-0_2/.

Karp:2012:TCE

[702] Richard M. Karp. Theory of computation as an enabling tool for the sciences. Lecture Notes in Computer Science, 7287:11, 2012.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-29952-0_3.

Li:2012:ICI

[703] Deyi Li and Liwei Huang. Interaction and collective intelligence on the Internet. Lecture Notes in Computer Science, 7287:12–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29952-0_4/.

Lampson:2012:WCD

[704] Butler Lampson. What computers do: Model, connect, engage. Lecture Notes in Computer Science, 7287:23–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29952-0_5/.

Li:2012:RCL

[705] Wei Li. R-calculus: a logical inference system for scientific discovery. Lecture Notes in Computer Science, 7287:27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/accesspage/chapter/10.1007/978-3-642-29952-0_6.

Yao:2012:QCG

[706] Andrew Chi-Chih Yao. Quantum computing: a great science in the making. Lecture Notes in Computer Science, 7287:28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/accesspage/chapter/10.1007/978-3-642-29952-0_7.

Kleinberg:2012:CST

[707] Jon Kleinberg. The convergence of social and technological networks. Lecture Notes in Computer Science, 7287:29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/accesspage/chapter/10.1007/978-3-642-29952-0_8.

Pan:2012:PNC

[708] Yicheng Pan. Principles of network computing. Lecture Notes in Computer Science, 7287:30–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29952-0_9/.

Peng:2012:SCP

[709] Pan Peng. The small community phenomenon in networks: Models, algorithms and applications. Lecture Notes in Computer Science, 7287:

40–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29952-0_10/.

Bonato:2012:VPH

[710] A. Bonato, D. Mitsche, and P. Prałat. Vertex-pursuit in hierarchical social networks. Lecture Notes in Computer Science, 7287:50–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29952-0_11/.

Zhang:2012:SAP

[711] Zipeng Zhang, Xinyu Feng, Ming Fu, Zhong Shao, and Yong Li. A structural approach to prophecy variables. Lecture Notes in Computer Science, 7287:61–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29952-0_12/.

Wang:2012:AGB

[712] Shuling Wang, Naijun Zhan, and Dimitar Guelev. An assume/guarantee based compositional calculus for hybrid CSP. Lecture Notes in Computer Science, 7287:72–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29952-0_13/.

Olderog:2012:AVR

[713] Ernst-Rüdiger Olderog. Automatic verification of real-time systems with rich data: An overview. Lecture Notes in Computer Science, 7287: 84–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29952-0_14/.

Kapur:2012:PAU

[714] Deepak Kapur. Program analysis using quantifier-elimination heuristics. Lecture Notes in Computer Science, 7287:94–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29952-0_15/.

Lawrence:2012:ETM

[715] Albert F. Lawrence, Séastien Phan, and Mark Ellisman. Electron tomography and multiscale biology. Lecture Notes in Computer Science, 7287: 109–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29952-0_16/.

Ito:2012:CTA

[716] Hiro Ito, Susumu Kiyoshima, and Yuichi Yoshida. Constant-time approximation algorithms for the knapsack problem. Lecture Notes in Computer Science, 7287:131–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29952-0_17/.

Bi:2012:LBS

[717] Jingguo Bi and Qi Cheng. Lower bounds of shortest vector lengths in random NTRU lattices. Lecture Notes in Computer Science, 7287: 143–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29952-0_18/.

Cerny:2012:PTC

[718] Michal Černý and Miroslav Rada. Polynomial time construction of ellipsoidal approximations of zonotopes given by generator descriptions. Lecture Notes in Computer Science, 7287:156–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29952-0_19/.

Popa:2012:HAA

[719] Alexandru Popa, Prudence W. H. Wong, and Fencol C. C. Yung. Hardness and approximation of the asynchronous border minimization problem. Lecture Notes in Computer Science, 7287:164–176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29952-0_20/.

Anonymous:2012:FMbo

[720] Anonymous. Front matter. Lecture Notes in Computer Science, 7287: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29952-0/1.

Flocchini:2012:DAF

[721] Paola Flocchini. Distributed algorithms by forgetful mobile robots. Lecture Notes in Computer Science, 7288:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/accesspage/chapter/10.1007/978-3-642-30347-0_1.

Persiano:2012:SML

[722] Giuseppe Persiano. Stability and metastability of the logit dynamics of strategic games. Lecture Notes in Computer Science, 7288:2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-30347-0_2.

Urrutia:2012:AGM

[723] Jorge Urrutia. Art galleries, k-modems, and k-convexity. Lecture Notes in Computer Science, 7288:3, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-30347-0_3.

Anderson:2012:VGK

[724] Terry Anderson and Therese Biedl. The Vulcan game of Kal-Toh: Finding or making triconnected planar subgraphs. Lecture Notes in Computer Science, 7288:4–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30347-0_4/.

Arkin:2012:STT

[725] Esther M. Arkin, Alon Efrat, George Hart, and Irina Kostitsyna. Scandinavian thins on top of cake: On the smallest one-size-fits-all box. Lecture Notes in Computer Science, 7288:16–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30347-0_5/.

Bender:2012:KPH

[726] Michael A. Bender, Ritwik Bose, Rezaul Chowdhury, and Samuel Mc-Cauley. The kissing problem: How to end a gathering when everyone kisses everyone else goodbye. Lecture Notes in Computer Science, 7288:28–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30347-0_6/.

Bruckdorfer:2012:MEC

[727] Till Bruckdorfer and Michael Kaufmann. Mad at edge crossings? Break the edges! Lecture Notes in Computer Science, 7288:40–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30347-0_7/.

Brueggeman:2012:TMF

[728] Leo Brueggeman, Michael Fellows, Rudolf Fleischer, and Martin Lackner. Train marshalling is fixed parameter tractable. Lecture Notes in Computer Science, 7288:51–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30347-0_8/.

Cannon:2012:CFG

[729] Sarah Cannon, Mashhood Ishaque, and Csaba D. Tóth. Conflict-free graph orientations with parity constraints. Lecture Notes in Computer Science, 7288:57–68, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30347-0_9/.

Cicalese:2012:MIU

[730] Ferdinando Cicalese. The multi-interval Ulam–Rényi game. Lecture Notes in Computer Science, 7288:69–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30347-0_10/.

Demaine:2012:PHP

[731] Erik D. Demaine, Martin L. Demaine, and Yair N. Minsky. Picturehanging puzzles. Lecture Notes in Computer Science, 7288:81–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30347-0_11/.

Dobrev:2012:OSN

[732] Stefan Dobrev, Lata Narayanan, and Jaroslav Opatrny. Optimal sensor networks for area monitoring using rotating and beam sensors. Lecture Notes in Computer Science, 7288:94–106, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30347-0_12/.

Dubois:2012:BBP

[733] Swan Dubois, Sébastien Tixeuil, and Nini Zhu. The Byzantine brides problem. Lecture Notes in Computer Science, 7288:107–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30347-0_13/.

Elmasry:2012:LPB

[734] Amr Elmasry and Jyrki Katajainen. Lean programs, branch mispredictions, and sorting. Lecture Notes in Computer Science, 7288:119– 130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30347-0_14/.

Engel:2012:CIR

[735] Jakob Engel, Markus Holzer, Oliver Ruepp, and Frank Sehnke. On computer integrated rationalized crossword puzzle manufacturing. Lecture Notes in Computer Science, 7288:131–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30347-0_15/.

Eppstein:2012:SSD

[736] David Eppstein. Solving single-digit Sudoku subproblems. Lecture Notes in Computer Science, 7288:142–153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30347-0_16/.

Flocchini:2012:FGC

[737] Paola Flocchini, Matthew Kellett, Peter C. Mason, and Nicola Santoro. Finding good coffee in Paris. Lecture Notes in Computer Science, 7288: 154–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30347-0_17/.

Fomin:2012:SIW

[738] Fedor V. Fomin, Frédéric Giroire, Alain Jean-Marie, and Dorian Mazauric. To satisfy impatient Web surfers is hard. Lecture Notes in Computer Science, 7288:166–176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30347-0_18/.

Fomin:2012:MLE

[739] Fedor V. Fomin, Pinar Heggernes, and Erik Jan van Leeuwen. Making life easier for firefighters. Lecture Notes in Computer Science, 7288: 177–188, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30347-0_19/.

Furer:2012:CPM

[740] Martin Fürer. Counting perfect matchings in graphs of degree 3. Lecture Notes in Computer Science, 7288:189–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30347-0_20/.

Anonymous:2012:FMbp

[741] Anonymous. Front matter. Lecture Notes in Computer Science, 7288: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30347-0/1.

Lee:2012:EUA

[742] HyunYong Lee and Akihiro Nakao. Efficient user-assisted content distribution over information-centric network. Lecture Notes in Computer Science, 7289:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30045-5_1/.

Katsaros:2012:IDN

[743] Konstantinos V. Katsaros, Nikos Fotiou, Xenofon Vasilakos, and Christopher N. Ververidis. On inter-domain name resolution for informationcentric networks. Lecture Notes in Computer Science, 7289:13–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30045-5_2/.

Chai:2012:CLM

[744] Wei Koong Chai, Diliang He, Ioannis Psaras, and George Pavlou. Cache “less for more” in information-centric networks. Lecture Notes in Computer Science, 7289:27–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30045-5_3/.

Guo:2012:CFC

[745] Shuo Guo, Haiyong Xie, and Guangyu Shi. Collaborative forwarding and caching in content centric networks. Lecture Notes in Computer Science, 7289:41–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30045-5_4/.

Blenn:2012:CDC

[746] Norbert Blenn, Christian Doerr, Bas Van Kester, and Piet Van Mieghem. Crawling and detecting community structure in online social networks using local information. Lecture Notes in Computer Science, 7289: 56–67, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30045-5_5/.

Jiang:2012:DCB

[747] Jin Jiang and Claudio E. Casetti. Distributed content backup and sharing using social information. Lecture Notes in Computer Science, 7289:

68–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30045-5_6/.

Apolonia:2012:TSN

[748] Nuno Apolónia, Paulo Ferreira, and Luís Veiga. Trans-social networks for distributed processing. Lecture Notes in Computer Science, 7289: 82–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30045-5_7/.

Blenn:2012:CSS

[749] Norbert Blenn, Kassandra Charalampidou, and Christian Doerr. Contextsensitive sentiment classification of short colloquial text. Lecture Notes in Computer Science, 7289:97–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30045-5_8/.

Vasconcelos:2012:RCN

[750] Marcelo F. Vasconcelos and Ronaldo M. Salles. Resilience in computer network management. Lecture Notes in Computer Science, 7289: 109–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30045-5_9/.

Genge:2012:ESI

[751] Béla Genge and Christos Siaterlis. An experimental study on the impact of network segmentation to the resilience of physical processes. Lecture Notes in Computer Science, 7289:121–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30045-5_10/.

Ben-Porat:2012:VHH

[752] Udi Ben-Porat, Anat Bremler-Barr, Hanoch Levy, and Bernhard Plattner. On the vulnerability of hardware hash tables to sophisticated attacks. Lecture Notes in Computer Science, 7289:135–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30045-5_11/.

Li:2012:DPE

[753] Cong Li, Huijuan Wang, and Piet Van Mieghem. Degree and principal eigenvectors in complex networks. Lecture Notes in Computer Science, 7289:149–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30045-5_12/.

Barla:2012:RVN

[754] Isil Burcu Barla, Dominic A. Schupke, and Georg Carle. Resilient virtual network design for end-to-end cloud services. Lecture Notes in Computer Science, 7289:161–174, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30045-5_13/.

Janssens:2012:DSC

[755] Nico Janssens, Xueli An, Koen Daenen, and Claudio Forlivesi. Dynamic scaling of call-stateful SIP services in the cloud. Lecture Notes in Computer Science, 7289:175–189, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30045-5_14/.

Mann:2012:RNA

[756] Vijay Mann, Akanksha Gupta, Partha Dutta, Anilkumar Vishnoi, and Parantapa Bhattacharya. Remedy: Network-aware steady state VM management for data centers. Lecture Notes in Computer Science, 7289: 190–204, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30045-5_15/.

Liu:2012:BFS

[757] Junjie Liu, Yingke Xie, Gaogang Xie, Layong Luo, Fuxing Zhang, and Xiaolong Wu. Building a flexible and scalable virtual hardware data plane. Lecture Notes in Computer Science, 7289:205–216, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30045-5_16/.

Vo:2012:PRI

[758] Hung Quoc Vo, Olav Lysne, and Amund Kvalbein. Permutation routing for increased robustness in IP networks. Lecture Notes in Computer Science, 7289:217–231, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30045-5_17/.

Shen:2012:RDT

[759] Meng Shen, Hongying Liu, Ke Xu, Ning Wang, and Yifeng Zhong. Routing on demand: Toward the energy-aware traffic engineering with OSPF. Lecture Notes in Computer Science, 7289:232–246, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30045-5_18/.

Giroire:2012:MNP

[760] Frédéric Giroire, Joanna Moulierac, Truong Khoa Phan, and Frédéric Roudaut. Minimization of network power consumption with redundancy elimination. Lecture Notes in Computer Science, 7289:247–258, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30045-5_19/.

Xiang:2012:SWY

[761] Yang Xiang, Zhiliang Wang, Jianping Wu, Xingang Shi, and Xia Yin. Sign what you really care about — secure BGP AS paths efficiently. Lecture Notes in Computer Science, 7289:259–273, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30045-5_20/.

Anonymous:2012:FMbq

[762] Anonymous. Front matter. Lecture Notes in Computer Science, 7289: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30045-5/1.

Ruckert:2012:QAP

[763] Julius Rückert, Osama Abboud, Thomas Zinner, Ralf Steinmetz, and David Hausheer. Quality adaptation in P2P video streaming based on objective QoE metrics. Lecture Notes in Computer Science, 7290: 1–14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30054-7_1/.

Hecht:2012:PPL

[764] Fabio V. Hecht, Thomas Bocek, Flávio Roberto Santos, and Burkhard Stiller. Playback policies for live and on-demand P2P video streaming. Lecture Notes in Computer Science, 7290:15–28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-30054-7_2/.

Roverso:2012:SHL

[765] Roberto Roverso, Sameh El-Ansary, and Seif Haridi. SmoothCache: HTTP-live streaming goes peer-to-peer. Lecture Notes in Computer Science, 7290:29–43, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30054-7_3/.

Hwang:2012:LVV

[766] K.-W. Hwang, D. Applegate, A. Archer, V. Gopalakrishnan, S. Lee, and V. Misra. Leveraging video viewing patterns for optimal content placement. Lecture Notes in Computer Science, 7290:44–58, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30054-7_4/.

Wang:2012:ETL

[767] Haiyang Wang, Feng Wang, Jiangchuan Liu, and Ke Xu. Enhancing traffic locality in BitTorrent via shared trackers. Lecture Notes in Computer Science, 7290:59–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30054-7_5/.

Zhao:2012:TBM

[768] Yong Zhao, Zhibin Zhang, Ting He, Alex X. Liu, Li Guo, and Binxing Fang. A task-based model for the lifespan of peer-to-peer swarms. Lecture Notes in Computer Science, 7290:71–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30054-7_6/.

Goncalves:2012:UCM

[769] Glauber D. Gonçalves, Anna Guimarães, Alex Borges Vieira, and Ítalo Cunha. Using centrality metrics to predict peer cooperation in live streaming applications. Lecture Notes in Computer Science, 7290:84– 96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30054-7_7/.

Kim:2012:CPD

[770] Seungbae Kim, Jinyoung Han, Taejoong Chung, Hyun chul Kim, and Ted “Taekyoung” Kwon. Content publishing and downloading practice in BitTorrent. Lecture Notes in Computer Science, 7290:97–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30054-7_8/.

Mikians:2012:TSC

[771] Jakub Mikians, Amogh Dhamdhere, Constantine Dovrolis, and Pere Barlet-Ros. Towards a statistical characterization of the interdomain traffic matrix. Lecture Notes in Computer Science, 7290:111–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30054-7_9/.

Liu:2012:CID

[772] Yujing Liu, Xiapu Luo, Rocky K. C. Chang, and Jinshu Su. Characterizing inter-domain rerouting after Japan earthquake. Lecture Notes in Computer Science, 7290:124–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30054-7_10/.

Dhamdhere:2012:MEI

[773] Amogh Dhamdhere, Himalatha Cherukuru, Constantine Dovrolis, and Kc Claffy. Measuring the evolution of Internet peering agreements. Lecture Notes in Computer Science, 7290:136–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30054-7_11/.

Hasan:2012:OGD

[774] Syed Hasan and Sergey Gorinsky. Obscure giants: Detecting the provider-free ASes. Lecture Notes in Computer Science, 7290:149–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30054-7_12/.

Hommes:2012:DSB

[775] Stefan Hommes, Radu State, and Thomas Engel. Detecting stealthy backdoors with association rule mining. Lecture Notes in Computer Science, 7290:161–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30054-7_13/.

Yang:2012:SAH

[776] Zichao Yang and John C. S. Lui. Security adoption in heterogeneous networks: the influence of cyber-insurance market. Lecture Notes in Computer Science, 7290:172–183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30054-7_14/.

Jerschow:2012:SCP

[777] Yves Igor Jerschow and Martin Mauve. Secure client puzzles based on random beacons. Lecture Notes in Computer Science, 7290:184– 197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30054-7_15/.

Djatmiko:2012:HSM

[778] Mentari Djatmiko, Mathieu Cunche, Roksana Boreli, and Aruna Seneviratne. Heterogeneous secure multi-party computation. Lecture Notes in Computer Science, 7290:198–210, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30054-7_16/.

Jimenez:2012:CAC

[779] Tania Jiménez, Yezekael Hayel, and Eitan Altman. Competition in access to content. Lecture Notes in Computer Science, 7290:211–222, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30054-7_17/.

Landa:2012:MTO

[780] Raul Landa, Eleni Mykoniati, Richard G. Clegg, David Griffin, and Miguel Rio. Modelling the tradeoffs in overlay-ISP cooperation. Lecture Notes in Computer Science, 7290:223–237, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30054-7_18/.

Gkorou:2012:RHD

[781] Dimitra Gkorou, Tamás Vinkó, Nitin Chiluka, Johan Pouwelse, and Dick Epema. Reducing the history in decentralized interaction-based reputation systems. Lecture Notes in Computer Science, 7290:238–251, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30054-7_19/.

Amigo:2012:PRS

[782] Isabel Amigo, Pablo Belzarena, and Sandrine Vaton. On the problem of revenue sharing in multi-domain federations. Lecture Notes in Computer Science, 7290:252–264, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30054-7_20/.

Anonymous:2012:FMbr

[783] Anonymous. Front matter. Lecture Notes in Computer Science, 7290: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30054-7/1.

Shi:2012:FMR

[784] Huaizhou Shi, R. Venkatesha Prasad, Vijay S. Rao, and I. G. M. M. Niemegeers. A fairness model for resource allocation in wireless networks. Lecture Notes in Computer Science, 7291:1–9, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30039-4_1/.

Ciancaglini:2012:ECM

[785] Vincenzo Ciancaglini, Luigi Liquori, Giang Ngo Hoang, and Petar Maksimović. An extension and cooperation mechanism for heterogeneous overlay networks. Lecture Notes in Computer Science, 7291:10–18, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30039-4_2/.

Xiao:2012:DPC

[786] Yong Xiao and Luiz A. DaSilva. Dynamic pricing coalitional game for cognitive radio networks. Lecture Notes in Computer Science, 7291: 19–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30039-4_3/.

Sabella:2012:ESL

[787] Dario Sabella, Marco Caretti, William Tomaselli, Valerio Palestini, and Bruno Cendón. Evaluation of ON-OFF schemes and linear prediction methods for increasing energy efficiency in mobile broadband networks. Lecture Notes in Computer Science, 7291:27–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30039-4_4/.

Martin-Hernandez:2012:CAT

[788] Javier Martín-Hernández, Christian Doerr, Johannes Lessmann, and Marcus Schöller. Challenge-aware traffic protection in wireless mobile backhaul networks. Lecture Notes in Computer Science, 7291:35–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30039-4_5/.

Abbassi:2012:DSF

[789] Mohtasim Abbassi, Shahbaz Khan, and M. Rahman. A distributed signaling fast mobile IPv6 scheme for next generation heterogeneous IP networks. Lecture Notes in Computer Science, 7291:43–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30039-4_6/.

Serrador:2012:EEG

[790] António Serrador and Luís M. Correia. Energy efficiency gains using VHOs in heterogeneous networks. Lecture Notes in Computer Science, 7291:52–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30039-4_7/.

Misseri:2012:IDR

[791] Xavier Misseri, Ivan Gojmerac, and Jean-Louis Rougier. Inter-domain route diversity for the Internet. Lecture Notes in Computer Science, 7291:63–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30039-4_8/.

Ghezzi:2012:PBM

[792] Antonio Ghezzi. A proposal of business model design parameters for future Internet carriers. Lecture Notes in Computer Science, 7291: 72–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30039-4_9/.

Lamali:2012:RAL

[793] Mohamed Lamine Lamali, Dominique Barth, and Johanne Cohen. Reputation-aware learning for SLA negotiation. Lecture Notes in Computer Science, 7291:80–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30039-4_10/.

Sackl:2012:QEW

[794] Andreas Sackl, Patrick Zwickl, and Peter Reichl. From quality of experience to willingness to pay for interconnection service quality. Lecture Notes in Computer Science, 7291:89–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30039-4_11/.

Zwickl:2012:IBA

[795] Patrick Zwickl and Peter Reichl. An instance-based approach for the quantitative assessment of key value network dependencies. Lecture Notes in Computer Science, 7291:97–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30039-4_12/.

Raju:2012:EIO

[796] Anand Raju, Vânia Gonçalves, Sven Lindmark, and Pieter Ballon. Evaluating impacts of oversubscription on future Internet business models. Lecture Notes in Computer Science, 7291:105–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30039-4_13/.

Agiatzidou:2012:IDC

[797] Eleni Agiatzidou, Costas Courcoubetis, Olivier Dugeon, and Finn-Tore Johansen. Inter-domain coordination models. Lecture Notes in Computer Science, 7291:113–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30039-4_14/.

Altman:2012:WCS

[798] Eitan Altman. In which content to specialize? A game theoretic analysis. Lecture Notes in Computer Science, 7291:121–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30039-4_15/.

Wollenberg:2012:EEN

[799] Till Wollenberg. Estimation of expectable network quality in wireless mesh networks. Lecture Notes in Computer Science, 7291:126–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30039-4_16/.

Krejcar:2012:DLM

[800] Ondrej Krejcar. Development of localization module for various smart devices platforms. Lecture Notes in Computer Science, 7291:133–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30039-4_17/.

Dvorsky:2012:IGB

[801] Marek Dvorsky, Libor Michalek, Pavel Moravec, and Roman Sebesta. Improved GSM-based localization by incorporating secondary network characteristics. Lecture Notes in Computer Science, 7291:139–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30039-4_18/.

Zelinka:2012:VCN

[802] Ivan Zelinka, Donald Davendra, and Lenka Skanderova. Visualization of complex networks dynamics: Case study. Lecture Notes in Computer Science, 7291:145–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30039-4_19/.

Kavlak:2012:PPS

[803] Hakan Kavlak and Hakki Ilk. PCI planning strategies for long term evolution networks. Lecture Notes in Computer Science, 7291:151–156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30039-4_20/.

Anonymous:2012:BMq

[804] Anonymous. Back matter. Lecture Notes in Computer Science, 7291: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-30039-4/1.

Anonymous:2012:FMbs

[805] Anonymous. Front matter. Lecture Notes in Computer Science, 7291: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30039-4/1.

Al-Turaiki:2012:TBA

[806] Isra Al-Turaiki, Ghada Badr, and Hassan Mathkour. Trie-based apriori motif discovery approach. Lecture Notes in Computer Science, 7292: 1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30191-9_1/.

Bulteau:2012:IED

[807] Laurent Bulteau and Minghui Jiang. Inapproximability of (1,2)-exemplar distance. Lecture Notes in Computer Science, 7292:13–23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30191-9_2/.

Catanzaro:2012:MIP

[808] Daniele Catanzaro and Martine Labbé. A mixed integer programming model for the parsimonious loss of heterozygosity problem. Lecture Notes in Computer Science, 7292:24–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30191-9_3/.

Chen:2012:RTR

[809] Xi Chen, Chen Wang, and Ayesha N. Shajahan. Reconstruction of transcription regulatory networks by stability-based network component analysis. Lecture Notes in Computer Science, 7292:36–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30191-9_4/.

Christinat:2012:TPE

[810] Yann Christinat and Bernard M. E. Moret. A transcript perspective on evolution. Lecture Notes in Computer Science, 7292:48–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30191-9_5/.

Crosby:2012:FAC

[811] Ralph W. Crosby and Tiffani L. Williams. A fast algorithm for computing the quartet distance for large sets of evolutionary trees. Lecture Notes in Computer Science, 7292:60–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30191-9_6/.

Damaschke:2012:EPS

[812] Peter Damaschke and Leonid Molokov. Error propagation in sparse linear systems with peptide-protein incidence matrices. Lecture Notes in Computer Science, 7292:72–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30191-9_7/.

DasGupta:2012:MAT

[813] Bhaskar DasGupta. Models and algorithmic tools for computational processes in cellular biology: Recent developments and future directions. Lecture Notes in Computer Science, 7292:84–86, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30191-9_8/.

Deepak:2012:IRT

[814] Akshay Deepak and Jianrong Dong. Identifying rogue taxa through reduced consensus: NP-hardness and exact algorithms. Lecture Notes in Computer Science, 7292:87–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30191-9_9/.

Gibas:2012:AAE

[815] Cynthia J. Gibas. Analytics approaches for the era of 10,000* genomes *and counting. Lecture Notes in Computer Science, 7292:99–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30191-9_10/.

Gorecki:2012:GSU

[816] Paweł Górecki and J. Gordon Burleigh. GTP supertrees from unrooted gene trees: Linear time algorithms for NNI based local searches. Lecture Notes in Computer Science, 7292:102–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30191-9_11/.

Gorecki:2012:RFM

[817] Paweł Górecki and Oliver Eulenstein. A Robinson–Foulds measure to compare unrooted trees with rooted trees. Lecture Notes in Computer Science, 7292:115–126, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30191-9_12/.

Guo:2012:PBS

[818] Fei Guo, Shuai Cheng Li, and Lusheng Wang. P-Binder: a system for the protein-protein binding sites identification. Lecture Notes in Computer Science, 7292:127–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30191-9_13/.

Kirkpatrick:2012:NIP

[819] Bonnie Kirkpatrick. Non-identifiable pedigrees and a Bayesian solution. Lecture Notes in Computer Science, 7292:139–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30191-9_14/.

Li:2012:IPL

[820] Juntao Li and Kwok Pui Choi. Iterative piecewise linear regression to accurately assess statistical significance in batch confounded differential expression analysis. Lecture Notes in Computer Science, 7292: 153–164, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30191-9_15/.

Li:2012:RNE

[821] Si Li, Kwok Pui Choi, Taoyang Wu, and Louxin Zhang. Reconstruction of network evolutionary history from extant network topology and duplication history. Lecture Notes in Computer Science, 7292:165–176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30191-9_16/.

Missirian:2012:PPP

[822] Victor Missirian, Isabelle Henry, and Luca Comai. POPE: Pipeline of parentally-biased expression. Lecture Notes in Computer Science, 7292: 177–188, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30191-9_17/.

Novak:2012:ONM

[823] Jiří Novák, David Hoksza, and Jakub Lokoč. On optimizing the nonmetric similarity search in tandem mass spectra by clustering. Lecture Notes in Computer Science, 7292:189–200, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30191-9_18/.

Ouangraoua:2012:CSA

[824] Aïda Ouangraoua and Krister M. Swenson. On the comparison of sets of alternative transcripts. Lecture Notes in Computer Science, 7292: 201–212, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30191-9_19/.

Park:2012:MFH

[825] Hyun Jung Park and Luay Nakhleh. MURPAR: a fast heuristic for inferring parsimonious phylogenetic networks from multiple gene trees. Lecture Notes in Computer Science, 7292:213–224, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30191-9_20/.

Anonymous:2012:FMbt

[826] Anonymous. Front matter. Lecture Notes in Computer Science, 7292: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30191-9/1.

Codish:2012:PBS

[827] Michael Codish. Programming with Boolean satisfaction. Lecture Notes in Computer Science, 7294:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-29822-6_1.

Terauchi:2012:AVH

[828] Tachio Terauchi. Automated verification of higher-order functional programs. Lecture Notes in Computer Science, 7294:2, 2012. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-29822-6_2.

Weirich:2012:DTP

[829] Stephanie Weirich. Dependently-typed programming in GHC. Lecture Notes in Computer Science, 7294:3, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-29822-6_3.

Accattoli:2012:CVS

[830] Beniamino Accattoli and Luca Paolini. Call-by-value solvability, revisited. Lecture Notes in Computer Science, 7294:4–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-29822-6_4/.

Antoy:2012:CFL

[831] Sergio Antoy and Arthur Peters. Compiling a functional logic language: The basic scheme. Lecture Notes in Computer Science, 7294: 17–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29822-6_5/.

Ariola:2012:CCN

[832] Zena M. Ariola, Paul Downen, and Hugo Herbelin. Classical call-by-need sequent calculi: The unity of semantic artifacts. Lecture Notes in Computer Science, 7294:32–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29822-6_6/.

Biernacki:2012:NFB

[833] Dariusz Biernacki and Sergueï Lenglet. Normal form bisimulations for delimited-control operators. Lecture Notes in Computer Science, 7294: 47–61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29822-6_7/.

Bouma:2012:RTP

[834] Gerlof Bouma. Real-time persistent queues and deques with logic variables (declarative pearl). Lecture Notes in Computer Science, 7294: 62–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29822-6_8/.

Caballero:2012:DDW

[835] Rafael Caballero and Yolanda García-Ruiz. Declarative debugging of wrong and missing answers for SQL views. Lecture Notes in Computer Science, 7294:73–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29822-6_9/.

Castineiras:2012:IPF

[836] Ignacio Castiñeiras and Fernando Sáenz-Pérez. Improving the performance of FD constraint solving in a CFLP system. Lecture Notes in Computer Science, 7294:88–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29822-6_10/.

deGuzman:2012:GIF

[837] Pablo Chico de Guzmán and Manuel Carro. A general implementation framework for tabled CLP. Lecture Notes in Computer Science, 7294: 104–119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29822-6_11/.

Estevez-Martin:2012:ESE

[838] Sonia Estévez-Martín and Jesús Correas Fernández. Extending the T OY system with the ECL^{i}PS^{e} solver over sets of integers. Lecture Notes in Computer Science, 7294:120–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29822-6_12/.

Hamana:2012:CLA

[839] Makoto Hamana. Correct looping arrows from cyclic terms. Lecture Notes in Computer Science, 7294:136–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29822-6_13/.

Hirai:2012:LCG

[840] Yoichi Hirai. A lambda calculus for Gödel–Dummett logic capturing Waitfreedom. Lecture Notes in Computer Science, 7294:151–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29822-6_14/.

Kiselyov:2012:I

[841] Oleg Kiselyov. Iteratees. Lecture Notes in Computer Science, 7294: 166–181, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29822-6_15/.

Kriener:2012:MEI

[842] Jael Kriener and Andy King. Mutual exclusion by interpolation. Lecture Notes in Computer Science, 7294:182–196, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29822-6_16/.

Lobachev:2012:PCS

[843] Oleg Lobachev. Parallel computation skeletons with premature termination property. Lecture Notes in Computer Science, 7294:197–212, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29822-6_17/.

Morihata:2012:CDN

[844] Akimasa Morihata. Calculational developments of new parallel algorithms for size-constrained maximum-sum segment problems. Lecture Notes in Computer Science, 7294:213–227, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29822-6_18/.

Rose:2012:DFL

[845] Kristoffer H. Rose, Lionel Villard, and Naoto Sato. A data flow language for hybrid query and programming languages. Lecture Notes in Computer Science, 7294:228–242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29822-6_19/.

Saeedloei:2012:CCL

[846] Neda Saeedloei and Gopal Gupta. Coinductive constraint logic programming. Lecture Notes in Computer Science, 7294:243–259, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29822-6_20/.

Anonymous:2012:FMbu

[847] Anonymous. Front matter. Lecture Notes in Computer Science, 7294: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29822-6/1.

Bernstein:2012:SWL

[848] Abraham Bernstein. Semantic Web/LD at a crossroads: Into the garbage can or to theory? Lecture Notes in Computer Science, 7295:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-30284-8_1.

vanGrondelle:2012:NAO

[849] Jeroen van Grondelle. New audiences for ontologies: Dealing with complexity in business processes. Lecture Notes in Computer Science, 7295:2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-30284-8_2.

Halevy:2012:BWD

[850] Alon Halevy. Bringing (web) databases to the masses. Lecture Notes in Computer Science, 7295:3, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-30284-8_3.

Kolcz:2012:LSL

[851] Aleksander Kołcz. Large scale learning at Twitter. Lecture Notes in Computer Science, 7295:4, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-30284-8_4.

Lam:2012:MDM

[852] Monica S. Lam. Musubi: a decentralized mobile social Web. Lecture Notes in Computer Science, 7295:5, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-30284-8_5.

Nagy-Rothengass:2012:DVC

[853] Márta Nagy-Rothengass. Data value chain in Europe. Lecture Notes in Computer Science, 7295:6, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-30284-8_6.

vandeLaar:2012:CTN

[854] Julius van de Laar. Cutting through the noise: How to shape public perception, frame the debate and effectively engage your audience in the digital age. Lecture Notes in Computer Science, 7295:7, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-30284-8_7.

Hartig:2012:SWL

[855] Olaf Hartig. SPARQL for a Web of linked data: Semantics and computability. Lecture Notes in Computer Science, 7295:8–23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30284-8_8/.

Damljanovic:2012:LDB

[856] Danica Damljanovic, Milan Stankovic, and Philippe Laublet. Linked data-based concept recommendation: Comparison of different methods in open innovation scenario. Lecture Notes in Computer Science, 7295: 24–38, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30284-8_9/.

Stankovic:2012:FCS

[857] Milan Stankovic, Matthew Rowe, and Philippe Laublet. Finding co-solvers on twitter, with a little help from linked data. Lecture Notes in Computer Science, 7295:39–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30284-8_10/.

Wagner:2012:TKL

[858] Andreas Wagner, Thanh Tran Duc, Günter Ladwig, Andreas Harth, and Rudi Studer. Top- k linked data query processing. Lecture Notes in Computer Science, 7295:56–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30284-8_11/.

Stolpe:2012:PIC

[859] Audun Stolpe and Martin G. Skjæveland. Preserving information content in RDF using bounded homomorphisms. Lecture Notes in Computer Science, 7295:72–86, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30284-8_12/.

Gueret:2012:ALD

[860] Christophe Guéret, Paul Groth, Claus Stadler, and Jens Lehmann. Assessing linked data mappings using network measures. Lecture Notes in Computer Science, 7295:87–102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30284-8_13/.

Sah:2012:NCB

[861] Melike Sah and Vincent Wade. A novel concept-based search for the Web of data using UMBEL and a fuzzy retrieval model. Lecture Notes in Computer Science, 7295:103–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30284-8_14/.

Nikolov:2012:ULL

[862] Andriy Nikolov, Mathieu d’Aquin, and Enrico Motta. Unsupervised learning of link discovery configuration. Lecture Notes in Computer Science, 7295:119–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30284-8_15/.

Losch:2012:GKR

[863] Uta Lösch, Stephan Bloehdorn, and Achim Rettinger. Graph kernels for RDF data. Lecture Notes in Computer Science, 7295:134–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30284-8_16/.

Ngomo:2012:EEA

[864] Axel-Cyrille Ngonga Ngomo and Klaus Lyko. EAGLE: Efficient active learning of link specifications using genetic programming. Lecture Notes in Computer Science, 7295:149–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30284-8_17/.

Jiang:2012:CIE

[865] Xueyan Jiang, Yi Huang, Maximilian Nickel, and Volker Tresp. Combining information extraction, deductive reasoning and machine learning for relation prediction. Lecture Notes in Computer Science, 7295: 164–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30284-8_18/.

Cruz:2012:ACS

[866] Isabel F. Cruz, Alessio Fabiani, Federico Caimi, Cosmin Stroe, and Matteo Palmonari. Automatic configuration selection using ontology matching task profiling. Lecture Notes in Computer Science, 7295: 179–194, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30284-8_19/.

Rubiera:2012:TRB

[867] Emilio Rubiera, Luis Polo, Diego Berrueta, and Adil El Ghali. TELIX: An RDF-based model for linguistic annotation. Lecture Notes in Computer Science, 7295:195–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30284-8_20/.

Anonymous:2012:FMbv

[868] Anonymous. Front matter. Lecture Notes in Computer Science, 7295: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30284-8/1.

Banditwattanawong:2012:WCC

[869] Thepparit Banditwattanawong. From Web cache to cloud cache. Lecture Notes in Computer Science, 7296:1–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30767-6_1/.

Wang:2012:PAR

[870] Jianzong Wang, Yanjun Chen, Daniel Gmach, Changsheng Xie, Jiguang Wan, and Rui Hua. pCloud: An adaptive I/O resource allocation algorithm with revenue consideration over public clouds. Lecture Notes in Computer Science, 7296:16–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30767-6_2/.

Lim:2012:GBM

[871] JongBeom Lim, Kwang-Sik Chung, Sung-Ho Chin, and Heon-Chang Yu. A Gossip-based mutual exclusion algorithm for cloud environments. Lecture Notes in Computer Science, 7296:31–45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30767-6_3/.

Qin:2012:EPA

[872] Zhuoran Qin, Jixian Zhang, and Xuejie Zhang. An effective partition approach for elastic application development on mobile cloud computing. Lecture Notes in Computer Science, 7296:46–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-30767-6_4/.

Ruan:2012:MVM

[873] Li Ruan, Huixiang Wang, Limin Xiao, Mingfa Zhu, and Feibo Li. Memory virtualization for MIPS processor based cloud server. Lecture Notes in Computer Science, 7296:54–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30767-6_5/.

Yang:2012:IDD

[874] Chao-Tung Yang, Wen-Chung Shih, and Chih-Lin Huang. Implementation of a distributed data storage system with resource monitoring on cloud computing. Lecture Notes in Computer Science, 7296:64–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30767-6_6/.

Abidi:2012:DVP

[875] Leila Abidi, Christophe Cérin, and Kais Klai. Design, verification and prototyping the next generation of desktop grid middleware. Lecture Notes in Computer Science, 7296:74–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30767-6_7/.

Gu:2012:RMM

[876] Pingli Gu, Yanlei Shang, Junliang Chen, Bo Cheng, and Yan Jiang. A request multiplexing method based on multiple tenants in SaaS. Lecture Notes in Computer Science, 7296:89–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30767-6_8/.

Mannava:2012:ADP

[877] Vishnuvardhan Mannava and T. Ramesh. An adaptive design pattern for genetic algorithm-based composition of Web services in autonomic computing systems using SOA. Lecture Notes in Computer Science, 7296: 98–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30767-6_9/.

Pan:2012:SOO

[878] Weisen Pan, Shizhan Chen, and Zhiyong Feng. Service-oriented ontology and its evolution. Lecture Notes in Computer Science, 7296: 109–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30767-6_10/.

Liang:2012:EEA

[879] Yunji Liang, Xingshe Zhou, Zhiwen Yu, Bin Guo, and Yue Yang. Energy efficient activity recognition based on low resolution accelerometer in smart phones. Lecture Notes in Computer Science, 7296:122– 136, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30767-6_11/.

Cao:2012:EEA

[880] Jian Cao, Yihua Wu, and Minglu Li. Energy efficient allocation of virtual machines in cloud computing environments based on demand forecast. Lecture Notes in Computer Science, 7296:137–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30767-6_12/.

Chandrasekar:2012:ECM

[881] Ashok Chandrasekar, Karthik Chandrasekar, Harini Ramasatagopan, and Rafica Abdul Rahim. Energy conservative mobile cloud infrastructure. Lecture Notes in Computer Science, 7296:152–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30767-6_13/.

Lee:2012:PCA

[882] Junghoon Lee, Gyung-Leen Park, Ho-Young Kwak, and Jikwang Han. Power-constrained actuator coordination for agricultural sensor networks. Lecture Notes in Computer Science, 7296:162–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-30767-6_14/.

Kim:2012:DEM

[883] Jennifer Kim. Design and evaluation of mobile applications with full and partial offloadings. Lecture Notes in Computer Science, 7296:172– 182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30767-6_15/.

Zhang:2012:CLS

[884] Fu-Quan Zhang and Inwhee Joe. A cross-layer scheme to improve TCP performance in wireless multi-hop networks. Lecture Notes in Computer Science, 7296:183–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30767-6_16/.

Xue:2012:FAV

[885] Jianxin Xue and Xiaoju Dong. A fully abstract view for local cause semantics. Lecture Notes in Computer Science, 7296:198–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30767-6_17/.

Kakkad:2012:ECP

[886] Jignesh Kakkad and Nandan Parameswaran. Efficiency considerations in policy based management in resource constrained devices. Lecture Notes in Computer Science, 7296:210–220, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30767-6_18/.

Zheng:2012:ABQ

[887] Di Zheng, Jun Wang, and Ke rong Ben. Agent based quality management middleware for context-aware pervasive applications. Lecture Notes in Computer Science, 7296:221–230, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30767-6_19/.

Cui:2012:VFS

[888] Yabing Cui, Chunming Hu, Tianyu Wo, and Hanwen Wang. A virtual file system for streaming loading of virtual software on Windows NT. Lecture Notes in Computer Science, 7296:231–243, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30767-6_20/.

Anonymous:2012:FMbw

[889] Anonymous. Front matter. Lecture Notes in Computer Science, 7296: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30767-6/1.

Yang:2012:AAC

[890] Wan-Shiou Yang and Shi-Xin Weng. Application of the ant colony optimization algorithm to competitive viral marketing. Lecture Notes in Computer Science, 7297:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30448-4_1/.

Pothitos:2012:CPC

[891] Nikolaos Pothitos, George Kastrinis, and Panagiotis Stamatopoulos. Constraint propagation as the core of local search. Lecture Notes in Computer Science, 7297:9–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30448-4_2/.

Kollia:2012:SQA

[892] Ilianna Kollia, Kostas Rapantzikos, Giorgos Stamou, and Andreas Stafylopatis. Semantic query answering in digital libraries. Lecture Notes in Computer Science, 7297:17–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30448-4_3/.

Revithis:2012:SBV

[893] Spyridon Revithis and Georgios Tagalakis. A SOM-based validation approach to a neural circuit theory of autism. Lecture Notes in Computer Science, 7297:25–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30448-4_4/.

Panagiotopoulou:2012:WSD

[894] Vicky Panagiotopoulou, Iraklis Varlamis, Ion Androutsopoulos, and George Tsatsaronis. Word sense disambiguation as an integer linear programming problem. Lecture Notes in Computer Science, 7297:33– 40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30448-4_5/.

Kokkinos:2012:PLC

[895] Yiannis Kokkinos and Konstantinos Margaritis. Parallelism, localization and chain gradient tuning combinations for fast scalable probabilistic neural networks in data mining applications. Lecture Notes in Computer Science, 7297:41–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30448-4_6/.

Rigas:2012:SOO

[896] Emmanouil Rigas, Georgios Meditskos, and Nick Bassiliades. SWRL2COOL: Object-oriented transformation of SWRL in the CLIPS production rule engine. Lecture Notes in Computer Science, 7297:49–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30448-4_7/.

Zouboulidis:2012:FFF

[897] Elias Zouboulidis and Sotiris Kotsiantis. Forecasting fraudulent financial statements with committee of cost-sensitive decision tree classifiers. Lecture Notes in Computer Science, 7297:57–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30448-4_8/.

Deligianni:2012:FCB

[898] Despina Deligianni and Sotiris Kotsiantis. Forecasting corporate bankruptcy with an ensemble of classifiers. Lecture Notes in Computer Science, 7297:65–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30448-4_9/.

Tzortzis:2012:GUM

[899] Grigorios Tzortzis and Aristidis Likas. Greedy unsupervised multiple kernel learning. Lecture Notes in Computer Science, 7297:73–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30448-4_10/.

Mimis:2012:AAM

[900] Angelos Mimis, Antonis Rovolis, and Marianthi Stamou. An AZP-ACO method for region-building. Lecture Notes in Computer Science, 7297: 81–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30448-4_11/.

Gavriilidis:2012:EQM

[901] Vasileios Gavriilidis and Anastasios Tefas. Exploiting quadratic mutual information for discriminant analysis. Lecture Notes in Computer Science, 7297:90–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30448-4_12/.

Kalles:2012:ESS

[902] Dimitris Kalles, Vassiliki Mperoukli, and Andreas Papandreadis. Emerge-Sort: Swarm intelligence sorting. Lecture Notes in Computer Science, 7297:98–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30448-4_13/.

Anastassiou:2012:ILC

[903] Vassileios-Marios Anastassiou, Panagiotis Diamantopoulos, and Stavros Vassos. iThink: a library for classical planning in video-games. Lecture Notes in Computer Science, 7297:106–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30448-4_14/.

Giannakopoulos:2012:DHF

[904] George Giannakopoulos, Vangelis Karkaletsis, and George A. Vouros. Detecting human features in summaries — symbol sequence statistical regularity. Lecture Notes in Computer Science, 7297:114–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30448-4_15/.

Lagani:2012:LME

[905] Vincenzo Lagani, Ioannis Tsamardinos, and Sofia Triantafillou. Learning from mixture of experimental data: a constraint–based approach. Lecture Notes in Computer Science, 7297:124–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30448-4_16/.

Vouros:2012:ISS

[906] George A. Vouros. Information sharing and searching via collaborative reinforcement learning. Lecture Notes in Computer Science, 7297: 132–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30448-4_17/.

Santipantakis:2012:DIR

[907] George M. Santipantakis and George A. Vouros. Distributed instance retrieval in E_{HQ^{+}}^{DDL}SHIQ representation framework. Lecture Notes in Computer Science, 7297:141–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30448-4_18/.

Galanopoulos:2012:EON

[908] Damianos Galanopoulos, Christos Athanasiadis, and Anastasios Tefas. Evolutionary optimization of a neural network controller for car racing simulation. Lecture Notes in Computer Science, 7297:149–156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30448-4_19/.

Papageorgiou:2012:BNH

[909] Elpiniki I. Papageorgiou, Panagiotis Oikonomou, and Arthi Kannappan. Bagged nonlinear Hebbian learning algorithm for fuzzy cognitive maps working on classification tasks. Lecture Notes in Computer Science, 7297: 157–164, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30448-4_20/.

Anonymous:2012:FMbx

[910] Anonymous. Front matter. Lecture Notes in Computer Science, 7297: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30448-4/1.

Araya:2012:CBC

[911] Ignacio Araya and Gilles Trombettoni. A contractor based on convex interval Taylor. Lecture Notes in Computer Science, 7298:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29828-8_1/.

Bardin:2012:FCA

[912] Sébastien Bardin and Arnaud Gotlieb. fdcc: a combined approach for solving constraints over finite domains and arrays. Lecture Notes in Computer Science, 7298:17–33, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29828-8_2/.

Bergman:2012:VOA

[913] David Bergman and Andre A. Cire. Variable ordering for the application of BDDs to the maximum independent set problem. Lecture Notes in Computer Science, 7298:34–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29828-8_3/.

Bergman:2012:GCF

[914] David Bergman and John N. Hooker. Graph coloring facets from all-different systems. Lecture Notes in Computer Science, 7298:50–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29828-8_4/.

Billaut:2012:CCN

[915] Jean-Charles Billaut and Emmanuel Hebrard. Complete characterization of near-optimal sequences for the two-machine flow shop scheduling problem. Lecture Notes in Computer Science, 7298:66–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29828-8_5/.

Bonfietti:2012:GCC

[916] Alessio Bonfietti and Michele Lombardi. Global cyclic cumulative constraint. Lecture Notes in Computer Science, 7298:81–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29828-8_6/.

Cambazard:2012:CGB

[917] Hadrien Cambazard and Deepak Mehta. A computational geometry-based local search algorithm for planar location problems. Lecture Notes in Computer Science, 7298:97–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29828-8_7/.

Chabert:2012:CIM

[918] Gilles Chabert and Sophie Demassey. The conjunction of interval A mong constraints. Lecture Notes in Computer Science, 7298:113–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29828-8_8/.

Cire:2012:FBC

[919] Andre A. Cire and Elvin Coban. Flow-based combinatorial chance constraints. Lecture Notes in Computer Science, 7298:129–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29828-8_9/.

Downing:2012:EFB

[920] Nicholas Downing and Thibaut Feydy. Explaining flow-based propagation. Lecture Notes in Computer Science, 7298:146–162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29828-8_10/.

Farnqvist:2012:COP

[921] Tommy Färnqvist. Constraint optimization problems and bounded treewidth revisited. Lecture Notes in Computer Science, 7298:163–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29828-8_11/.

Fontaine:2012:HLL

[922] Daniel Fontaine and Laurent Michel. A high level language for solver independent model manipulation and generation of hybrid solvers. Lecture Notes in Computer Science, 7298:180–194, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29828-8_12/.

Gange:2012:EPD

[923] Graeme Gange and Peter J. Stuckey. Explaining propagators for s-DNNF circuits. Lecture Notes in Computer Science, 7298:195–210, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29828-8_13/.

Heinz:2012:RMI

[924] Stefan Heinz and J. Christopher Beck. Reconsidering mixed integer programming and MIP-based hybrids for scheduling. Lecture Notes in Computer Science, 7298:211–227, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29828-8_14/.

Michel:2012:ABS

[925] Laurent Michel and Pascal Van Hentenryck. Activity-based search for black-box constraint programming solvers. Lecture Notes in Computer Science, 7298:228–243, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29828-8_15/.

Malitsky:2012:ISA

[926] Yuri Malitsky and Meinolf Sellmann. Instance-specific algorithm configuration as a method for non-model-based portfolio generation. Lecture Notes in Computer Science, 7298:244–259, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29828-8_16/.

Massen:2012:PBH

[927] Florence Massen and Yves Deville. Pheromone-based heuristic column generation for vehicle routing problems with black box feasibility. Lecture Notes in Computer Science, 7298:260–274, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29828-8_17/.

Masson:2012:STP

[928] Renaud Masson, Fabien Lehuédé, and Olivier Péton. Simple temporal problems in route scheduling for the Dial-a-Ride Problem with transfers. Lecture Notes in Computer Science, 7298:275–291, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29828-8_18/.

Pham:2012:SLS

[929] Quang Dung Pham and Yves Deville. Solving the longest simple path problem with constraint-based techniques. Lecture Notes in Computer Science, 7298:292–306, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29828-8_19/.

Ponte:2012:BSM

[930] Aníbal Ponte, Luís Paquete, and José R. Figueira. On beam search for multicriteria combinatorial optimization problems. Lecture Notes in Computer Science, 7298:307–321, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29828-8_20/.

Anonymous:2012:FMby

[931] Anonymous. Front matter. Lecture Notes in Computer Science, 7298: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29828-8/1.

Thuraisingham:2012:CCA

[932] Bhavani Thuraisingham, Vaibhav Khadilkar, Jyothsna Rachapalli, and Tyrone Cadenhead. Cloud-centric assured information sharing. Lecture Notes in Computer Science, 7299:1–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30428-6_1/.

Yunos:2012:IAT

[933] Zahri Yunos, Rabiah Ahmad, Syahrir Mat Ali, and Solahuddin Shamsuddin. Illicit activities and terrorism in cyberspace: An exploratory study in the Southeast Asian region. Lecture Notes in Computer Science, 7299:27–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30428-6_2/.

Chang:2012:CTP

[934] Weiping Chang, Yungchang Ku, Sinru Wu, and Chaochang Chiu. CybercrimeIR — a technological perspective to fight cybercrime. Lecture Notes in Computer Science, 7299:36–44, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30428-6_3/.

Xia:2012:ICT

[935] Xin Xia, Xiaohu Yang, Chao Wu, Shanping Li, and Linfeng Bao. Information credibility on Twitter in emergency situation. Lecture Notes in Computer Science, 7299:45–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30428-6_4/.

Li:2012:MLN

[936] Chao Li, Jun Luo, Joshua Zhexue Huang, and Jianping Fan. Multilayer network for influence propagation over microblog. Lecture Notes in Computer Science, 7299:60–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30428-6_5/.

Li:2012:HSO

[937] Tim M. H. Li, Michael Chau, Paul W. C. Wong, and Paul S. F. Yip. A hybrid system for online detection of emotional distress. Lecture Notes in Computer Science, 7299:73–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30428-6_6/.

Huynh:2012:PMS

[938] Viet H. Huynh and An N. T. Le. Process mining and security: Visualization in database intrusion detection. Lecture Notes in Computer Science, 7299:81–95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30428-6_7/.

Faisal:2012:SAM

[939] Mustafa Amir Faisal, Zeyar Aung, John R. Williams, and Abel Sanchez. Securing advanced metering infrastructure using intrusion detection system with data stream mining. Lecture Notes in Computer Science, 7299: 96–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30428-6_8/.

Bogdanov:2012:UTC

[940] Dan Bogdanov, Roman Jagomägis, and Sven Laur. A universal toolkit for cryptographically secure privacy-preserving data mining. Lecture Notes in Computer Science, 7299:112–126, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30428-6_9/.

Valcik:2012:IWC

[941] Jakub Valcik, Jan Sedmidubsky, Michal Balazia, and Pavel Zezula. Identifying walk cycles for human recognition. Lecture Notes in Computer Science, 7299:127–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30428-6_10/.

Ling:2012:CON

[942] Yajuan Ling, Jing Yang, and Liang He. Chinese organization name recognition based on multiple features. Lecture Notes in Computer Science, 7299:136–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30428-6_11/.

Jadalla:2012:PDS

[943] Ameera Jadalla and Ashraf Elnagar. A plagiarism detection system for Arabic text-based documents. Lecture Notes in Computer Science, 7299: 145–153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30428-6_12/.

Sharef:2012:ETC

[944] Nurfadhlina Mohd Sharef and Khairul Azhar Kasmiran. Examining text categorization methods for incidents analysis. Lecture Notes in Computer Science, 7299:154–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30428-6_13/.

Sun:2012:MCT

[945] Weiwei Sun, Bo Yang, Qiong Huang, Sha Ma, and Ximing Li. Multicommitter threshold commitment scheme from lattice. Lecture Notes in Computer Science, 7299:162–175, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30428-6_14/.

Chen:2012:CKS

[946] Zhenhua Chen, Chunying Wu, Daoshun Wang, and Shundong Li. Conjunctive keywords searchable encryption with efficient pairing, constant ciphertext and short trapdoor. Lecture Notes in Computer Science, 7299: 176–189, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30428-6_15/.

Yu:2012:SME

[947] Jia Yu, Shuguang Wang, Huawei Zhao, Minglei Shu, Jialiang Lv, and Qiang Guo. A simultaneous members enrollment and revocation protocol for secret sharing schemes. Lecture Notes in Computer Science, 7299: 190–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-30428-6_16/.

Jiang:2012:SDB

[948] Biao Jiang, Eul Gyu Im, and Yunmo Koo. SaaS-driven botnets. Lecture Notes in Computer Science, 7299:198–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30428-6_17/.

Anonymous:2012:BMr

[949] Anonymous. Back matter. Lecture Notes in Computer Science, 7299: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-30428-6/1.

Anonymous:2012:FMbz

[950] Anonymous. Front matter. Lecture Notes in Computer Science, 7299: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30428-6/1.

Berthold:2012:BKD

[951] Michael R. Berthold, editor. Bisociative Knowledge Discovery: An Introduction to Concept, Algorithms, Tools, and Applications, volume 7250 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31829-0 (print), 3-642-31830-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-31830-6.

Donnelly:2012:IAS

[952] Mark Donnelly, Cristiano Paggetti, Chris Nugent, and Mounir Mokhtari, editors. Impact Analysis of Solutions for Chronic Disease Prevention and Management: 10th International Conference on Smart Homes and Health Telematics, ICOST 2012, Artiminio, Italy, June 12–15, 2012. Proceedings, volume 7251 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30778-7 (print), 3-642-30779-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-30779-9.

Salerno:2012:CIM

[953] Emanuele Salerno, A. Enis Çetin, and Ovidio Salvetti, editors. Computational Intelligence for Multimedia Understanding: International Workshop, MUSCLE 2011, Pisa, Italy, December 13–15, 2011, Revised Selected Papers, volume 7252 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32435-5 (print), 3-642-32436-3 (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-32436-9.

Arbab:2012:FAC

[954] Farhad Arbab and Peter Csaba Ölveczky, editors. Formal Aspects of Component Software: 8th International Symposium, FACS 2011, Oslo, Norway, September 14–16, 2011, Revised Selected Papers, volume 7253 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35742-3 (print), 3-642-35743-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-35743-5.

Cranefield:2012:COI

[955] Stephen Cranefield, M. Birna van Riemsdijk, Javier Vázquez-Salceda, and Pablo Noriega, editors. Coordination, Organizations, Institutions, and Norms in Agent System VII: COIN 2011 International Workshops, COIN@AAMAS 2011, Taipei, Taiwan, May 3, 2011, COIN@WI-IAT 2011, Lyon, France, August 22, 2011, Revised Selected Papers, volume 7254 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35544-7 (print), 3-642-35545-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
35545-5.

Peters:2012:TRS

[956] James F. Peters and Andrzej Skowron, editors. Transactions on Rough Sets XV, volume 7255 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31902-5 (print), 3-642-31903-3 (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-31903-7.

Fernandez-Baca:2012:LTI

[957] David Fernández-Baca, editor. LATIN 2012: Theoretical Informatics: 10th Latin American Symposium, Arequipa, Peru, April 16–20, 2012. Proceedings, volume 7256 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29343-3 (print), 3-642-29344-1 (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-29344-3.

Gill:2012:IAF

[958] Andy Gill and Jurriaan Hage, editors. Implementation and Application of Functional Languages: 23rd International Symposium, IFL 2011, Lawrence, KS, USA, October 3–5, 2011, Revised Selected Papers, volume 7257 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34406-2 (print), 3-642-34407-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
34407-7.

Okumura:2012:NFA

[959] Manabu Okumura, Daisuke Bekki, and Ken Satoh, editors. New Frontiers in Artificial Intelligence: JSAI-isAI 2011 Workshops, LENLS, JURISIN, ALSIP, MiMI, Takamatsu, Japan, December 1–2, 2011. Revised Selected Papers, volume 7258 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32089-9 (print), 3-642-32090-2 (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-32090-3.

Kim:2012:ISC

[960] Howon Kim, editor. Information Security and Cryptology — ICISC 2011: 14th International Conference, Seoul, Korea, November 30–December 2, 2011. Revised Selected Papers, volume 7259 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31911-4 (print), 3-642-31912-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-31912-9.

Dusterhoft:2012:CMT

[961] Antje Düsterhöft, Meike Klettke, and Klaus-Dieter Schewe, editors. Conceptual Modelling and Its Theoretical Foundations: Essays Dedicated to Bernhard Thalheim on the Occasion of His 60th Birthday, volume 7260 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28278-4 (print), 3-642-28279-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-28279-9.

Eder:2012:HSV

[962] Kerstin Eder, João Lourenço, and Onn Shehory, editors. Hardware and Software: Verification and Testing: 7th International Haifa Verification Conference, HVC 2011, Haifa, Israel, December 6–8, 2011, Revised Selected Papers, volume 7261 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34187-X (print), 3-642-34188-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-34188-5.

Chor:2012:RCM

[963] Benny Chor, editor. Research in Computational Molecular Biology: 16th Annual International Conference, RECOMB 2012, Barcelona, Spain, April 21–24, 2012. Proceedings, volume 7262 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29626-2 (print), 3-642-29627-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-29627-7.

Linte:2012:AEC

[964] Cristian A. Linte, John T. Moore, Elvis C. S. Chen, and David R. Holmes III, editors. Augmented Environments for Computer-Assisted Interventions: 6th International Workshop, AE-CAI 2011, Held in Conjunction with MICCAI 2011, Toronto, ON, Canada, September 22, 2011, Revised Selected Papers, volume 7264 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32629-3 (print), 3-642-32630-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-32630-1.

Erdem:2012:CRE

[965] Esra Erdem, Joohyung Lee, Yuliya Lierler, and David Pearce, editors. Correct Reasoning: Essays on Logic-Based AI in Honour of Vladimir Lifschitz, volume 7265 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30742-6 (print), 3-642-30743-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-30743-0.

Vinel:2012:CTV

[966] Alexey Vinel, Rashid Mehmood, Marion Berbineau, Cristina Rico Garcia, Chung-Ming Huang, and Naveen Chilamkurti, editors. Communication Technologies for Vehicles: 4th International Workshop, Nets4Cars/ Nets4Trains 2012, Vilnius, Lithuania, April 25–27, 2012. Proceedings, volume 7266 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29666-1 (print), 3-642-29667-X (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-29667-3.

Rutkowski:2012:AISa

[967] Leszek Rutkowski, Marcin Korytkowski, Rafał Scherer, Ryszard Tadeusiewicz, Lotfi A. Zadeh, and Jacek M. Zurada, editors. Artificial Intelligence and Soft Computing: 11th International Conference, ICAISC 2012, Zakopane, Poland, April 29–May 3, 2012, Proceedings, Part I, volume 7267 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29346-8 (print), 3-642-29347-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-29347-4.

Rutkowski:2012:AISb

[968] Leszek Rutkowski, Marcin Korytkowski, Rafał Scherer, Ryszard Tadeusiewicz, Lotfi A. Zadeh, and Jacek M. Zurada, editors. Artificial Intelligence and Soft Computing: 11th International Conference, ICAISC 2012, Zakopane, Poland, April 29–May 3, 2012, Proceedings, Part II, volume 7268 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29349-2 (print), 3-642-29350-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-29350-4.

Rutkowski:2012:SEC

[969] Leszek Rutkowski, Marcin Korytkowski, Rafał Scherer, Ryszard Tadeusiewicz, Lotfi A. Zadeh, and Jacek M. Zurada, editors. Swarm and Evolutionary Computation: International Symposia, SIDE 2012 and EC 2012, Held in Conjunction with ICAISC 2012, Zakopane, Poland, April 29–May 3, 2012. Proceedings, volume 7269 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29352-2 (print), 3-642-29353-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-29353-5.

Nguyen:2012:TCC

[970] Ngoc Thanh Nguyen, editor. Transactions on Computational Collective Intelligence VII, volume 7270 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32065-1 (print), 3-642-32066-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-32066-8.

Leavens:2012:TAO

[971] Gary T. Leavens, Shigeru Chiba, Michael Haupt, Klaus Ostermann, and Eric Wohlstadter, editors. Transactions on Aspect-Oriented Software Development IX, volume 7271 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35550-1 (print), 3-642-35551-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-35551-6.

Goschka:2012:DAI

[972] Karl Michael Göschka and Seif Haridi, editors. Distributed Applications and Interoperable Systems: 12th IFIP WG 6.1 International Conference, DAIS 2012, Stockholm, Sweden, June 13–16, 2012. Proceedings, volume 7272 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30822-8 (print), 3-642-30823-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
30823-9.

Giese:2012:FTD

[973] Holger Giese and Grigore Rosu, editors. Formal Techniques for Distributed Systems: Joint 14th IFIP WG 6.1 International Conference, FMOODS 2012 and 32nd IFIP WG 6.1 International Conference, FORTE 2012, Stockholm, Sweden, June 13–16, 2012. Proceedings, volume 7273 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30792-2 (print), 3-642-30793-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-30793-5.

Sirjani:2012:CML

[974] Marjan Sirjani, editor. Coordination Models and Languages: 14th International Conference, COORDINATION 2012, Stockholm, Sweden, June 14– 15, 2012. Proceedings, volume 7274 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30828-7 (print), 3-642-30829-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-30829-1.

Schindler:2012:CSC

[975] Werner Schindler and Sorin A. Huss, editors. Constructive Side-Channel Analysis and Secure Design: Third International Workshop, COSADE 2012, Darmstadt, Germany, May 3–4, 2012. Proceedings, volume 7275 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29911-3 (print), 3-642-29912-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-29912-4.

Klasing:2012:EAI

[976] Ralf Klasing, editor. Experimental Algorithms: 11th International Symposium, SEA 2012, Bordeaux, France, June 7–9, 2012. Proceedings, volume 7276 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30849-X (print), 3-642-30850-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
30850-5.

Koucheryavy:2012:WWI

[977] Yevgeni Koucheryavy, Lefteris Mamatas, Ibrahim Matta, and Vassilis Tsaoussidis, editors. Wired/Wireless Internet Communication: 10th International Conference, WWIC 2012, Santorini, Greece, June 6–8, 2012. Proceedings, volume 7277 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30629-2 (print), 3-642-30630-6 (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-30630-3.

Domenach:2012:FCA

[978] Florent Domenach, Dmitry I. Ignatov, and Jonas Poelmans, editors. Formal Concept Analysis: 10th International Conference, ICFCA 2012, Leuven, Belgium, May 7–10, 2012. Proceedings, volume 7278 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29891-5 (print), 3-642-29892-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-29892-9.

Sadre:2012:DNS

[979] Ramin Sadre, Jiří Novotný, Pavel Čeleda, Martin Waldburger, and Burkhard Stiller, editors. Dependable Networks and Services: 6th IFIP WG 6.6 International Conference on Autonomous Infrastructure, Management, and Security, AIMS 2012, Luxembourg, Luxembourg, June 4–8, 2012. Proceedings, volume 7279 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30632-2 (print), 3-642-30633-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-30633-4.

Helleseth:2012:STA

[980] Tor Helleseth and Jonathan Jedwab, editors. Sequences and Their Applications — SETA 2012: 7th International Conference, Waterloo, ON, Canada, June 4–8, 2012. Proceedings, volume 7280 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-30614-4 (print), 3-642-30615-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-30615-0.

Alvarez:2012:FIF

[981] Federico Álvarez, Frances Cleary, Petros Daras, John Domingue, Alex Galis, Ana Garcia, Anastasius Gavras, Stamatis Karnourskos, Srdjan Krco, Man-Sze Li, Volkmar Lotz, Henning Müller, Elio Salvadori, Anne-Marie Sassen, Hans Schaffers, Burkhard Stiller, Georgios Tselentis, Petra Turkama, and Theodore Zahariadis, editors. The Future Internet: Future Internet Assembly 2012: From Promises to Reality, volume 7281 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30240-8 (print), 3-642-30241-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-30241-1.

Isokoski:2012:HPDa

[982] Poika Isokoski and Jukka Springare, editors. Haptics: Perception, Devices, Mobility, and Communication: International Conference, EuroHaptics 2012, Tampere, Finland, June 13–15, 2012. Proceedings, Part I, volume 7282 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31400-7 (print), 3-642-31401-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
31401-8.

Isokoski:2012:HPDb

[983] Poika Isokoski and Jukka Springare, editors. Haptics: Perception, Devices, Mobility, and Communication: International Conference, EuroHaptics 2012, Tampere, Finland, June 13–15, 2012 Proceedings, Part II, volume 7283 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31403-1 (print), 3-642-31404-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
31404-9.

Bang:2012:PTD

[984] Magnus Bang and Eva L. Ragnemalm, editors. Persuasive Technology. Design for Health and Safety: 7th International Conference, PERSUASIVE 2012, Linköping, Sweden, June 6–8, 2012. Proceedings, volume 7284 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31036-2 (print), 3-642-31037-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-31037-9.

Snoeyink:2012:FAA

[985] Jack Snoeyink, Pinyan Lu, Kaile Su, and Lusheng Wang, editors. Frontiers in Algorithmics and Algorithmic Aspects in Information and Management: Joint International Conference, FAW-AAIM 2012, Beijing, China, May 14–16, 2012. Proceedings, volume 7285 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29699-8 (print), 3-642-29700-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-29700-7.

Peffers:2012:DSR

[986] Ken Peffers, Marcus Rothenberger, and Bill Kuechler, editors. Design Science Research in Information Systems. Advances in Theory and Practice: 7th International Conference, DESRIST 2012, Las Vegas, NV, USA, May 14–15, 2012. Proceedings, volume 7286 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29862-1 (print), 3-642-29863-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-29863-9.

Agrawal:2012:TAM

[987] Manindra Agrawal, S. Barry Cooper, and Angsheng Li, editors. Theory and Applications of Models of Computation: 9th Annual Conference, TAMC 2012, Beijing, China, May 16–21, 2012. Proceedings, volume 7287 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29951-2 (print), 3-642-29952-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
29952-0.

Kranakis:2012:FAI

[988] Evangelos Kranakis, Danny Krizanc, and Flaminia Luccio, editors. Fun with Algorithms: 6th International Conference, FUN 2012, Venice, Italy, June 4–6, 2012. Proceedings, volume 7288 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-30346-3 (print), 3-642-30347-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-30347-0.

Bestak:2012:NIIa

[989] Robert Bestak, Lukas Kencl, Li Erran Li, Joerg Widmer, and Hao Yin, editors. NETWORKING 2012: 11th International IFIP TC 6 Networking Conference, Prague, Czech Republic, May 21–25, 2012, Proceedings, Part I, volume 7289 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30044-8 (print), 3-642-30045-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-30045-5.

Bestak:2012:NIIb

[990] Robert Bestak, Lukas Kencl, Li Erran Li, Joerg Widmer, and Hao Yin, editors. NETWORKING 2012: 11th International IFIP TC 6 Networking Conference, Prague, Czech Republic, May 21–25, 2012, Proceedings, Part II, volume 7290 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30053-7 (print), 3-642-30054-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-30054-7.

Becvar:2012:NWI

[991] Zdenek Becvar, Robert Bestak, and Lukas Kencl, editors. NETWORK-ING 2012 Workshops: International IFIP TC 6 Workshops, ETICS, HetsNets, and CompNets, Held at NETWORKING 2012, Prague, Czech Republic, May 25, 2012. Proceedings, volume 7291 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30038-3 (print), 3-642-30039-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-30039-4.

Bleris:2012:BRA

[992] Leonidas Bleris, Ion Măndoiu, Russell Schwartz, and Jianxin Wang, editors. Bioinformatics Research and Applications: 8th International Symposium, ISBRA 2012, Dallas, TX, USA, May 21–23, 2012. Proceedings, volume 7292 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30190-8 (print), 3-642-30191-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
30191-9.

Schrijvers:2012:FLP

[993] Tom Schrijvers and Peter Thiemann, editors. Functional and Logic Programming: 11th International Symposium, FLOPS 2012, Kobe, Japan, May 23–25, 2012. Proceedings, volume 7294 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29821-4 (print), 3-642-29822-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-29822-6.

Simperl:2012:SWR

[994] Elena Simperl, Philipp Cimiano, Axel Polleres, Oscar Corcho, and Valentina Presutti, editors. The Semantic Web: Research and Applications: 9th Extended Semantic Web Conference, ESWC 2012, Heraklion, Crete, Greece, May 27–31, 2012. Proceedings, volume 7295 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30283-1 (print), 3-642-30284-X (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-30284-8.

Li:2012:AGP

[995] Ruixuan Li, Jiannong Cao, and Julien Bourgeois, editors. Advances in Grid and Pervasive Computing: 7th International Conference, GPC 2012, Hong Kong, China, May 11–13, 2012. Proceedings, volume 7296 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30766-3 (print), 3-642-30767-1 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-30767-6.

Maglogiannis:2012:AIT

[996] Ilias Maglogiannis, Vassilis Plagianakos, and Ioannis Vlahavas, editors. Artificial Intelligence: Theories and Applications: 7th Hellenic Conference on AI, SETN 2012, Lamia, Greece, May 28–31, 2012. Proceedings, volume 7297 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30447-8 (print), 3-642-30448-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-30448-4.

Beldiceanu:2012:IAT

[997] Nicolas Beldiceanu, Narendra Jussien, and Éric Pinson, editors. Integration of AI and OR Techniques in Contraint Programming for Combinatorial Optimzation Problems: 9th International Conference, CPAIOR 2012, Nantes, France, May 28–June 1, 2012. Proceedings, volume 7298 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29827-3 (print), 3-642-29828-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-29828-8.

Chau:2012:ISI

[998] Michael Chau, G. Alan Wang, Wei Thoo Yue, and Hsinchun Chen, editors. Intelligence and Security Informatics: Pacific Asia Workshop, PAISI 2012, Kuala Lumpur, Malaysia, May 29, 2012. Proceedings, volume 7299 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30427-3 (print), 3-642-30428-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
30428-6.